A Complete Bibliography of Publications in the IEEE symposia on Foundations of Computer Science (FOCS)

for 1990–1999

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

08 August 2024 Version 2.05

Title word cross-reference

(2 + ϵ)n [DZ96]. (q) [GK91]. 0 [MS99]. 1 [AF98, MS99]. 2 [yCLZ94, Fre91, Hås93, Jai98, NZ97]. 3 [BE95, Blu90b, Gar96, GR98, GLST98, HJP93, Kog96]. 5/2n^{2} [Bla99]. 6/5 [Ben95]. 7/8 [KZ97]. 8 [ENZ96]. ^{0} [AMRT96, BPR97]. 1 [SYY99]. ^{3/2} [SZ95]. _{e} [SZ95]. N C [KK94]. N P [PT96]. K [KT99a, AGHP90, BR92, BK93, CG99, CHLT93, CT96, Dey97, Epp94, FRR90, Fre91, Gar96, JV99, Kou99, LS92a, MPA92, PPSZ98, SV91, Sch99]. L^{1} [FKSV99]. l_{1} [GNRS99]. L_{p} [AAG^{+}95]. ≤ k [Cha98]. log n + O(1) [RZ98]. m [Gro92, GT98]. n^{1−ϵ} [Hås96]. ≠ [Hak95]. n log n [RS91]. n × n [Bla99]. O(1.3446^{n}) [BE95]. O(distance + 1/sessionrate) [AFHB^{+}97]. O(lg^{3/2 mod V} mod) [JM91]. O(log^{1.5} n) [NSW92]. O(n^{1+ϵ} log b) [NR94]. O(n log^{2} n) [AW92]. O(n log^{3} n) [Rei93]. o(n log n) [BAG93]. ω [Kla91]. O˜(mR∘) [Fei93]. p [Egi93, Gro92, GT98]. p < 0 [BO94]. s [BE93]. Σ_{2}^{p} [Uma99]. t [BE93, Coh93]. θ [Sze94]. θ(log n log log n) [AMRT96].

-adic [Egi93]. -approximation [ENZ96, Gar96, KZ97, NZ97]. -automata [Kla91]. -coloring [BE95, Blu90b]. -conjecture [RS93]. -connected [CT96].

-cover [RV93]. -cuts [SV91]. -difference [FKSV99]. -edge-connectivity [Fre91]. -embeddings [GNRS99]. -flows [Wei94]. -Frege [BPR97]. -generated [yCLZ94]. -level [KT99a]. -levels [Cha98, Dey97]. -link [MPA92]. -lower [Bla99]. -matrix [Bla99]. -median [CG99, JV99]. -SAT [PPSZ98, Sch99]. -server [BR92, FRR90, Kou99]. -set [CHLT93, RV93]. -sets [Dey97]. -spanners [Coh93]. -way [AF98]. -wise [AGHP90].

31st [IEE90]. 32nd [IEE91]. 33rd [IEE92]. 34th [IEE93]. 35th [Gol94]. 36th [IEE95]. 37th [IEE96]. 38th [IEE97]. 39th [IEE98]. 3SAT [KZ97].

40th [IEE99].

Abelian [VSR96]. abstract [Gar92]. AC [AMRT96, BS90]. ACC [BT91, Yao90]. accept [GP93]. access [AZ98, FMRW95, MKR95, Nay99, PR96]. according [KL96]. Achievements [Pnu97]. action [ACM98]. acyclic [GLS98]. Adaptive [AF91, AST99, BV99, Sah99]. addition [HI91, PPZ90]. adic [Egi93]. admission [AGLR94]. adversarial [ACBFS95, Gam98]. adversaries [FKM^{+}95, Kou99]. affine [ES93a]. against [AN90, FKM^{+}95]. agreement [CHLT93, KA94, Sch95, GMY95]. Algebraic [BO94, Ede95, LFKN90, Yao92, AB91, ALRS92, BPR94, BCE91, GKV95, GS98, KP97b, Koi97, LT90, PRT92, Rab96b, Wat99a]. algebraic-geometric [GS98]. algebras [GR98]. algorithm [AM97, AFB93, AS90, AL93, Bar93, Bas97, BE95, BC93, BFKV96, Bou98, Cha91, Cha97, DKLR95, ET91, ENZ96, FKSV99, GKP93, Gar92, Gaz91, GK91, HP95, HJM94, HR91, Jac94, Jai98, JV92, KZ97, KS93, NGM90, NZ97, NR94, OP98, PPSZ98, Rei93, Sch99, VF90, Var98, VSR96, Vem97].

Algorithmic [Adl94, GIP99, Alo91, Alo92, AV99, Bar96, LLR94]. Algorithms [Bea95, Sho94, Sze98, AB91, AADW94, AMW95, ACM98, AV94, AGR94, Amb99, AEIS99, ABZ96, AMS94, ALS90, BR92, BV99, BB93, BJK^{+}90, BG99, BCF^{+}99, BM95, BCDZ99, CFKL94, CDZ98, CG99, Cla94, Coh93, CR90, CCG^{+}93, DG95, Epp92, ENRS95, FCI99, FRR90, FFK^{+}91, FW90, FKM^{+}95, FKV98, FLPR99, Gab93, GK98, GZ94, Gie92, GMV91, Goo91, HI91, Ind98a, JV99, Kan94, KA94, Kin99, KT99b, Koi97, KS97, Kos95, KP94, Lag90, LS91b, LRT93, LRS95, MR95, Mul92, Ogi95, PRT92, PST91, RS98, Rad93, RV93, Zwi98]. all-nearest-neighbors [Cal93]. all-pairs [KKP91, Kin99]. allegory [BPT90]. allocation [AA94, CS97, Goo91, LS91b, LU99]. Almost [Sha93, AM90, AGHP90, DKS98, DHZ96, FGL^{+}91, MSP90, Zwi98, vD98]. almost-polynomial [DKS98]. alphabet [GP92]. alphabet-independent [GP92]. alphabets [Far97]. Alternating [AHK97]. Alternating-time [AHK97]. alternative [Bea95]. Alto [IEE93, IEE98]. Ambivalent [Fre91]. Amortization [Oka95]. Amortized [FKN91, Cha99, ST98]. Amplification [DZ92, GKS90, GIL^{+}90]. amplitudes [Wat99a]. analog [TV94]. analysis [AM97, AADW94, AMW95, AB97, EYFKT92, FR92, KP94, KRRT98, RSW98, Tor95, VF90]. analytic [Smo90]. ancestor [AHR98]. angles [Mit93]. annealing [JS93]. Annual [Gol94, IEE90, IEE92, IEE93, IEE95, IEE96, IEE97, IEE98, IEE99, IEE91]. anyway [PT98]. apparatus [MY98]. Apple [HLL92]. Application [BGV95, CDZ98, ES93b, NR95]. Application-controlled [BGV95]. applications [AB91, AST99, AB97, AS94, AFK96, Bar96, BBR97, BBB^{+}96, Ben95, FK98, Ind97, Kar93, KT99a, Kla91, KD90, LLR94, Lok95, Pla95, SW98, Ver91]. approach [BFU96, Cha94, DH90, EC91, FGY93, MTV91, Vem98]. Approximability [Hal95, BGS95, KMSV94]. Approximate [BF91, CJS96, CL90, Coh92, EKR96, FCI99, KMS94, KR96a, ABSS93, Blu90b, Cha93a, FKSV99, Hås96, Ind98b, Mic98, PRT92, SV96, Sch95]. Approximating [CCG^{+}98, CT96, DKS98, FGL^{+}91, Fle99, Koi95, TY93, VA97, DMS99, Fur95, Rav94, Uma99]. Approximating-CVP [DKS98]. Approximation [ABC^{+}99, BGK^{+}98, KARR90, KT99b, AFW94, AM96, ALM^{+}92, Aro96, Aro97, AL93, Bar96, CDZ98, CG93, ENRS95, ENZ96, FK96, Gar96, GGR96, Goo91, GKP95, GK91, Ind99, Jai98, JV99, JV92, JKL98, KZ97, KS97, MR95, NZ97, PST91, RV93, TSSW96, Ung98, dlVK98]. approximations [FKV98, MWW91, NS97, Sri97]. arbitrary [Bla99, GK91]. area [BB90]. area-universal [BB90]. arithmetic [BF90, CS92, Ebe92, GR98, NW95]. armed [ACBFS95]. arrangement [AFK96, HP99]. arrays [KKL^{+}90]. Art [Mat98, HKK91]. Arthur [MV99]. Artificial [Kea98]. aspects [Alo92, GIP99]. assignment [AFK96, Pap91]. assumption [IW98]. asymmetric [AM98]. asymmetry [Voc99]. asymptotic [MPT92]. Asymptotically [ABO91, KKL^{+}90]. Asynchronous [MSP90, AR92, AKPR93, IL94, Wan91]. attributes [DH94]. augmentation [BJJ97, HR91]. Augmenting [Fra90]. automata [AF98, BBB^{+}96, BPT90, DS96, EJ91, ELW90, Gac97, KF90, Kla91, KW97, MP90, Nay99, Wat95]. automaton [Wan91]. autoreducibility [BFT95]. availability [NW94]. average [ABC^{+}99, Bar95, CFKL94, CN97, Gur90, MRSG99, Orl91, Sho91, VF90]. average-case [CFKL94, CN97, Orl91, Sho91]. axis [dBO90]. axis-parallel [dBO90].

Back [AH92]. balanced [LU99, MR97, Orl91]. balancing [AAG^{+}95, ABK92, GI99, KRT99, Mit96, Mul91b, RSW98, Voc99]. band [Ebe92]. bandit [ACBFS95]. bandwidth [Adl96, MadHVW97, MKR95, Ung98]. barrier [GR97a]. based [FLS90, KST99, MR95, Vem97]. basis [BMP^{+}99]. be [CGG98]. Beach [IEE97]. beam [RTY90]. becomes [Kog96]. behavior [GP93]. best [AW95, Sho91]. Better [ES93a, Kah91, Rad91, Amb99, DZF91, Imp90, Sho91]. between [BEGJ98, CG90]. Beyond [GR97a, KP94, GMY95]. bi [HF93]. bi-weighted [HF93]. biconnectivity [HK95, Rau92]. bijections [SZ90]. bin [Sho91]. Binary [AGMV96, HF93]. bisection [JS93, Kun91]. bisimulation [Sen98]. bit [BCJL93, CG98, Fri92, KWZ95]. bit-reversal [CG98]. bits [BGS95, HN98, NRS94]. Bivariate [Bla98]. blocking [LRT93, LU99]. Blum [Koi93]. body [And96a]. Boolean [Abr90, AT96, Ajt99, AGMN92, Blu90a, BBL98, DZ92, Gal91, HW97, KLM94, KP95b, MSS91, PPZ90, TV94]. Boosting [KS99, AD93]. bootstrapping [BST92]. both [BCJL93]. bottleneck [FFM98]. bottlenecks [Hak95]. bound [Ajt99, Amb99, Bla99, Blu90a, CHLT93, ES93b, Gil93, GKV95, Ind98a, Kun91, MR96, PR99, RS91, Sun91, Sze94, Yao94]. boundary [OP98]. bounded [BFU96, CCM98, CG93, CR93, FKM^{+}95, Kil94, Kle91, LMS92, MWW91, MMS98, Mor91, PS95, TV94, Wat99a, adHS96]. bounded-degree [LMS92]. bounded-interaction [Kil94]. Bounds [BCDZ99, CG90, AKRS92, ADFK97, AD93, BE93, BBC^{+}98, BIK^{+}94, BIP95, BP96, BGP95, BAG91, BKRS92, Cha94, CH92, CS92, Dey97, DH90, EIRS91, ES93a, Eri96, Fis97, Gal91, GV94, GR98, Gri98, GT98, HJP93, HR92, HS93a, HRG96, KKL^{+}90, KMPS94, KKP91, KPR98, KW91, MP90, Nay99, NW95, PPS92, Pla95, Poo93, Rad91, RTS97, RS98, RSW94, Sha93, Sho91, SS91, Smo90, Vig99]. BPP [AS94, ACRT97, Zuc91]. branching [Ajt99, BST98, ST94]. broadcast [Rav94]. buffers [BFU96, adHS96]. building [AV91]. bulk [AA97]. Burlington [IEE96]. butterfly [CMS95, Tam92]. Buy [AA97]. Buy-at-bulk [AA97]. Byzantine [GMY95].

Cache [FLPR99, BGV95]. Cache-oblivious [FLPR99]. caching [KK96, Tor95]. calculus [BSI99, BG99, EJ91, PU92]. calendar [ELL94]. California [IEE93, IEE98]. can [BAG93, BS94, CGG98]. capacity [CR98, GR97b, NW94]. Carlo [BCF^{+}99, DKLR95, FKV98]. Carnap [Got93]. cascade [BCK96]. cascaded [MP90]. case [Bar95, CFKL94, CN97, GRS95, Gur90, Orl91, Sho91, VF90]. cases [KP92]. casino [ACBFS95]. catenable [BST92]. catenation [Oka95]. cell [Bas98]. cellular [DS96, Gac97, Wan91, Wat95]. certain [Vai90]. chain [BCF^{+}99, KR99, LRS95]. chains [BD97, Kan94, LS90, RSW98]. Challenges [Pnu97]. channels [AM98]. chaos [BPT90]. character [AFB93]. characteristic [BCH^{+}95, BO94, KP92, Kog96]. characteristics [Yao92]. characterization [AS92, BF90, GLST98, KC91, Lei98, MMS98, TW96]. checkable [ST98]. Checking [BEG^{+}91, AS92, AV91, APSV91, BW94, EKR96]. Chernoff [Gil93]. Chinese [AB99]. choosing [KP93]. chordal [KST94]. chosen [Sah99]. chosen-ciphertext [Sah99]. chromatic [Fur95]. ciphertext [Sah99]. circuit [AGLR94, BT91, CHH99, EIRS91, LP90, Yao93]. circuits [CSS97, Gal91, GKS90, GR98, Gro92, GT98, HG90, HJP93, KLM94, KW91, LMP91, MSS91, NW95, PPZ90, PS90, RS91, Smo90, TV94, Yao90]. circulant [AB91]. City [IEE99]. clairvoyance [KP95a]. class [CC92]. classes [AS94, BFT95, Pap90]. classical [Wat99a]. classification [KT99b]. clause [MP92]. clauses [AFP90]. Clique [Hås96, FGL^{+}91]. Clock [AR92]. closed [Bas97]. closest [LS92a]. closure [Coh94, DDPY94, GM92, HK95, Kin99]. clustering [Ind99]. clusters [Sze98]. CMOS [CHH99]. CNFs [BSI99]. coalitions [AN90]. code [DMS99]. codes [AEL95, ABSS93, Aro95, GS98, Nay99, SS94, Sud96, VSR96]. Coding [OR95, PPZ97, Adl96, Ind97]. Cognitive [Val95]. Coin [AN90, AKV96]. Coin-flipping [AN90]. coin-weighing [AKV96]. collect [AST99]. Coloring [Ira90, AM97, BE95, Blu90b, FK98, KMS94, Rab96a, RS98, Sub95, Vis90]. colorings [Vig99]. combinatorial [ASWZ96, Bar95, BPR94, Bas98, CG99, Rad92]. combining [PRT92]. commitment [BCJL93]. commodities [Fle99]. common [AS91]. Communication [ACK90, BTY90, EIRS91, LT90, WK91, AM98, ASTS^{+}98, AGLR94, CG90, FKN91, Gro92, KW91, Lok95, Orl91, RS93]. Communication-optimal [ACK90]. Communication-space [BTY90]. commutative [yCLZ94]. compact [MKS93]. comparison [FT94]. comparisons [DZ96]. Competitive [EYFKT92, FRR90, FFK^{+}91, FMRW95, AADW94, AB97, AKRS92, ABF93, AAP93, AA94, BR92, BBK99, KP94]. complementation [Kla91]. complete [FGL^{+}91, Gro96, Gur90, SV97]. completeness [KMO94, VVAG92]. completion [ABC^{+}99, KST94]. complex [NR94]. complexities [Gro92]. Complexity [GV94, KP90, Adl94, AS94, ASTS^{+}98, BPR94, Bas98, BT91, BFT95, yCLZ94, CG90, CR92a, CH92, CGG90, DC91, EIRS91, Egi93, FKN91, Gal91, GMY95, GP91, Gol97, GLS98, GR98, HF93, HLP97, IPZ98, JL93, Kil94, Lat95, LMN97, LV99, Lok95, LS92b, LT90, MT90, MP90, MPT92, Orl91, Pap90, PR99, PT96, PU92, RS93, RTY90, Sch98, ST98, Tod90, Ung98, Vai90, WK91, Yao93]. components [JM91, KTDC91]. compression [BST92, Cov96, VK91]. computability [Hav97, RTY90]. Computable [Hav97, CIK93]. Computation [Mat98, BO94, CG96a, CR93, GGL91, GV94, HS93b, LT90, MMS98, RS91, Sho94, Sho96, Sim94, Spi96, TV94, Val95]. computational [BG92, GTVV93, HLP97, KMSV94, MSS91, MMP^{+}91, PT98]. Computationally [Rab96b, BIN97, FKM^{+}95, KO97]. computationally-private [KO97]. computations [BL94, CGGK90, CS92, KMO94]. Computer [Gol94, IEE90, IEE91, IEE92, IEE93, IEE95, IEE96, IEE97, IEE98, IEE99]. computers [ABO96, FRU92]. Computing [Blo91, ELW90, GI91, HRG96, HHK95, KP97b, Kog96, Luk92, MPA92, SZ94, AGLR94, BMP^{+}99, DC91, KKL^{+}90, KP95b, LS90, OR95, PRT92]. concave [KT99a, Tol92]. Concentrated [Kun91]. concentrators [Mor91]. concepts [AV99, BCH94, KS90]. concrete [BCK96, BDJR97]. Concurrent [dAHK98, GP93, PR96]. conductance [KR99]. congestion [LU99]. congruences [HW96]. conjecture [CS95, DH90, FFK92, KMPS94, RS93]. conjunctions [AFP90]. conjunctive [GLS98]. Connected [JM91, CT96, KTDC91]. connecting [Hsu92]. connection [CN97, GGR96, LPR94, Pla95]. connection-oriented [LPR94]. connectivity [BJJ97, BE93, BIP95, Ben95, Fra90, Fre91, HRG96, Kar93, NGM90, NSW92, Poo93, Sch94, Yao94]. conquer [ENRS95, Var98, WK91]. consensus [AW92]. Constant [CSS97, AM90, ANKS99, FR99, GMV91, GM97, KLOW94, Mic98, Wat99b]. constant-round [Wat99b]. constants [Gut98, Pla99]. constraint [Sch99]. constraints [CP94, GTT93]. Constructing [Ist90, Coh93, VSR96]. construction [AGHP90, AR92, BCK96, Cha98, Far97, FFM98, KK94, Kos95, Mor91, NR95, PR99, Wag92]. constructions [ABCP93, NR97]. constructive [BMP^{+}99]. contact [HRS93]. contain [KLM94]. Contention [GM97, PS95, AAF^{+}96, HSW91]. contention-resolution [AAF^{+}96]. context [HJM94]. context-free [HJM94]. contribution [Adl94]. control [AL93, AGLR94, BBR97, MPS98]. Controllability [Lat95]. controlled [BGV95]. convergence [ABDCBH93]. conversion [ISW99]. convex [AGR94, AS93b, Cha99, Cha91, CR93, Ede95, KN98, KP97b, MKS93]. convolution [Ind98a]. coordination [IR93]. core [Imp95, KS99, LRT93]. correcting [Wan91]. correction [APSV91, JKL98, KS96]. correctness [BEG^{+}91]. correlated [Orl91]. cost [ABNK^{+}92, ABCP93, Gab93, Var98]. cost-scaling [Gab93]. counterexamples [MT90]. Counting [CEG^{+}90, Hak95, HI93, Bar93, DFJ99, HSW91]. coupled [BPT90]. coupling [BD97, KR99]. cover [RV93]. covering [PST91, RV93]. covers [ABCP93, LRT93]. CPU [YDS95]. creating [VVAG92]. CREW [JM91]. crossing [PT98]. cryptocomputing [SYY99]. cryptography [Gol97, MY98]. cryptosystems [FGMY97]. CS [Mic94]. cubes [GZ94, MS99]. curve [HI91]. curves [DG95, HI93, vzGS95]. cut [NZ97, dlVK98]. Cuts [GNRS99, Ben95, GSV94, SV91]. cutting [BEGJ98, CEG^{+}90]. cuttings [Cha91]. CVP [DKS98]. cycle [Hen94]. cycle-equivalence [Hen94]. cycles [CEG^{+}90, UL97].

Data [BST92, AEIS99, And96a, ALRS92, AB96, AS91, BAG91, Cov96, FKSV99, Fre91, JKL98, Kun91, MadHVW97, VK91]. database [KO97, Yan95]. databases [KMRV97]. Datalog [GM92]. dates [ABC^{+}99]. de-randomization [IW98]. deadlock [AA94]. Decades [Pnu97]. Decidability [Sen98]. Deciding [HR90, SS97, HJM94]. Decision [LMSS90, BM95, DS96, GKV95, LNNW91, NRS94, RS91, Yao92]. decoding [GS98, Sud96]. decohered [ABO96]. decomposition [BKRS92, Cal93, Ede95, GR97a, Gur90, Lag90, MP90]. decrease [CR93]. definable [Koi95]. degeneracies [EC91, ES93a]. degree [LMS92, Sen98, Smo93]. delay [AFHB^{+}97, GM97, PS95]. Delayed [ACM98]. Delsarte [Sze94]. denest [Blo92]. dense [AFW94, AFK96, FK96, KP93, OP98]. densely [KT95]. density [KL96, Mit96]. dependent [Mit96]. depends [AS93a]. depth [BIP95, BCE91, CSS97, Coh92, CS92, EIRS91, ES93b, GR98, HG90, HJP93, HH90, KW91, KLOW94, Lok95, RTS97, Yao94]. depth-two [RTS97]. derandomization [BCM93, NSS95]. Derandomizing [MR95, MV99]. derivatives [NW95]. design [AZ98, AA97, CHH99, Goo91, KLM94]. detecting [ES93a]. determinacy [EJ91]. determinant [Cla92]. determine [MMP^{+}91]. Deterministic [BB90, Ind97, adHS96, And96b, AMS94, Cha97, Gaz91, Koi97]. detour [Kle91]. diagnosis [BHK95]. diagram [MKS93]. dial [CR98]. dial-a-ride [CR98]. diameter [AMS94, BHK^{+}90, Ist90, LS92b]. diameters [BGK^{+}98]. dichotomous [FW90]. dictionaries [AO91, AMRT96]. dictionary [AF91, Sun91]. difference [FKSV99]. difficult [Kog96]. difficulty [CR93]. digraphs [Kin99]. dimension [AM90, Bar93, Koi97, MWW91, Tol92]. dimensional [AGR94, Bar95, Epp91, GP92, KT99a, Wan91, Wat95]. dimensions [ABDCBH93, AS93b, Cha98, CCG^{+}93, Sha93]. dining [AS90]. Directed [HRS93, BE93, BS94, ELW90, MKR95, NZ97, Zwi98, FR99]. directions [Gol97]. Discrepancy [ASWZ96, MWW91, Cha93b]. discrete [Sho94]. discretized [BCH94]. Disjoint [KT95, AAB^{+}92, Sri97, Wei99]. disjunction [AW95]. disk [ABZ96]. distance [BIP95, DMS99, HP95]. Distributed [AV91, PR90, AADW94, AKRS92, AS91, ABF93, ABCP93, AA94, BR92, FGY93, GKP93, LS91b, NR91, PR99, Pla95]. distribution [Blu90a, BK93, Jac94, JL93, KS90]. distribution-free [Blu90a, KS90]. distributional [LPR94]. distributions [Imp95, Orl91]. divergence [RSW98]. Divide [ENRS95, Var98, WK91]. Divide-and-conquer [ENRS95, Var98, WK91]. division [BP92]. DNA [KW90, Li90]. DNF [AP91, AHP92, Jac94, Uma98]. Does [BIN97]. doesn’t [BGK^{+}98]. domain [Blu90a]. domains [CR92a]. double [BST92, CHH99]. down [HJP93]. Drawing [FHH^{+}90, Kan92]. DSPACE [SZ95]. dual [JV99, RV93]. Dump [ABF93]. Dynamic [AEM92, Cha99, Epp91, FST91, FMS93, Sch91, AFHB^{+}97, AV96, BFU96, CJ90, Epp92, FT94, FMRW95, Fre91, Hen94, HK95, Kin99, LS91b, Mul91b, Mul91a, Rau92, SV96]. dynamical [RSW92].

easier [BBFG91]. easy [Sch91].

Eavesdropping [FGY93]. Edge [BJJ97, Wei99, Ben95, Fra90, Fre91, NGM90, Sri97]. Edge-connectivity [BJJ97, Fra90, NGM90]. Edge-disjoint [Wei99, Sri97]. effective [Cla92, TW96]. efficiency [KST99]. Efficient [ABNK^{+}92, AKFS99, AEIS99, CFKL94, EHJ^{+}96, FKM^{+}95, HS93b, HI91, KS90, KPS92, KS96, Lag90, LS91b, LRT93, MKR95, Rei95, SV96, ST94, Tam92, vzG91, AN94, AKPR93, BR94, Ebe92, Jac94, KP92, KA94, NR97, Ogi95, VSR96]. Efficiently [SZ90]. eigenvalue [Kah92, Rei95]. eigenvector [Rei95]. election [RZ98]. element [ST96]. elementary [GV94]. eliminate [Mit93]. elimination [BPR94, Bas97, BMMR97]. embedded [KT95]. embedding [Tam92]. embeddings [FCI99, GNRS99]. emptiness [Eri96]. emulation [ABO91]. emulations [KKR93]. encoding [ABE^{+}94]. encryption [BDJR97]. ended [BST92]. energy [YDS95]. entropy [CR93]. Enumerating [LS92a]. envelopes [Sha93]. environment [DKP91, HS93a]. equalities [Ge93]. equational [Sen98, Ver91]. equations [ABSS93, EKR96, Gut98, Pla99, Rub94, Woo98]. Equivalence [Gro96, AHP92, Hen94, HJM94, Sen98, SPLK93]. equivalent [Uma98]. erasure [AEL95]. Error [RRV99, BIN97, BCDZ99]. Estimating [Coh94]. estimation [DKLR95]. Euclidean [Aro96, Aro97, DZF91, HS93b, Ind98b, SPLK93]. Euler [CHH99, Yao92]. evaluation [Cla92, KD90, Oka95, PRT92, SS91]. even [CCKV97]. evolution [ADFK97]. Evolutionary [CGG98, FT94, KW90, KA94]. evolving [KMRV97]. Exact [AP91, Bsh93, FR92, GKS90, BC92, CH92, CGG90, Zwi98]. example [Epp99]. execution [AKPR93]. Expander [SS94, Gil93, KR96b]. expanders [AKS92]. expansion [Kah91, Kah92]. expected [AW92, CL90, GM97, Sub95]. Explicit [Mor91]. Exploiting [MadHVW97]. exploration [BS94]. Exploring [DP90]. exponent [Hås93]. Exponential [BEGJ98, GR98, BFL90, Gut98, IPZ98, JV92, PPSZ98]. exponential-time [PPSZ98]. exponentiation [vzG91]. expression [KD90]. extension [AV90]. extensions [BP92]. External [GTVV93, AV96]. External-memory [GTVV93]. extraction [Fri92]. extractors [RRV99].

facility [CG99, JV99]. factor [Jai98]. Factoring [MR96, SS97, Sho94]. factorization [OP98]. factors [DKS98]. fairly [LP90]. Fairness [KRT99]. fan [KW91, TV94]. fan-in [KW91, TV94]. Fast [Coh93, FKV98, Gie92, IL94, KLOW94, PR96, PST91, Yap92, ALS90, CR93, CCG^{+}93, NGM90]. Faster [AO91, And96b, DGM90, GK98, Ind98a, Kos95, PPZ90, Rad93, Sch95, Cha97]. Fat [MMP^{+}91, AGMV96]. Fault [AAB^{+}92, AU90, AB96, BHK95, GGL91, JCT92, Sho96, BMP^{+}99, KP95c, LMP91, LMS92, PR96, Spi96]. Fault-tolerant [GGL91, JCT92, Sho96, BMP^{+}99, LMP91, PR96, Spi96]. faults [CMS95, Tam92]. faulty [ABO91, KKL^{+}90]. Fe [Gol94]. feasible [BPR97, KLOW94]. feedback [ENZ96]. fiber [MKR95]. field [Woo98, vzG91]. fields [Bas97, Bla99, GR98, HI93, KT99b, Kog96, vzGS95]. file [NR91]. files [Orl91]. fill [BMMR97, KST94]. fill-in [KST94]. financial [EYFKT92]. Finding [CHH99, CCKV97, Epp94, GSV94, KKP91, KK99, SV91, vzGS95, FK98, FKV98, Tod90]. Finely [BBK99]. Finely-competitive [BBK99]. finger [CR90]. Finite [KF90, AF98, BHK^{+}90, BF91, CCG^{+}98, CR98, GR98, Gro96, HHK95, HI93, KW97, Sen98, ST96, Sze98, VSR96, Woo98, vzG91, vzGS95]. Finite-memory [KF90]. finite-variable [Gro96]. fit [Sho91]. fixed [AFB93, AM90, Bar93, GKS90, Pla95, Tol92]. fixed-connection [Pla95]. flash [BHK95]. flipping [AN90]. Florida [IEE97]. flow [AL93, BBR97, Coh92, DGG98, Fle99, Gab93, GK98, GR97a, Hal95, KP97a, KARR90, Kle96, KS97, Rad93, Sri97]. Flows [GR97b, Wei94]. fluid [Gam98]. forms [Gie92]. formula [DDPY94]. formulae [BCE91, PPZ90, PZ91]. formulas [AP91, BC92, ES93b, HW97, Rad91]. Foundations [Gol94, IEE90, IEE91, IEE92, IEE93, IEE95, IEE96, IEE97, IEE98, IEE99]. four [Hsu92, KTDC91]. four-connected [KTDC91]. four-connecting [Hsu92]. fractional [Fle99, GK98, PST91, Rad92]. framework [BV99, Gab93]. Free [BGS95, ALS90, Blu90a, Hav97, HJM94, JCT92, KS90, KP90]. Frege [BPR97]. full [GGL91]. Fully [Hen94, HK95, Kin99, LS91a, Rau92, AR92]. function [GV94]. functional [EKR96, Rub94]. functionals [CIK93, KC91]. functions [AAB^{+}92, ALRS92, BCK96, BFKV96, BBL98, BS90, Cha93a, DZ92, DNRS99, GKS90, GR98, HRS93, KST99, KP95b, MRV99, NR95, NR97, PPZ90, ST94, Smo90, Tol92, TV94]. further [Mul91a]. future [AH92].

Gadgets [TSSW96]. gallery [HKK91]. Gambling [ACBFS95]. game [Pla95]. games [AN90, EYFKT92, FGY93, FKM^{+}95, IR93, MV99, dAHK98]. Gates [GP93, Gal91, GV94, KLM94, KW91, Smo90]. gathering [EHJ^{+}96]. Gauss [VF90]. Gaussians [Das99]. GCD [SPLK93]. General [AFHB^{+}97, AD93, Zuc90, Bou98, BFU96, CGG98, EC91, RS91, Zuc91]. generalizations [AMW95, AF98]. generalized [Ist90, Jai98, Rad93, Sze94]. generate [Imp90]. generated [yCLZ94]. generation [BEG90]. generators [ASWZ96, Mul92, RSC95]. Genome [BP93]. genomic [HP95]. Geometric [Cha93b, Mat98, SW98, AEM92, Aro96, Aro97, AMS94, BCH94, Cla94, GS98, MTV91, Mul92, Sch91]. geometry [AKV96, Bar95, GTVV93, LLR94, MMP^{+}91, PT98]. gets [CR92b]. getting [vD98]. GF [GK91].

Gilbert [DH90]. given [SZ90]. Global [BBR97, AA94]. Go [AV94]. good [BEG90, MSP90]. grained [AKPR93]. Graph [Sch94, AM97, AFK96, BBR^{+}90, CCKV97, DP90, Epp92, FFK^{+}91, FGY93, GKP95, Hsu92, JS93, KTDC91, KMS94, Kar93, MTV91, Mit93, Pap90, PRST94, Poo93, Vis90]. graph-theoretic [FGY93, Pap90]. graphs [AAB^{+}92, AFW94, AKFS99, BS94, CHH99, DFJ99, FK98, FHH^{+}90, Fra90, GSV94, Gaz91, Gil93, GNRS99, Hen94, HHK95, Ira90, Ist90, Kah91, Kah92, Kan92, KST94, KS93, KR96b, KT95, LLR94, Lyn90, MR96, MR97, Rau92, Sen98, SZ99, ST96, Sub95, Tho98, UL97, Wei99, Zwi98]. graphs-exact [Zwi98]. greedy [AAF^{+}96]. grid [Sze98, UL97]. grids [Kun91]. group [CFKL94]. groups [BHK^{+}90, BF91, BB93, Bea95, KP90, Luk92, VSR96]. guaranteed [GM97]. guarantees [AFHB^{+}97].

Hajos [PU92]. half [AEM92, Vem97, vD98]. half-space [AEM92]. half-spaces [Vem97]. halfspace [Cha98, Eri96]. halfspaces [BK93, Mat91]. Halvers [AKS92]. Hamiltonian [UL97]. Hard [Imp95, AHP92, BSI99, DKS98, Hås96, Imp90, JL93, KS99, Mic98, Ogi95, PR90, Rab96b]. Hard-core [Imp95, KS99]. Hardness [DMS99, Uma99, ALM^{+}92, ABSS93, Fur95, GIL^{+}90, ISW99]. Hartmanis [CS95]. hash [AAB^{+}92, KST99, BF90]. hashing [GPY94, Sun91]. haste [HR92]. Hausdorff [FCI99]. having [KW91]. heap [BST92]. Heat [ABF93]. heaven [Gol97]. help [BGK^{+}98, NRS94]. helps [Voc99]. heuristic [DZF91]. Heuristics [FK98]. Hidden [dBO90, KKP91]. Hierarchies [GM92, ACF90]. hierarchy [RM97, TW96]. high [AGLR94, FHH^{+}90]. higher [AGR94, Bar95, CIK93, Sha93]. higher-dimensional [AGR94, Bar95]. Highly [AKPR93, LMP91, Spi96, GRS95].

Hilbert [BIK^{+}94]. history [KW90, Ver91]. hitting [ACRT97, MV99]. holds [FFK92]. hole [CCKV97]. holes [HKK91, MMP^{+}91]. homology [DC91]. Horn [AFP90, MP92]. Horn-clause [MP92]. hot [FR92]. hot-potato [FR92]. hull [Cha99, Cha91, CR93, Wag92]. hulls [AGR94]. hypercubes [ABO91]. hypergraph [RS98]. hypothesis [AD93, CSS97].

identification [GKS90]. identities [RS96]. identity [AB99]. ignore [Goo91]. II [KP92]. immune [AN90]. imperfect [MY98]. implicants [Uma98]. implication [MP92]. Improved [AMW95, BIP95, BP92, CG99, Dey97, Fur95, GKV95, GS98, KS97, NS97, PPS92, RS98, Sri97, Vig99, Adl96, Bas97, BP96, CN97, PRT92, PPSZ98]. improving [BC93]. in-approximations [MWW91]. inapproximability [Aro95]. Incoercible [CG96a]. increasing [NGM90]. incremental [Fis97]. independent [AGHP90, DFJ99, FK98, Fle99, GP92]. index [GPY94]. individual [HN98]. inductive [Ful90, Ira90]. inequality [ABNK^{+}92, LS90]. inference [Ful90, KPS92]. Inferring [KW90]. infinitary [GM92]. infinite [HHK95, Sze98]. Information [BH98, ACM98, ACK90, BBR97, CGKS95, EHJ^{+}96, GGL91, GPY94, IR93, KO97, RZ98, vD98]. inputs [CR93]. instances [Imp90]. integer [RV93, SPLK93, SZ99, Yap92]. integers [AB91, CGGK90]. Integral [Bar95, Bar93, KP97b]. Intelligence [Kea98]. interaction [DP92, Kil94]. Interactive [Orl91, BFL90, BGG90, FLS90, LFKN90, Sah99, SYY99, Sha90, Wat99b]. Internet [EHJ^{+}96, KPR98]. interpolation [BPR97, OP98, SS91, Smo90]. interrogation [vD98]. intersection [BK93, KWZ95, Vem97]. intertwines [GI91]. interval [AV96, KST94]. inversion [OP98]. inverting [SZ90]. IP [LPR94]. IP=PSPACE [Sha90]. irrelevant [DH94]. isolation [Nef90]. isomorphism [AT96, FFK92, Gaz91]. isoperimetric [LS90]. Issues [Kea98, Lat95]. iterative [RSW98].

Jacobian [HI91]. JAG [BE93]. JAGs [Poo93]. Jerrum [KR99]. Jerrum-Sinclair [KR99]. Jitter [MPS98]. job [AKRS92]. Juan [IEE91]. jump [Mit96].

key [FGMY97]. kinetic [AEGH98]. Klee [GZ94]. Klee-Minty [GZ94]. knapsack [MS99]. knot [HLP97]. knowledge [BBP91, DP92, FLS90, GP91, Kil94, KPR98, PT96, SV97, Sah99].

labeling [KT99b, SV96]. Languages

[BBFG91, ANKS99, AH92, VVAG92]. large [AKFS99, AKPR93, Far97, FK98, HW96, Mit93]. large-grained [AKPR93]. later [Gol97]. lattice [CN97, LRS95, Mic98, VF90]. lattices [ABSS93, Yap92]. layered [FFK^{+}91]. layout [Vem98]. lazy [Mul91b, Oka95]. leader [RZ98]. lean [IL94]. learn [AHP92, BS94, DKP91]. learnability [ABDCBH93, ADFK97]. learned [CGG98]. Learning [AFP90, BK93, Coh97, Das99, FJK96, GRS95, AP91, AV90, AV99, AD93, BBB^{+}96, Blu90a, BFKV96, BBL98, BC92, Bsh93, BCH94, BM95, DH94, FKM^{+}95, GGR96, HLL92, Jac94, KS90, Li90, MT90, Vem97]. Lemma [PPZ97, Alo91, Alo92, FK96, KT99a]. lemmata [Pap90]. length [ET91]. level [KT99a]. levels [Cha98, Dey97]. light [NS97]. likelihood [Sud96]. limited [MadHVW97]. Limits [KST99]. line [ACM98, AAP93, AGLR94, ABK92, BF90, BKM^{+}91, BB90, BC93, FG95, Ira90, KTDC91, KP94, Mit93, SWW91, SZ90].

Linear [AEL95, Ajt99, AN90, AM90, And96b, ABSS93, Aro97, ABCP93, BFKV96, BMMR97, BST92, Cha90, CHH99, Coh94, CC92, DMS99, DS96, Epp91, FJK96, GKP93, HR91, Kah92, KP92, KS93, KK99, KS96, LRT93, LMSS90, MKS93, SPLK93, Tho97, Tol92, TSSW96, Vai90, Wei99]. linear-processor [KS93]. linear-sized [AN90]. Linearity [BCH^{+}95]. linearizable [HSW91]. linearly [KTI92, MMP^{+}91]. lines [CEG^{+}90]. linguistic [MMS98]. link [HLP97, MPA92]. linking [Oka95]. list [AM97, Amb99]. list-coloring [AM97]. lists [Oka95]. lived [AST99]. lmc [Kan92]. lmc-ordering [Kan92]. Load [AAG^{+}95, Mit96, ABK92, GI99, KRT99, NW94, RSW98, Voc99]. load-balancing [RSW98]. loading [BKN95]. Local [AA94, KN98, RSW98, Alo91, APSV91, AL93, BBR97]. local-control [AL93]. locality [FM97, KP95c, MadHVW97]. localization [Kle94]. location [AEIS99, CG99, CJ90, JV99]. log [RS93, Wag92, Wei94]. logarithmic [BJK^{+}90, Cha99]. logarithms [Sho94]. Logic [Pnu97, AHK97, CC92, Got93, GM92, Kla91, LMSS90, TW96]. Logical [Cos93]. logics [Gro96]. Long [AST99, KPS94]. Long-lived [AST99]. loose [HR92]. Louis [IEE90]. Lovasz [KT99a, Sze94]. Low [HSW91, FKV98, Smo93, ST98]. low-degree [Smo93]. low-rank [FKV98]. Lower [AKRS92, BIK^{+}94, BGP95, BAG91, CS92, Fis97, Gal91, GT98, KPR98, NW95, SS91, Ajt99, Amb99, BE93, BBC^{+}98, BIP95, BP96, BIN97, Bla99, CHLT93, Cha94, DH90, EIRS91, ES93a, Eri96, ES93b, GV94, GKV95, GR98, Gri98, HJP93, KW91, Nay99, PR99, PPS92, RS91, Sha93, Smo90, Sun91, Yao94]. LP [AB91]. Lutz [CSS97].

machine [Sku98]. machines [FST91, Pla95, SWW91]. made [Gol97, Sch91]. Magic [DNRS99]. maintaining [HF93, Kin99]. maintenance [ACK90, AS91, KTDC91, Sch91]. makes [HR92]. Making [RA97]. malleable [Sah99]. management [AV96, KMRV97, MadHVW97]. many [MMP^{+}91]. Map [Tho98]. mapping [KST94]. Marked [AHR98]. Markov [BCF^{+}99, BD97, CGG98, Kan94, KPR92, KT99b, LS90, LRS95, Mit96, RSW98]. Markovian [AV90, KR99]. massive [FKSV99]. match [Gol97]. matching [ABNK^{+}92, AF91, AAL^{+}97, CS98, CL90, CT96, CH92, CCG^{+}93, CGG90, CG96b, DGM90, FG95, GP92, Ind97, Ind98a, KD90, SV96, Var98]. matrices [AB91, yCLZ94, KW91, OP98]. Matrix [Gur90, Abr90, AGMN92, BB93, Bea95, Bla99, Ebe92, Gie92, Lok95, Luk92]. matroids [Kar93]. max [Coh92, KZ97, dlVK98]. MAX-CUT [dlVK98]. maximal [KK99]. Maximizing [Tol92]. Maximum [KA94, Sud96, Wei94, KTI92]. may [Goo91]. Measure [AS94, RSC95]. measures [Kla91]. Median [DZ96, CG99, JV99]. medians [Tod90]. meet [Fra90]. meets [BW94]. Mehlhorn [KC91]. membership [AHP92, yCLZ94, GKV95, Jac94, MT90]. membership-query [Jac94]. memories [BEG^{+}91]. memory [ACF90, AV96, BV99, CCM98, FFM98, GTVV93, KF90, LS91b]. memory-adaptive [BV99]. memory-bounded [CCM98]. men [HP95]. merging [MPT92]. Merlin [MV99]. mesh [BEG90, Rab96a]. meshes [ST96]. message [DDWY90, GMY95]. method [AM96, Rad92, Sch98]. methods [Lok95, LFKN90]. metric [Bar96, CCG^{+}98, Ind99, JV99, KT99b, dlVK98]. metrics [CCG^{+}98, ENRS95, FCI99, KA94]. Mexico [Gol94]. Miami [IEE97]. mice [HP95]. Mick [CR92b]. mildly [JV92]. Milwaukee [IEE95]. min [Var98]. min-cost [Var98]. minimal [VSR96]. minimization [HW97, Uma99]. minimize [MRSG99]. Minimizing [KP97a, ABC^{+}99]. Minimum [Sub95, AEM92, AEGH98, ABNK^{+}92, Cha97, CT96, DZF91, DMS99, FW90, GKP93, Gar96, HH90, KST94, Kar93, KTI92, Rav94, Uma98]. minimum-size [CT96]. minimum-weight [GKP93]. minmax [ET91]. Minty [GZ94]. MIS [BE95]. Missouri [IEE90]. mistake [Blu90a]. mistake-bound [Blu90a]. mixed [ALRS92]. mixing [BCF^{+}99, BD97, LS90, MR96]. mixtures [Das99]. mobile [AKRS92, Kle94]. MOD [GT98, Gro92, GT98]. modal [Got93]. model [AV90, CGG98, GGL91, Koi93, LNNW91, Sun91, YDS95]. models [BE93, Blu90a, Gam98]. modular [CFKL94]. modulo [HW96]. monadic [Cos93, Sch94]. monotone [BGP95, BBL98, CS92, DDPY94, Hak95, HRS93, RM97, Yao94, Bsh93]. monotonicity [GGLR98]. Monte [BCF^{+}99, DKLR95, FKV98]. Monte-Carlo [FKV98]. Morgan [PZ91]. Motion [PRST94, HS93a, Lat95]. moving [KTI92]. MST [PR99]. mu [EJ91]. mu-calculus [EJ91]. Multi [GPY94, ACBFS95, KMO94, LS91a, RV93, Sch95]. multi-armed [ACBFS95]. Multi-index [GPY94]. multi-party [KMO94]. multi-writer [Sch95]. multicommodity [AL93, Fle99, GK98, KARR90]. multidimensional [Mul91b, Mul91a]. multiparty [CG96a]. Multiple [FLS90, PPZ90]. multiplication [Abr90, AGMN92, Bla98, Bla99, PPZ90]. multiplications [And95]. Multiplicative [Sch98, Ge93]. multiplicity [BBB^{+}96]. multipoint [PRT92]. multipole [PRT92]. multithreaded [BL94]. multivariate [BM95, KL96]. multiway [NZ97].

n [Wag92, And96a]. N-body [And96a]. named [Poo93]. natural [Mor91, RSC95]. navigation [BC93]. NC [Lei98, Nef90, RM97, SYY99, SPLK93]. Near [ABCP93, HS93a, ISW99, KK96, Cha99, Kun91, NSS95, PR99]. Near-linear [ABCP93]. near-logarithmic [Cha99]. Near-optimal [ISW99, KK96, NSS95]. Near-quadratic [HS93a]. near-tight [PR99]. nearest [Cal93, FCI99, Ind98b]. Nearly [ADFK97, Aro97, RSW94, AEL95, GMV91, HLL92]. necessary [AMRT96]. needed [KO97]. negated [GTT93]. neighbor [FCI99]. neighborhood [ABCP93]. neighbors [Cal93, Ind98b]. nested [Blo92, HH90]. net [Bou98]. Network [AP90a, AZ98, AU90, AA97, Jai98, LU99, Rad93, FR99]. networks [BB90, Coh92, CMS95, ELW90, GR97b, HRS93, KLOW94, LMS92, LPR94, MPS98, MPT92, NR91, Tam92, Wei94]. Newton [Rad92]. NEXP [LS91a]. NEXP-time [LS91a]. NEXPTIME [CP94]. No [BPR97, Imp90, BE95]. no-MIS [BE95]. node [Poo93]. node-named [Poo93]. node-ordered [Poo93]. noise [AD93]. noisy [BFKV96, Coh97, ES93b, Gal91, GRS95, RS91]. Non [Sah99, SYY99, Ajt99, BGS95, Ede95, FLS90, Ind98b, Tol92]. non-approximability-towards [BGS95]. non-convex [Ede95]. non-Euclidean [Ind98b]. Non-interactive [SYY99, FLS90, Sah99]. non-linear [Ajt99, Tol92]. Non-malleable [Sah99]. nonclairvoyantly [KP97a]. nonconvex [VA97]. Noncryptographic [Fei99]. nondeterminism [RA97]. Nondeterministic [BFL90]. noninteractive [Kil94]. nonregular [HR90]. norm [AAG^{+}95]. normal [Gie92, KL96]. normed [HJM94]. norms [BS90]. note [Sze94]. novel [BMP^{+}99]. November [Gol94, IEE93, IEE98]. NP [AS92, Cos93, DKS98, FGL^{+}91, Got93, GLST98, Hak95, Imp90, Sch94]. NP-complete [FGL^{+}91]. NP-hard [DKS98]. Nullstellensatz [BIK^{+}94, Gri98]. Number [NR97, Adl94, AFB93, ANKS99, CCG^{+}98, FR99, Fle99, Fur95, GK91, PT98, Sze94]. Number-theoretic [NR97]. numbers [Egi93]. numerical [PRT92].

O [Wag92, Wei94]. object [VVAG92]. object-creating [VVAG92]. objectives [AA94]. objects [JCT92, PR96]. Oblivious [CCM98, BR94, FLPR99, ST94]. obstructions [Hav97]. occupancy [KMPS94]. October [IEE90, IEE91, IEE92, IEE95, IEE96, IEE97, IEE99]. odds [CR92b]. offline [adHS96]. offs [PR98]. old [HRG96]. On-line [AGLR94, ABK92, BKM^{+}91, KTDC91, ACM98, AAP93, BB90, BC93, FG95, Ira90, KP94, SWW91]. one [CCG^{+}93, HLL92, KST99, Wat95]. one-dimensional [Wat95]. one-sided [HLL92]. one-way [KST99]. Online [CR90, MRSG99, AB97, FM97, Sho91, Vis90, adHS96]. operations [AW92, Cha99]. optical [MKR95, RTY90]. optima [ABSS93]. Optimal [AV96, Cal93, DG95, FT94, Far97, FG95, FGMY97, HF93, NR91, Nay99, PR98, VK91, Wag92, AEL95, ABO91, ACK90, CG98, Cha91, CG96b, DKLR95, GSV94, ISW99, KK96, NSS95, SV91]. Optimal-resilience [FGMY97]. Optimally [CCG^{+}93, LS92a, NGM90]. optimization [AEM92, AA94, BBR97, Bar95, Gar92, Rad92]. Optimizing [ELL94]. option [CJS96]. oracle [FFK92, MMS98, vD98]. Orchestrating [JKL98]. ordered [Amb99, BST92, Poo93]. ordering [Kan92]. orders [BMMR97]. organization [Gac97]. oriented [LPR94]. other [Aro96, Aro97, GK98]. out-degree [Sen98]. out-of-core [LRT93]. output [Cla94]. output-sensitive [Cla94]. Overcoming [FFM98]. overhead [AP90a]. overload [BKM^{+}91].

P [BF90, Hak95, Ogi95]. P-hard [Ogi95]. PAC [AD93, DH94]. pack [Sho91]. packet [AFHB^{+}97, BFU96]. packing [GK98, KR96a, PST91, Sho91]. paging [ABF93, BGV95, BBK99, FM97, KPR92, LPR94, Tor95]. pair [Cal93]. pairs [DHZ96, KKP91, Kin99, LS92a, SZ99, Zwi98]. pairwise [KT99b]. Palo [IEE93, IEE98]. papers [IEE92, IEE93]. paradigm [AV91, SV96]. Parallel [AM90, AGR94, CIK93, Alo91, AR92, AKPR93, BIN97, BJK^{+}90, BP92, BC92, Cal93, CR92a, CCG^{+}93, FST91, FRU92, Gaz91, GMV91, Goo91, HR92, JM91, KP92, KK96, Lag90, NR95, Rei95, SWW91, Sku98, Spi96, Vai90, Wag92, WK91, dBO90]. parallelized [LS91a]. Parallelizing [BMMR97]. parameterized [KST94, Kos95]. parameters [Epp99]. Parametric [AEGH98]. parentheses [MR97]. parity [PS90]. partial [HF93, KPS92, NW95]. partially [KLM94]. parties [BCJL93]. partitioning [KD90, ST96]. partitions [AGMV96, AP90b]. party [KMO94]. Path [BD97, Rab96a, BST92, CG96b, KKP91, MPA92]. path-matching [CG96b]. paths [AAB^{+}92, AGMN92, Coh93, DHZ96, Epp94, FW90, HS93b, KKP91, Kin99, KS93, KR96b, KT95, SZ99, Sri97, Tho97, VA97, Zwi98]. Pattern [AAL^{+}97, CS98, CCG^{+}93, DGM90, GP92, Ind97]. patterns [SV96]. PCPs [Aro95, BGS95, GLST98]. Pennsylvania [IEE92]. per-packet [AFHB^{+}97]. perceptron [Coh97]. perceptrons [Coh97]. percolation [DZ92]. Perfect [RZ98, AFB93, AAB^{+}92, Var98]. Perfectly [DDWY90]. performance [AGLR94, BC93]. periodic [KLOW94]. permanent [JV92]. permanents [Kog96]. permutation [CG98, KST99]. permutation-based [KST99]. Permuting [FMP90]. persistence [Oka95]. Perspectives [Yan95]. Petri [Bou98]. philosophers [AS90]. phylogeny [AFB93]. physical [KST94]. physics [BCF^{+}99]. picking [Imp90]. piecewise [MKS93]. piecewise-linear [MKS93]. Pittsburgh [IEE92]. planar [Cha99, CHH99, CJ90, Dey97, GSV94, Gaz91, GKP95, GI91, HP99, Kan92, KS93, LRS95, Mit93, MR97, SPLK93, ST96, Wei94, Yap92]. plane [FHH^{+}90, HS93b, MKS93, Var98, Wei99]. planes [BEGJ98]. planning [HS93a, Lat95, PRST94]. play [FKM^{+}95]. point [CS98, CJ90]. points [Bar93, GKS90, HI93, KTI92, KP97b, Mat91, vzGS95]. Pollak [DH90]. polygon [Cha90, HS93a, KK92, MPA92]. polygonal [HS93a]. polygons [HKK91, Wag92]. polyhedra [AS93b, Bar93, Ede95, dBO90]. polyhedron [GKV95, VA97]. polylog [KS93]. polylog-time [KS93]. polylogarithmic [AP90a]. Polynomial [AB91, ABO96, AFW94, Aro96, BS90, AFB93, AS90, Bar93, BSI99, BP92, Bla98, Blo91, BFKV96, BG99, Coh97, CGG98, DKS98, Ge93, Gro96, HP95, HJM94, HW96, Ist90, Kan94, KC91, MMS98, Nef90, PRT92, SS91, Sub95, Tho98]. polynomial-time [AFB93, BFKV96, HJM94]. polynomials [BBC^{+}98, BM95, EKR96, GRS95, GK91, HH90, KP95b, SS97, Smo93]. polytopes [Bar95]. popular [LMS92]. portfolio [Cov96]. positive [KP92]. potato [FR92]. Potential [AM96]. power [AS93a, BS94, HG90, KW97, MSS91, PRT92, Sim94]. powerful [KP95a]. PRAM [ABO91, AR92, JM91]. PRAMs [MSP90]. precision [Nef90]. predictability [BPT90]. prediction [CR93]. prefetching [KK96, VK91]. prefix [CR92a]. preprocessing [CCG^{+}93]. presence [BKM^{+}91]. preserving [BJJ97, GIL^{+}90]. price [vD98]. pricing [CJS96]. Primal [JV99, RV93]. Primal-dual [JV99, RV93]. Primality [AB99]. prime [HW96]. primitive [Bou98]. Priority [ABE^{+}94, CR93]. privacy [FGY93]. Private [CGGK90, CGKS95, KMO94, KO97]. proactive [FGMY97]. Probabilistic [AS92, Bar96, Kea98, DZ92, KS90, KRRT98, MMS98, Sch99]. Probabilistically [ST98]. Probabilities [Lyn90]. probability [LU99]. problem [AFB93, AT96, ABZ96, AZ98, AFK96, ACBFS95, BR92, BKN95, Bou98, yCLZ94, CR98, CG96b, DGG98, ENZ96, FR99, Fri92, Gur90, HS93a, HP95, HW97, HI91, Jai98, KWZ95, Kle94, Kou99, LMN97, MP92, NZ97, NR94, Rad93, Rei93, SV97, Sun91, Uma98, Ung98]. problems [AMW95, AKV96, AHR98, AEIS99, ALM^{+}92, AFK96, Aro96, Aro97, BAG91, BKRS92, CN97, CG99, CR92a, Cha93a, CC92, FMS93, FK96, Gab93, GK98, Gar92, GI99, HLP97, Imp95, IPZ98, Ind98a, JV99, JL93, KP95a, KST94, KT99b, KS97, Lag90, LMSS90, LNNW91, OP98, PST91, Poo93, Rab96b, Rei95, Sch99, SS91, Sri97, Uma99, Vai90]. procedure [AFK96, DS96]. Proceedings [Gol94, IEE90, IEE91, IEE92, IEE93, IEE98]. processes [CS97, HJM94, Mit96, Wat99a]. Processor [KP92, AW92, Goo91, KS93]. Processor-efficient [KP92]. processors [KKL^{+}90]. Product [BCM93]. Products [NRS94, TW96]. Program [BW94, AV91, CG98]. programming [AM90, CC92, Epp91, FT94, KMS94, MR95, SPLK93, TSSW96, Vai90]. programs [Ajt99, AKPR93, BF90, BST98, BGP95, HR90, RV93, ST94, SZ90]. Progress [Kla91]. projection [AV99, Vem98]. projections [CP94]. promise [SV97]. Proof [ALM^{+}92, BEGJ98, BG99, BMP^{+}99, LFKN90, Wat99b]. proof=polynomial [Sha90]. proofs [AS92, BIK^{+}94, BBFG91, BGG90, BPR97, DP92, FLS90, Gri98, Kil94, Mic94, RSC95, ST98]. propagation [ES93b]. properties [HR90, SS97, Smo90]. Property [GGR96]. propositional [BIK^{+}94, LMSS90]. protein [GIP99].

Protocols [AM98, AAF^{+}96, AV91, BFL90, BTY90, BIN97, Fei99, IL94, LS91a, adHS96]. Provably [BEG90, BCJL93]. prover [BFL90, LS91a]. proving [BD97, DH90]. proximity [AEIS99]. pseudo [ISW99, Mul92, NR95, NR97]. pseudo-random [Mul92, NR95, NR97]. pseudo-randomness [ISW99]. Pseudorandom [BCK96, RSC95, ASWZ96]. PSPACE [Pla99, Wat99b]. public [FGMY97]. public-key [FGMY97]. Puerto [IEE91].

QoS [MPS98]. quadrangle [ABNK^{+}92]. Quadratic [RSW92, ET91, HS93a]. quantifier [BPR94, Bas97]. Quantifying [GP91]. Quantum [BBC^{+}98, MY98, Yao93, vD98, ABO96, ASTS^{+}98, AF98, Amb99, BMP^{+}99, BCJL93, BCDZ99, DS96, KW97, Nay99, Sho94, Sho96, Sim94, Wat95, Wat99b, Wat99a]. quartets [JKL98]. queries [AHP92, ANKS99, CG93, GRS95, GLS98, MT90]. query [AMRT96, AD93, GLST98, Jac94, ST98, VVAG92]. queue [CR93]. queues [BST92, ELL94, Gam98]. quorum [NW94].

radicals [Blo91, Blo92, HH90]. radix [AN94]. Ramanujan [Blo92, Kah91]. ramification [Rav94]. RAMs [AMRT96, BAG91]. Random [BSI99, Cha98, Kar93, MS99, Vem98, AM97, AGHP90, ACRT97, AV99, CMS95, FLS90, FK98, Gil93, Imp90, KT99b, Lyn90, MRV99, Mul92, NR95, NR97, Nay99, SZ94, Sub95, Tam92, Vem97, Zuc90, Zuc91]. randomised [AFW94]. randomization [BGK^{+}98, IW98]. Randomized [AMS94, AW92, CS97, GZ94, Koi97, Mul91b, Mul91a, Mul92, Vis90, BKRS92, Fei93, KK94, LS92b, dlVK98]. Randomness [BGG90, BR94, IW98, CG90, ISW99]. Randomness-efficient [BR94]. range [BCM93, Cha98]. rank [Bla99, FKV98, RS93]. ranks [KW91]. Rapid [Rav94, BD97]. rate [LS90]. rates [MR96]. ratio [AKRS92, DH90]. rational [yCLZ94, HI93, OP98]. reachability [Bou98, dAHK98]. reactive [PR90]. Read [AHP92, AP91]. Read-thrice [AHP92]. read-twice [AP91]. real [Bas97, KP95b, Rei93]. rearrangement [FMRW95]. rearrangements [BP93]. receiver [CCM98]. recognizability [Lat95]. Recommendation [KRRT98]. Reconstructing [ALRS92, MS95]. recovery [AEL95]. rectangles [AGMV96, ASWZ96, KK92]. recurrence [Lei98]. recurrences [KS96]. recursive [Bou98]. reduced [YDS95]. Reducibility [KMO94, Cos93]. Reducing [LU99, Vai90]. reduction [RRV99, VF90, Yap92]. Reductions [Aro95]. Reed [GS98, Sud96]. reference [FM97]. Refining [Mit93]. register [AS93a]. registers [Sch95]. Regular [ANKS99, AH92, AEIS99, Kah92, Kun91]. regularity [Alo92, FK96]. related [GI99, Lag90, Sri97]. relationships [GTT93, KT99b]. relative [FFK92]. relaxation [LRT93]. relaxations [Sku98]. release [ABC^{+}99]. Reliable [Gac97, RS91, Gal91, KLM94]. remaindering [AB99]. removal [dBO90]. removing [EC91]. repeated [FKM^{+}95]. repetition [BIN97]. repetitions [KK99]. replicated [ACK90]. Replication [KO97]. Reporting [Mat91, AEM92, Cha98]. representation [BF91, Ben95, MR97]. representations [Ge93, Smo93]. represented [AO91]. requirements [Fra90]. requires [DZ96]. resilience [FGMY97]. resolution [AAF^{+}96, AA94, BP96, BEGJ98, BG99, CS95, FHH^{+}90, GM97, PS95]. Resolving [GMY95]. resource [AA94]. respect [Jac94]. response [AS90]. restricted [BEGJ98]. restriction [PZ91]. result [BW94]. result-checking [BW94]. results [AAF^{+}96, BGS95, Cha91, CJ90, Fur95, Mul91a]. Retrieval [BH98, CGKS95, GPY94, KO97]. reversal [CG98]. reversals [BP93]. revisited [BCK96, Cha93b]. Rico [IEE91]. ride [CR98]. Riemann [HI91]. rigged [ACBFS95]. rigidity [Lok95]. rings [MKR95]. risk [AB97]. RNC [RV93]. robot [Lat95]. robots [BS94, Kle94]. Robust [Ful90, AV99]. robustness [Rub94]. Roch [HI91]. rods [CEG^{+}90]. root [Nef90, Rei93]. roots [NR94]. round [Wat99b]. rounding [AFK96]. rounds [MS95, RZ98]. Routing [CMS95, AFHB^{+}97, AAP93, AGLR94, BB90, BFU96, FR92, FRU92, KRT99, Kun91, MKR95, RSW94, Sri97, Wei99, adHS96]. RSA [HN98]. RSPACE [SZ95]. rumor [Rav94].

S [SZ95, Wei94]. Safe [Cla92]. sample [KK94]. Sampling [KL96, Mat98, ASTS^{+}98, BR94, BCM93, Cha98, Kar93, MS99, Mul91a, Vem97, Vig99]. San [IEE91]. Santa [Gol94]. SAT [LV99, PPSZ98, Sch99]. satisfaction [Sch99]. Satisfiability [Gut98, PPZ97, Pla99, CR92b, KMPS94]. satisfying [Pap91]. Scale [ABDCBH93]. Scale-sensitive [ABDCBH93]. scaling [Gab93]. Scheduling [BL94, SWW91, AKRS92, ABZ96, BKM^{+}91, BKN95, FST91, Hal95, KP95a, MRSG99, Sku98, YDS95]. scheme [BCJL93, GKP95, Ind99, dlVK98]. schemes [ABC^{+}99, AFW94, Aro96, Aro97, Fis97, FK96, RSW98]. Science [Gol94, IEE90, IEE91, IEE92, IEE93, IEE95, IEE96, IEE97, IEE98, IEE99]. Search [LNNW91, BG99, FG95, KN98, Mul91b, Mul91a]. searching [Amb99, And95, And96b, CR90]. second [Kah92]. secretary [AMW95]. secure [DDWY90]. Security [GIL^{+}90, BCK96, BDJR97, HN98, Sah99]. selecting [Pap91]. selection [Cov96, DZ96, Fei99, Pla95]. Self [APSV91, AV91, Gac97, IL94, KS96, Tam92, Wan91]. self-correcting [Wan91]. self-correction [KS96]. self-embedding [Tam92]. self-organization [Gac97]. Self-stabilization [APSV91]. self-stabilizing [AV91, IL94]. self-testing [KS96]. self-testing/self-correction [KS96]. semi [FK98, KP97b, KP90, Sub95]. semi-algebraic [KP97b]. semi-groups [KP90]. semi-random [FK98, Sub95]. Semidefinite [Sku98, KMS94, MR95]. semidirect [TW96]. semigroups [yCLZ94]. Sensitive [Cha93a, ABDCBH93, BCM93, Cla94]. sentences [Lyn90]. separate [BFT95]. separated [Cal93]. Separating [Blu90a, Gro92]. Separation [RM97]. separations [BEGJ98, Ful90]. separator [GSV94, SW98]. separators [MTV91]. sequences [Ist90, KW90]. sequencing [Li90]. sequential [ABCP93]. server [BR92, BKRS92, FRR90, Kou99]. Set [CP94, CHLT93, ENZ96, GTT93, KA94, LMN97, RV93]. set-union [LMN97]. sets [ACRT97, ASWZ96, CS98, Dey97, DFJ99, FK98, KN98, KP97b, KS99, Koi95, MV99, Ogi95]. Setting [Epp99]. shallow [NS97]. shallow-light [NS97]. shared [BGV95, JCT92, PR96]. sharing [NR91]. Shellsort [PPS92]. shifts [Sch98]. shop [Hal95]. Shor [BMP^{+}99]. Short [KR96b, KPS94]. shorter [PPZ90]. shortest [AGMN92, DHZ96, Epp94, FW90, HS93b, KKP91, Kin99, KS93, Mic98, MPA92, SZ99, TY93, Tho97, Uma98, VA97, Zwi98]. show [Hak95]. Shrinkage [PZ91, Hås93]. Shub [Koi93]. shuffling [Mul91b]. side [CR92b]. sided [HLL92]. sigmoid [MSS91]. Signal [ES93b]. signature [Fis97]. similarity [GIP99]. Simple [AGHP90, BM95, AL93, Cha90, LP90, Wag92]. simpler [GK98]. simplex [GZ94]. simplicity [BJJ97]. Simplified [BP96]. Simplifying [HH90]. Simulated [JS93]. Simulating [Zuc91]. simulation [And96a, AR92]. simulations [ABO96, ACRT97, HHK95]. Sinclair [KR99]. Single [Kle96, Bas98, DGG98, FLS90, KO97, Tho97]. Single-source [Kle96, DGG98]. singular [KP92]. sites [MKS93]. Size [BCE91, AS93a, CT96, Coh94, Fis97, Ist90, Lok95, Smo90]. Size-depth [BCE91, Lok95]. sized [AN90]. slowdown [KKR93]. Smale [Koi93]. small [AS94, AMS94, BIP95, BCDZ99, CCG^{+}98, CR92a, Coh92, HG90, Ist90, KK94]. small-depth [HG90]. small-error [BCDZ99]. smallest [Fre91]. smooth [KN98]. snakes [ELW90]. solid [UL97]. Solomon [GS98, Sud96]. solution [DH90, KP92, Rei95]. solutions [MS99]. solvable [CC92, Luk92]. Solving [GTT93, HW96, HH90]. Some [BJK^{+}90, Blu90b, BCF^{+}99, CR92b, Gol97, LMS92, LLR94, Mic98, Smo90]. somewhat [Imp95]. sort [AN94, BAG93, CR93]. sorting [And96b, AU90, BP93, HR92, Kun91, KLOW94, LMP91, PR98]. sorts [LP90]. sound [BIN97]. source [DGG98, Kle96, Tho97, Zuc91]. sources [ACRT97, SZ94, Zuc90]. Space [Poo93, Abr90, AEM92, AGMV96, And96b, BE93, BTY90, BBR^{+}90, BST98, CEG^{+}90, Fei93, Gut98, NSW92, Ogi95, PR98, Sha90, Wat99a]. space-bounded [Wat99a]. space-efficient [Ogi95]. spaces [Bar96, BCM93, Ind98b, Ind99, KK94, Vem97]. span [BGP95]. spanners [AMS94, Coh93]. spanning [AEM92, AEGH98, Cha97, CT96, Fre91, FW90, GKP93, Gar96, Kar93, KTI92, NS97]. Sparse [AP90b, Ogi95, ABCP93, DFJ99, FT94, KP95b, Lyn90, Rei95]. Sparsification [Epp92]. spatial [CS98]. special [Smo90]. Specified [Nef90].

Spectral [Lok95, ST96, Cha94]. spectrum [BS90]. Speed [KP95a]. speeding [Epp92]. spherical [ES93a]. Splitters [NSS95]. spreading [ENRS95]. St [IEE90]. Stability [Gam98, AAF^{+}96]. stabilization [APSV91]. stabilizing [AV91, IL94]. state [CGG98, KW97]. states [AFB93]. Static [AMRT96, ELL94, MR97]. statistical [AD93, BG92, BCF^{+}99, SV97]. stealing [BL94]. Steiner [FR99, DH90, DZF91, Jai98]. Stochastic [GI99]. Storage [KMRV97, FMRW95]. straight [BF90, Mit93, SZ90]. straight-line [Mit93]. stratified [GM92]. streams [FKSV99, Kun91]. street [Kle91]. strengths [AF98]. stretch [Coh93, MRSG99]. string [CL90, CH92, CGG90, FLS90, FG95, Ind98a, Li90]. strings [MS95]. strip [KR96a]. strongly [IPZ98]. structural [BST92]. structure [BAG91, GIP99]. structured [OP98]. structures [AEIS99, And96a, AB96, Fre91, LRS95, Sch91]. study [BG99]. sub [GKP93]. sub-linear [GKP93]. subexponential [Gar92]. subgraph [KP93]. subgraphs [CT96]. sublinear [CL90, FG95, Ind99]. Sublogarithmic [And95, KKR93]. submodular [Gab93]. Subquadratic [BBP91]. subring [AB91]. subset [ENZ96, GTT93]. substitution [Sch98]. substrings [MS95]. subtree [KA94]. Succinct [MR97]. sufficient [AMRT96]. suffix [Far97, FFM98, Kos95]. sums [Blo91, HF93]. superconcentrators [RTS97]. superfast [OP98]. superimposed [Ind97]. superstrings [KPS94, TY93]. surely [AM90]. surface [dBO90]. surfaces [DG95, KT99a]. swaps [AAL^{+}97]. switch [Wei99]. switching [AKS92]. symmetric [BDJR97, KW91, PPZ90, Smo90]. Symposium [Gol94, IEE90, IEE91, IEE92, IEE93, IEE95, IEE96, IEE97, IEE98, IEE99]. Synchronization [AS93a, AP90a]. synchronous [MSP90]. syntactic [KMSV94]. synthesize [PR90]. Synthesizers [NR95]. system [AS91, CDZ98]. systems [ABSS93, AR92, BEGJ98, FGY93, GTT93, HW96, KP92, KRRT98, LFKN90, MadHVW97, NW94, PR90, RSW92, Ver91, Wat99b, Woo98]. SZK [DDPY94].

t [Wei94]. Tail [KMPS94]. Taking [HP99, AB97]. tangential [OP98]. tasting [HLL92]. tautologies [Gri98]. Taylor [Sch98]. TC [BPR97]. TDI [CDZ98]. team [BS94]. technique [BD97, Epp92, KD90]. techniques [Adl96, HRG96, PRT92]. Temporal [Pnu97, TW96, AHK97, Kla91]. term [KD90]. terminals [FR99]. test [BW94]. testable [ANKS99]. Testing [Ge93, GGLR98, AB99, AKFS99, BCH^{+}95, BW94, GGR96, GKV95]. testing/self [KS96]. tests [BG92]. their [BBFG91, NR95]. theorem [BKRS92, HKK91]. theorems [SW98]. theoretic [FGY93, NR97, Pap90]. Theoretical [Kea98, BV99]. theory [Adl94, AADW94, AH92, AV99, BF91, BG92, BW94, Bsh93, Egi93, FRU92, GMV91, Gol97, Li90, RSC95, Ver91, Yan95]. theory-the [Adl94]. three [AS93b, Cha98, Epp91, KT99a]. three-dimensional [Epp91, KT99a]. threshold [BFKV96, BS90, HG90, HRS93, KMPS94, MSS91, PS90, Rad91, ST94, Yao90]. thrice [AHP92]. Throughput [AAP93]. Throughput-competitive [AAP93]. Tight [KP95c, MP90, Pla95, RTS97, ADFK97, BGS95, CHLT93, GLST98, HR92, KKL^{+}90, PR99, RSW94, Sha93, Sho91]. Tighter [CH92]. tile [Sze98]. Tiling [KK92]. Time [BE93, BBR^{+}90, BST98, Abr90, ABC^{+}99, AFB93, Ajt99, AM90, AFW94, AEL95, AHK97, AMRT96, Aro96, Aro97, AS90, BFL90, Bar93, BE95, BAG93, BKN95, Blo91, BFKV96, Cha99, CL90, Cha90, CHH99, Coh94, Coh97, CGG98, ET91, Fei93, FG95, FMRW95, GKP93, Ge93, GMV91, Gro96, HJM94, HR91, IW98, Ind99, JM91, KP97a, Kan94, KC91, KKP91, KS93, KK99, LS91a, MMS98, PR98, PPSZ98, PR99, Rav94, Sub95, Tho97, Tho98, Wag92, Wei94, Wei99]. Time-space [BE93, BBR^{+}90, BST98, Abr90, Fei93, PR98]. timed [AH92]. times [Ben95]. Tits [Bea95]. tolerance [LMS92]. tolerant [AAB^{+}92, AU90, AB96, BMP^{+}99, GGL91, JCT92, LMP91, PR96, Sho96, Spi96]. tools [Blu90b]. Top [HJP93]. Top-down [HJP93]. topological [Bas98]. Torpid [BCF^{+}99]. tours [KPS94]. tracing [RTY90]. Tracking [AW95, AKRS92]. Tractability [KST94]. tractable [FR99]. trade [PR98]. trade-offs [PR98]. tradeoff [Abr90, Fei93]. tradeoffs [BTY90, BBR^{+}90, BST98, BCE91, CG90, Lok95]. trading [AB97]. trails [CHH99]. Trans [FW90]. Trans-dichotomous [FW90]. transcendental [GV94]. transfer [CCM98]. transformations [FJK96]. Transforming [HP95]. transition [Wat99a]. transitive [Coh94, GM92, HK95, Kin99]. translator [BR92]. transmission [ABE^{+}94, DDWY90]. traversal [BBR^{+}90, FFK^{+}91]. traversing [Ist90]. treatment [BDJR97]. Tree [And96a, EJ91, CCG^{+}98, DGM90, FT94, Far97, FFM98, Gar96, KD90, Lag90, Lei98, LNNW91]. tree-decomposition [Lag90]. tree-partitioning [KD90]. trees [AEM92, AEGH98, BO94, BM95, Cha97, CGG98, DZF91, Fre91, FW90, GKP93, Got93, GV94, GKV95, GNRS99, HF93, Kar93, KTI92, KA94, Kos95, MKR95, Mul91b, Mul91a, MR97, NS97, NRS94, RS91, Yao92]. trees-metrics [KA94]. trellises [VSR96]. triangles [MMP^{+}91]. Triangulating [Cha90]. triangulation [DC91, ET91, Mit93]. triconnected [Hsu92]. triconnectivity [HR91]. triply [BJK^{+}90]. triply-logarithmic [BJK^{+}90].

Truly [FM97, GP92]. truncated [MS99]. truth [Pap91]. Tseitin [Gri98]. TSP [Aro96, Aro97, GKP95]. turning [adHS96]. Tutorial [Kea98, Mat98]. Tutte [AFW94]. twenty [Gol97]. twice [AP91, GSV94, SV91]. twist [Sch98]. Two [Pnu97, BFL90, BCH^{+}95, BS94, CCG^{+}93, CGG98, GP92, KW91, RTS97, RS98, Wan91]. two-coloring [RS98]. two-dimensional [GP92, Wan91]. two-prover [BFL90]. two-state [CGG98]. type [CIK93, DC91]. types [KPS92].

unambiguous [RA97]. unbounded [KW91]. unbreakable [BCJL93]. Undecidability [MP92]. Undirected [NSW92, Tho97, BBR^{+}90, GR97b, HRS93, SZ99]. unification [KP90]. unified [MTV91, OP98, Tor95]. Uniform [ACF90, ABDCBH93, BK93, IW98, Jac94]. uniformly [Imp90]. unimodular [Yap92]. union [AS93b, LMN97]. uniquely [AO91]. unit [GR97b]. unitary [DS96]. Universal [AAF^{+}96, Cov96, KKR93, BB90, BMP^{+}99, Ist90]. universality [BMP^{+}99]. unknown [DP90, DKP91, Kle91]. unlabeled [BS94]. unreliable [KLM94]. unrestricted [BTY90]. Unsatisfiable [Woo98]. unsplittable [DGG98, Kle96, KS97, Sri97]. until [TW96]. update [FG95]. upper [Sha93]. user [AKRS92]. Using [BFT95, CR93, Goo91, ABNK^{+}92, BBR97, Cal93, GKS90, Kan92, LRT93, MV99, SV96, Ver91, Zuc91]. USTCON [Fei93]. usually [LP90]. utilization [Adl96].

V [Wei94]. Valiant [AV90]. value [IR93]. variable [Gro96]. variables [AGHP90]. Variation [KW91]. varieties [Koi97]. VC [MWW91]. VC-dimension [MWW91]. vector [KWZ95, Mic98]. Vegas [BB93]. Verifiable [MRV99]. verification [ALM^{+}92]. Verifying [RS96]. Vermont [IEE96]. version [Alo91, Koi93]. versus [KMSV94, MSS91]. vertex [ENZ96, HRG96]. vertices [Gar96]. very [Lyn90, SZ94]. via [AB99, AD93, Bsh93, CT96, ENRS95, FCI99, FMRW95, Gam98, NW95, Oka95, VK91]. views [KMSV94]. VLSI [CHH99, Vem98]. volume [Koi95, LS90, LS92b]. Voronoi [MKS93]. vs [HRS93, IW98, KR99].

wait [ALS90, Hav97, JCT92]. wait-free [ALS90, Hav97, JCT92]. walk [HP99]. Walking [Kle91]. walks [Gil93, MS99]. Waste [HR92]. wavelength [MKR95]. way [AF98, KST99]. ways [Imp90]. Weak [ACRT97, Kou99, Koi93, Smo90, SZ94, Zuc90, Zuc91]. weaknesses [AF98]. Web [BH98]. weighing [AKV96]. weight [GKP93]. weighted [ABC^{+}99, HF93, Zwi98]. weights [SZ99]. well [Cal93]. well-separated [Cal93]. where [Kog96]. Which [IPZ98, PT98]. winners [AV94]. Wisconsin [IEE95]. wise [AGHP90]. within [Ben95, DKS98, GSV94, Hås96, Mic98, SV91]. without [And95, DP92, SS97]. Witnesses [AGMN92]. word [FMS93, Gut98, KK99, Pla99]. work [BL94]. works [ST96]. wormhole [FRU92, RSW94]. worst [CN97]. worst-case [CN97]. writer [Sch95].

years [Gol97]. yield [Ogi95]. York [IEE99].

Zero [DP92, BBP91, BCDZ99, FLS90, Kil94, KPR98, SV97, Sah99]. zero-error [BCDZ99]. Zero-knowledge [DP92, BBP91, Kil94, SV97]. zeros [GK91].

References

Awerbuch:1994:LOG

[AA94] B. Awerbuch and Y. Azar. Local optimization of global objectives: competitive distributed deadlock resolution and resource allocation. In Goldwasser [Gol94], pages 240– 249. CODEN ASFPDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Awerbuch:1997:BBN

[AA97] B. Awerbuch and Y. Azar.

Buy-at-bulk network design. In IEEE [IEE97], pages 542– 547. CODEN ASFPDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Ajtai:1992:FTG

[AAB^{+}92] M. Ajtai, N. Alon, J. Bruck, R. Cypher, C. T. Ho, M. Naor, and E. Szemeredi. Fault tolerant graphs, perfect hash functions and disjoint paths. In IEEE [IEE92], pages 693–702. CODEN ASFPDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Ajtai:1994:TCA

[AADW94] M. Ajtai, J. Aspnes, C. Dwork, and O. Waarts. A theory of competitive analysis for distributed algorithms. In Goldwasser [Gol94], pages 401–411. CODEN ASFPDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Andrews:1996:USR

[AAF^{+}96] M. Andrews, B. Awerbuch, A. Fernandez, J. Kleinberg, T. Leighton, and Zhiyong Liu. Universal stability results for greedy contention-resolution protocols. In IEEE [IEE96], pages 380–389. CODEN ASF-PDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Awerbuch:1995:LBN

[AAG^{+}95] B. Awerbuch, Y. Azar, E. F. Grove, Ming-Yang Kao, P. Krishnan, and J. S. Vitter. Load balancing in the L_{p} norm. In IEEE [IEE95], pages 383–391. CODEN ASF-PDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Amir:1997:PMS

[AAL^{+}97] A. Amir, Y. Aumann, G. M.

Landau, M. Lewenstein, and N. Lewenstein. Pattern matching with swaps. In IEEE [IEE97], pages 144–153. CO-DEN ASFPDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Awerbuch:1993:TCL

[AAP93] B. Awerbuch, Y. Azar, and S. Plotkin. Throughputcompetitive on-line routing. In IEEE [IEE93], pages 32–40. CODEN ASFPDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Adler:1991:PAL

[AB91] I. Adler and P. A. Beling. Polynomial algorithms for LP over a subring of the algebraic integers with applications to LP with circulant matrices. In IEEE [IEE91], pages 480–487. CODEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no.

Aumann:1996:FTD

[AB96] Y. Aumann and M. A. Bender. Fault tolerant data structures. In IEEE [IEE96], pages 580– 589. CODEN ASFPDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Al-Binali:1997:CAR

[AB97] S. Al-Binali. The competitive analysis of risk taking with applications to online trading. In IEEE [IEE97], pages 336– 344. CODEN ASFPDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Agrawal:1999:PIT

[AB99] M. Agrawal and S. Biswas. Primality and identity testing via Chinese remaindering. In IEEE [IEE99], pages 202– 208. CODEN ASFPDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Afrati:1999:ASM

[ABC^{+}99] F. Afrati, E. Bampis, C. Chekuri, D. Karger, C. Kenyon, S. Khanna, I. Milis, M. Queyranne, M. Skutella, C. Stein, and M. Sviridenko. Approximation schemes for minimizing average weighted completion time with release dates. In IEEE [IEE99], pages 32–43. CODEN ASFPDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Awerbuch:1993:NLC

[ABCP93] B. Awerbuch, B. Berger, L. Cowen, and D. Peleg. Near-linear cost sequential and distributed constructions of sparse neighborhood covers. In IEEE [IEE93], pages 638–647. CODEN ASFPDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Alon:1993:SSD

[ABDCBH93] N. Alon, S. Ben-David, N. Cesa-Bianchi, and D. Haussler. Scale-sensitive dimensions, uniform convergence, and learnability. In IEEE [IEE93], pages 292–301. CODEN ASF-PDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Albanese:1994:PET

[ABE^{+}94] A. Albanese, J. Blomer, J. Edmonds, M. Luby, and M. Sudan. Priority encoding transmission. In Goldwasser [Gol94], pages 604–612. CO-DEN ASFPDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Awerbuch:1993:HDC

[ABF93] B. Awerbuch, Y. Bartal, and A. Fiat. Heat and Dump: competitive distributed paging. In IEEE [IEE93], pages 22–31. CODEN ASFPDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Azar:1992:LLB

[ABK92] Y. Azar, A. Z. Broder, and A. R. Karlin. On-line load balancing. In IEEE [IEE92], pages 218–225. CODEN ASF-PDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0.

IEEE Computer Society Press Order Number 2900.

Aggarwal:1992:EMC

[ABNK^{+}92] A. Aggarwal, A. Bar-Noy, S. Khuller, D. Kravets, and B. Schieber. Efficient minimum cost matching using quadrangle inequality. In IEEE [IEE92], pages 583–592. CO-DEN ASFPDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Anumann:1991:AOP

[ABO91] Y. Anumann and M. Ben-Or. Asymptotically optimal PRAM emulation on faulty hypercubes. In IEEE [IEE91], pages 440–446. CODEN ASF-PDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Aharonov:1996:PSD

[ABO96] D. Aharonov and M. Ben-Or.

Polynomial simulations of decohered quantum computers. In IEEE [IEE96], pages 46– 55. CODEN ASFPDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Abrahamson:1990:TST

[Abr90] K. Abrahamson. A timespace tradeoff for Boolean matrix multiplication. In IEEE [IEE90], pages 412–419. CO-DEN ASFPDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Arora:1993:HAO

[ABSS93] S. Arora, L. Babai, J. Stern, and Z. Sweedyk. The hardness of approximate optima in lattices, codes, and systems of linear equations. In IEEE [IEE93], pages 724–733. CODEN ASFPDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Andrews:1996:NAD

[ABZ96] M. Andrews, M. A. Bender, and L. Zhang. New algorithms for the disk scheduling problem. In IEEE [IEE96], pages 550–559. CODEN ASF-PDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Auer:1995:GRC

[ACBFS95] P. Auer, N. Cesa-Bianchi, Y. Freund, and R. E. Schapire. Gambling in a rigged casino: The adversarial multi-armed bandit problem. In IEEE [IEE95], pages 322–331. CO-DEN ASFPDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Alpern:1990:UMH

[ACF90] B. Alpern, L. Carter, and E. Feig. Uniform memory hierarchies. In IEEE [IEE90], pages 600–608. CODEN ASF-PDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Awerbuch:1990:COM

[ACK90] B. Awerbuch, I. Cidon, and S. Kutten. Communicationoptimal maintenance of replicated information. In IEEE [IEE90], pages 492–502. CO-DEN ASFPDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Albers:1998:DIA

[ACM98] S. Albers, M. Charikar, and M. Mitzenmacher. Delayed information and action in online algorithms. In IEEE [IEE98], pages 71–80. CODEN ASFPDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Andreev:1997:WRS

[ACRT97] A. E. Andreev, A. E. F. Clementi, J. D. P. Rolim, and L. Trevisan. Weak random sources, hitting sets, and BPP simulations. In IEEE [IEE97], pages 264–272. CO-DEN ASFPDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Aslam:1993:GBS

[AD93] J. A. Aslam and S. E. Decatur.

General bounds on statistical query learning and PAC learning with noise via hypothesis boosting. In IEEE [IEE93], pages 282–291. CODEN ASF-PDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Ambainis:1997:NTB

[ADFK97] A. Ambainis, R. Desper, M. Farach, and S. Kannan. Nearly tight bounds on the learnability of evolution. In IEEE [IEE97], pages 524–533. CODEN ASFPDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

aufderHeide:1996:DRB

[adHS96] F. Meyer auf der Heide and C. Scheideler. Deterministic routing with bounded buffers: turning offline into online protocols. In IEEE [IEE96], pages 370–379. CODEN ASF-PDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Adleman:1994:ANT

[Adl94] L. M. Adleman. Algorithmic number theory-the complexity contribution. In Goldwasser [Gol94], pages 88–113. CODEN ASFPDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Adler:1996:NCT

[Adl96] M. Adler. New coding techniques for improved bandwidth utilization. In IEEE [IEE96], pages 173–182. CODEN ASF-PDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Agarwal:1998:PKM

[AEGH98] P. K. Agarwal, D. Eppstein, L. J. Guibas, and M. R. Henzinger. Parametric and kinetic minimum spanning trees. In IEEE [IEE98], pages 596– 605. CODEN ASFPDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Amir:1999:ERD

[AEIS99] A. Amir, A. Efrat, P. Indyk, and H. Samet. Efficient regular data structures and algorithms for location and proximity problems. In IEEE [IEE99], pages 160–170. CODEN ASFPDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Alon:1995:LTE

[AEL95] N. Alon, J. Edmonds, and M. Luby. Linear time erasure codes with nearly optimal recovery. In IEEE [IEE95], pages 512–519. CODEN ASF-PDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Agarwal:1992:DHS

[AEM92] P. K. Agarwal, D. Eppstein, and J. Matousek. Dynamic half-space reporting, geometric optimization, and minimum spanning trees. In IEEE [IEE92], pages 80–89. CODEN ASFPDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428.

LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Amir:1991:ADM

[AF91] A. Amir and M. Farach. Adaptive dictionary matching. In IEEE [IEE91], pages 760–766. CODEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Ambainis:1998:WQF

[AF98] A. Ambainis and R. Freivalds. 1-way quantum finite automata: strengths, weaknesses and generalizations. In IEEE [IEE98], pages 332–341. CO-DEN ASFPDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Agarwala:1993:PTA

[AFB93] R. Agarwala and D. Fernandez-Baca. A polynomial-time algorithm for the perfect phylogeny problem when the number of character states is fixed. In IEEE [IEE93], pages 140– 147. CODEN ASFPDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Andrews:1997:GDR

[AFHB^{+}97] M. Andrews, A. Fernandez, M. Harchol-Balter, T. Leighton, and L. Zhang. General dynamic routing with perpacket delay guarantees of O(distance + 1/sessionrate). In IEEE [IEE97], pages 294–302. CODEN ASFPDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Arora:1996:NRP

[AFK96] S. Arora, A. Frieze, and H. Kaplan. A new rounding procedure for the assignment problem with applications to dense graph arrangement problems. In IEEE [IEE96], pages 21– 30. CODEN ASFPDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Angluin:1990:LCH

[AFP90] D. Angluin, M. Frazier, and L. Pitt. Learning conjunctions of Horn clauses. In IEEE [IEE90], pages 186–192.

CODEN ASFPDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Alon:1994:PTR

[AFW94] N. Alon, A. Frieze, and D. Welsh. Polynomial time randomised approximation schemes for the Tutte polynomial of dense graphs. In Goldwasser [Gol94], pages 24–35. CODEN ASFPDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Alon:1990:SCA

[AGHP90] N. Alon, O. Goldreich, J. Håstad, and R. Peralta. Simple construction of almost k-wise independent random variables. In IEEE [IEE90], pages 544– 553. CODEN ASFPDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Awerbuch:1994:LAC

[AGLR94] B. Awerbuch, R. Gawlick, T. Leighton, and Y. Rabani. On-line admission control and circuit routing for high performance computing and communication. In Goldwasser [Gol94], pages 412–423. CO-DEN ASFPDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Alon:1992:WBM

[AGMN92] N. Alon, Z. Galil, O. Margalit, and M. Naor. Witnesses for Boolean matrix multiplication and for shortest paths. In IEEE [IEE92], pages 417– 426. CODEN ASFPDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Agarwal:1996:BSP

[AGMV96] P. K. Agarwal, E. F. Grove, T. M. Murali, and J. S. Vitter. Binary space partitions for fat rectangles. In IEEE [IEE96], pages 482–491. CO-DEN ASFPDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Amato:1994:PAH

[AGR94] N. M. Amato, M. T. Goodrich, and E. A. Ramos. Parallel algorithms for higher-dimensional convex hulls. In Goldwasser [Gol94], pages 683–694. CO-DEN ASFPDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Alur:1992:BFT

[AH92] R. Alur and T. A. Henzinger. Back to the future: towards a theory of timed regular languages. In IEEE [IEE92], pages 177–186. CODEN ASF-PDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Alur:1997:ATT

[AHK97] R. Alur, T. A. Henzinger, and O. Kupferman. Alternatingtime temporal logic. In IEEE [IEE97], pages 100–109. CO-DEN ASFPDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Aizenstein:1992:RTD

[AHP92] H. Aizenstein, L. Hellerstein, and L. Pitt. Read-thrice DNF is hard to learn with membership and equivalence queries. In IEEE [IEE92], pages 523– 532. CODEN ASFPDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Alstrup:1998:MAP

[AHR98] S. Alstrup, T. Husfeldt, and T. Rauhe. Marked ancestor problems. In IEEE [IEE98], pages 534–543. CODEN ASF-PDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Ajtai:1999:NLT

[Ajt99]  M. Ajtai.  A non-linear

time lower bound for Boolean branching programs. In IEEE [IEE99], pages 60–70. CODEN ASFPDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Alon:1999:ETL

[AKFS99] N. Alon, M. Krivelevich, E. Fischer, and M. Szegedy. Efficient testing of large graphs. In IEEE [IEE99], pages 656– 666. CODEN ASFPDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Aumann:1993:HEA

[AKPR93] Y. Aumann, Z. M. Kedum, K. V. Palem, and M. O. Rabin. Highly efficient asynchronous execution of largegrained parallel programs. In IEEE [IEE93], pages 271–280. CODEN ASFPDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Alon:1992:LBC

[AKRS92] N. Alon, G. Kalai, M. Ricklin, and L. Stockmeyer. Lower bounds on the competitive ratio for mobile user tracking and distributed job scheduling. In IEEE [IEE92], pages 334– 343. CODEN ASFPDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Ajtai:1992:HES

[AKS92] M. Ajtai, J. Komlos, and E. Szemeredi. Halvers and expanders (switching). In IEEE [IEE92], pages 686–692. CO-DEN ASFPDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Alon:1996:GCW

[AKV96] N. Alon, D. N. Kozlov, and V. H. Vu. The geometry of coin-weighing problems. In IEEE [IEE96], pages 524–532. CODEN ASFPDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Awerbuch:1993:SLC

[AL93] B. Awerbuch and T. Leighton.

A simple local-control approximation algorithm for multicommodity flow. In IEEE [IEE93], pages 459–468. CO-DEN ASFPDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Arora:1992:PVH

[ALM^{+}92] S. Arora, C. Lund, R. Motwani, M. Sudan, and Mario Szegedy. Proof verification and hardness of approximation problems. In IEEE [IEE92], pages 14–23. CODEN ASF-PDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Alon:1991:PAV

[Alo91] N. Alon. A parallel algorithmic version of the local lemma. In IEEE [IEE91], pages 586–593. CODEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Alon:1992:AAR

[Alo92] N. Alon. The algorithmic aspects of the regularity lemma. In IEEE [IEE92], pages 473– 481. CODEN ASFPDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Ar:1992:RAF

[ALRS92] S. Ar, R. J. Lipton, R. Rubinfeld, and M. Sudan. Reconstructing algebraic functions from mixed data. In IEEE [IEE92], pages 503–512. CO-DEN ASFPDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Attiya:1990:WFA

[ALS90] H. Attiya, N. Lynch, and N. Shavit. Are wait-free algorithms fast? In IEEE [IEE90], pages 55–64. CODEN ASFPDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Alon:1990:PLP

[AM90] N. Alon and N. Megiddo. Parallel linear programming in fixed dimension almost surely in constant time. In IEEE [IEE90], pages 574–582. CO-DEN ASFPDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Amano:1996:PAM

[AM96] K. Amano and A. Maruoka. Potential of the approximation method. In IEEE [IEE96], pages 431–440. CODEN ASF-PDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Achlioptas:1997:ALC

[AM97] D. Achlioptas and M. Molloy. The analysis of a list-coloring algorithm on a random graph. In IEEE [IEE97], pages 204– 212. CODEN ASFPDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Adler:1998:PAC

[AM98] M. Adler and B. M. Maggs. Protocols for asymmetric communication channels. In IEEE [IEE98], pages 522–533. CO-DEN ASFPDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Ambainis:1999:BLB

[Amb99] A. Ambainis. A better lower bound for quantum algorithms searching an ordered list. In IEEE [IEE99], pages 352–357. CODEN ASFPDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Andersson:1996:SDA

[AMRT96] A. Andersson, P. B. Miltersen, S. Riis, and M. Thorup. Static dictionaries on AC^{0} RAMs: query time θ(\sqrt{log n log log n}) is necessary and sufficient. In IEEE [IEE96], pages 441–450. CODEN ASFPDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Arya:1994:RDA

[AMS94] S. Arya, D. M. Mount, and M. Smid. Randomized and deterministic algorithms for geometric spanners of small diameter. In Goldwasser [Gol94], pages 703–712. CODEN ASF-PDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Ajtai:1995:IAA

[AMW95] M. Ajtai, N. Megiddo, and O. Waarts. Improved algorithms and analysis for secretary problems and generalizations. In IEEE [IEE95], pages 473–482. CODEN ASF-PDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Alon:1990:CFG

[AN90] N. Alon and M. Naor. Coinflipping games immune against linear-sized coalitions. In IEEE [IEE90], pages 46–54. CODEN ASFPDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Andersson:1994:NER

[AN94] A. Andersson and S. Nilsson. A new efficient radix sort. In Goldwasser [Gol94], pages 714– 721. CODEN ASFPDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Andersson:1995:SSM

[And95] A. Andersson. Sublogarithmic searching without multiplications. In IEEE [IEE95], pages 655–663. CODEN ASF-PDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Anderson:1996:TDS

[And96a] R. J. Anderson. Tree data structures for N-body simulation. In IEEE [IEE96], pages 224–233. CODEN ASF-PDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Andersson:1996:FDS

[And96b] A. Andersson. Faster deterministic sorting and searching in linear space. In IEEE [IEE96], pages 135–141. CO-DEN ASFPDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Alon:1999:RLT

[ANKS99] N. Alon, I. Newman, M. Krivelevich, and M. Szegedy. Regular languages are testable with a constant number of queries. In IEEE [IEE99], pages 645– 655. CODEN ASFPDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Andersson:1991:FUR

[AO91] A. Andersson and T. Ottmann. Faster uniquely represented dictionaries. In IEEE [IEE91], pages 642–649. CODEN ASF-PDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Awerbuch:1990:NSP

[AP90a] B. Awerbuch and D. Peleg. Network synchronization with polylogarithmic overhead. In IEEE [IEE90], pages 514–522. CODEN ASFPDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Awerbuch:1990:SP

[AP90b] B. Awerbuch and D. Peleg.

Sparse partitions. In IEEE [IEE90], pages 503–513. CO-DEN ASFPDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Aizenstein:1991:ELR

[AP91] H. Aizenstein and L. Pitt. Exact learning of read-twice DNF formulas. In IEEE [IEE91], pages 170–179. CODEN ASF-PDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Awerbuch:1991:SSL

[APSV91] B. Awerbuch, B. Patt-Shamir, and G. Varghese. Selfstabilization by local checking and correction. In IEEE [IEE91], pages 268–277. CO-DEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Aumann:1992:CCF

[AR92] Y. Aumann and M. O. Rabin. Clock construction in fully asynchronous parallel systems and PRAM simulation. In IEEE [IEE92], pages 147– 156. CODEN ASFPDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Arora:1995:RCP

[Aro95] S. Arora. Reductions, codes, PCPs, and inapproximability. In IEEE [IEE95], pages 404–413. CODEN ASF-PDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Arora:1996:PTA

[Aro96] S. Arora. Polynomial time approximation schemes for Euclidean TSP and other geometric problems. In IEEE [IEE96], pages 2–11. CODEN ASFPDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Arora:1997:NLT

[Aro97] S. Arora. Nearly linear time approximation schemes for Euclidean TSP and other geometric problems. In IEEE [IEE97], pages 554–563. CO-DEN ASFPDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Awerbuch:1990:DPA

[AS90] B. Awerbuch and M. Saks. A dining philosophers algorithm with polynomial response time. In IEEE [IEE90], pages 65– 74. CODEN ASFPDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Awerbuch:1991:MCD

[AS91] B. Awerbuch and L. J. Schulman. The maintenance of common data in a distributed system. In IEEE [IEE91], pages 505–514. CODEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Arora:1992:PCP

[AS92] S. Arora and S. Safra. Probabilistic checking of proofs; a new characterization of NP. In IEEE [IEE92], pages 2– 13. CODEN ASFPDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Aiek:1993:SPD

[AS93a] Y. Aiek and G. Stupp. Synchronization power depends on the register size. In IEEE [IEE93], pages 196–205. CO-DEN ASFPDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Aronov:1993:UCP

[AS93b] B. Aronov and M. Sharir. The union of convex polyhedra in three dimensions. In IEEE [IEE93], pages 518–527. CO-DEN ASFPDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Allender:1994:MSC

[AS94] E. Allender and M. Strauss. Measure on small complexity classes, with applications for BPP. In Goldwasser [Gol94], pages 807–818. CODEN ASF-PDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Afek:1999:LLA

[AST99] Y. Afek, G. Stupp, and D. Touitou. Long-lived adaptive collect with applications. In IEEE [IEE99], pages 262– 272. CODEN ASFPDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Ambainis:1998:QCC

[ASTS^{+}98] A. Ambainis, L. J. Schulman, A. Ta-Shma, U. Vazirani, and A. Wigderson. The quantum communication complexity of sampling. In IEEE [IEE98], pages 342–351. CODEN ASF-PDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Armoni:1996:DSP

[ASWZ96] R. Armoni, M. Saks, A. Wigderson, and Shiyu Zhou. Discrepancy sets and pseudorandom generators for combinatorial rectangles. In IEEE [IEE96], pages 412–421. CO-DEN ASFPDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Agrawal:1996:BIP

[AT96] M. Agrawal and T. Thierauf. The Boolean isomorphism problem. In IEEE [IEE96], pages 422–430. CO-DEN ASFPDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Assaf:1990:FTS

[AU90] S. Assaf and E. Upfal. Fault tolerant sorting network. In IEEE [IEE90], pages 275–284. CODEN ASFPDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Aldous:1990:MEV

[AV90] D. Aldous and U. Vazirani. A Markovian extension of Valiant’s learning model. In IEEE [IEE90], pages 392–396. CODEN ASFPDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Awerbuch:1991:DPC

[AV91] B. Awerbuch and G. Varghese.

Distributed program checking: a paradigm for building selfstabilizing distributed protocols. In IEEE [IEE91], pages 258–267. CODEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Aldous:1994:SWA

[AV94] D. Aldous and U. Vazirani.

“go with the winners” algorithms. In Goldwasser [Gol94], pages 492–501. CODEN ASF-PDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Arge:1996:ODI

[AV96] L. Arge and J. S. Vitter. Optimal dynamic interval management in external memory. In IEEE [IEE96], pages 560– 569. CODEN ASFPDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Arriaga:1999:ATL

[AV99] R. I. Arriaga and S. Vempala. An algorithmic theory of learning: robust concepts and random projection. In IEEE [IEE99], pages 616– 623. CODEN ASFPDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Aspnes:1992:RCE

[AW92] J. Aspnes and O. Waarts. Randomized consensus in expected O(n log^{2} n) operations per processor. In IEEE [IEE92], pages 137–146. CODEN ASF-PDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Auer:1995:TBD

[AW95] P. Auer and M. K. Warmuth. Tracking the best disjunction. In IEEE [IEE95], pages 312–321. CODEN ASF-PDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Andrews:1998:AND

[AZ98] M. Andrews and L. Zhang. The access network design problem. In IEEE [IEE98], pages 40– 49. CODEN ASFPDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Ben-Amram:1991:LBD

[BAG91] A. M. Ben-Amram and Z. Galil.

Lower bounds for data structure problems on RAMs. In IEEE [IEE91], pages 622–631. CODEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Ben-Amram:1993:WCW

[BAG93] A. M. Ben-Amram and Z. Galil.

What can we sort in o(n log n) time? In IEEE [IEE93], pages 538–546. CODEN ASF-PDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Barvinok:1993:PTA

[Bar93] A. I. Barvinok. A polynomial time algorithm for counting integral points in polyhedra when the dimension is fixed. In IEEE [IEE93], pages 566– 572. CODEN ASFPDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Barvinok:1995:IGH

[Bar95] A. I. Barvinok. Integral geometry of higher-dimensional polytopes and the average case in combinatorial optimization. In IEEE [IEE95], pages 275–283. CODEN ASF-PDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Bartal:1996:PAM

[Bar96] Y. Bartal. Probabilistic approximation of metric spaces and its algorithmic applications. In IEEE [IEE96], pages 184–193. CODEN ASF-PDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Basu:1997:IAQ

[Bas97] S. Basu. An improved algorithm for quantifier elimination over real closed fields. In IEEE [IEE97], pages 56– 65. CODEN ASFPDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Basu:1998:CTC

[Bas98] S. Basu. On the combinatorial and topological complexity of a single cell. In IEEE [IEE98], pages 606–616. CODEN ASFPDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Bay:1990:DLR

[BB90] P. Bay and G. Bilardi. Deterministic on-line routing on area-universal networks. In IEEE [IEE90], pages 297–306. CODEN ASFPDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Beals:1993:VAM

[BB93] R. Beals and L. Babai. Las Vegas algorithms for matrix groups. In IEEE [IEE93], pages 427–436. CODEN ASF-PDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Beimel:1996:AMA

[BBB^{+}96] A. Beimel, F. Bergadano, N. H. Bshouty, E. Kushilevitz, and S. Varricchio. On the applications of multiplicity automata in learning. In IEEE [IEE96], pages 349–358. CO-DEN ASFPDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Beals:1998:QLB

[BBC^{+}98] R. Beals, H. Buhrman, R. Cleve, M. Mosca, and R. de Wolf. Quantum lower bounds by polynomials. In IEEE [IEE98], pages 352–361. CODEN ASF-PDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Beigel:1991:LET

[BBFG91] R. Beigel, M. Bellare, J. Feigenbaum, and S. Goldwasser. Languages that are easier than their proofs. In IEEE [IEE91], pages 19–28. CODEN ASF-PDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Blum:1999:FCP

[BBK99] A. Blum, C. Burch, and A. Kalai. Finely-competitive paging. In IEEE [IEE99], pages 450–457. CODEN ASF-PDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Blum:1998:LMB

[BBL98] A. Blum, C. Burcht, and J. Langford. On learning monotone Boolean functions. In IEEE [IEE98], pages 408– 415. CODEN ASFPDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Boyar:1991:SZK

[BBP91] J. Boyar, G. Brassard, and R. Peralta. Subquadratic zeroknowledge. In IEEE [IEE91], pages 69–78. CODEN ASF-PDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Beame:1990:TST

[BBR^{+}90] P. Beame, A. Borodin, P. Raghavan, W. L. Ruzzo, and M. Tompa. Time-space tradeoffs for undirected graph traversal. In IEEE [IEE90], pages 429–438. CODEN ASF-PDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Bartal:1997:GOU

[BBR97] Y. Bartal, J. W. Byers, and D. Raz. Global optimization using local information with applications to flow control. In IEEE [IEE97], pages 303– 312. CODEN ASFPDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Bshouty:1992:ELF

[BC92] N. H. Bshouty and R. Cleve.

On the exact learning of formulas in parallel. In IEEE [IEE92], pages 513–522. CO-DEN ASFPDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Blum:1993:LAI

[BC93] A. Blum and P. Chalasani.

An on-line algorithm for improving performance in navigation. In IEEE [IEE93], pages 2–11. CODEN ASF-PDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Buhrman:1999:BSE

[BCDZ99] H. Buhrman, R. Cleve, R. De Wolf, and C. Zalka. Bounds for small-error and zero-error quantum algorithms. In IEEE [IEE99], pages 358–368. CO-DEN ASFPDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Bshouty:1991:SDT

[BCE91] N. H. Bshouty, R. Cleve, and W. Eberly. Size-depth tradeoffs for algebraic formulae. In IEEE [IEE91], pages 334–341. CODEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Borgs:1999:TMS

[BCF^{+}99] C. Borgs, J. T. Chayes, A. Frieze, Jeong Han Kim, P. Tetali, E. Vigoda, and Van Ha Vu. Torpid mixing of some Monte Carlo Markov chain algorithms in statistical physics. In IEEE [IEE99], pages 218–229. CODEN ASF-PDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Bshouty:1994:LDG

[BCH94] N. H. Bshouty, Zhixiang Chen, and S. Homer. On learning discretized geometric concepts. In Goldwasser [Gol94], pages 54–63. CODEN ASF-PDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Bellare:1995:LTC

[BCH^{+}95] M. Bellare, D. Coppersmith, J. Håstad, M. Kiwi, and M. Sudan. Linearity testing in characteristic two. In IEEE [IEE95], pages 432–441. CO-DEN ASFPDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Brassard:1993:QBC

[BCJL93] G. Brassard, C. Crepeau, R. Jozsa, and D. Langlois. A quantum bit commitment scheme provably unbreakable by both parties. In IEEE [IEE93], pages 362–371. CO-DEN ASFPDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Bellare:1996:PFR

[BCK96] M. Bellare, R. Canetti, and H. Krawczyk. Pseudorandom functions revisited: the cascade construction and its concrete security. In IEEE [IEE96], pages 514–523. CO-DEN ASFPDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Bronnimann:1993:PRS

[BCM93] H. Bronnimann, B. Chazelle, and J. Matousek. Product range spaces, sensitive sampling, and derandomization. In IEEE [IEE93], pages 400– 409. CODEN ASFPDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Bubley:1997:PCT

[BD97] R. Bubley and M. Dyer. Path coupling: A technique for proving rapid mixing in Markov chains. In IEEE [IEE97], pages 223–231. CODEN ASF-PDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Bellare:1997:CST

[BDJR97] M. Bellare, A. Desai, E. Jokipii, and P. Rogaway. A concrete security treatment of symmetric encryption. In IEEE [IEE97], pages 394–403. CODEN ASF-PDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Barnes:1993:TSL

[BE93] G. Barnes and J. A. Edmonds.

Time-space lower bounds for directed s–t connectivity on JAG models. In IEEE [IEE93], pages 228–237. CODEN ASF-PDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Beigel:1995:CTN

[BE95] R. Beigel and D. Eppstein. 3coloring in time O(1.3446^{n}): a no-MIS algorithm. In IEEE [IEE95], pages 444–452. CO-DEN ASFPDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Beals:1995:AMG

[Bea95] R. Beals. Algorithms for matrix groups and the Tits alternative. In IEEE [IEE95], pages 593–602. CODEN ASF-PDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Bern:1990:PGM

[BEG90] M. Bern, D. Eppstein, and J. Gilbert. Provably good mesh generation. In IEEE [IEE90], pages 231–241. CODEN ASF-PDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Blum:1991:CCM

[BEG^{+}91] M. Blum, W. Evans, P. Gemmell, S. Kannan, and M. Naor. Checking the correctness of memories. In IEEE [IEE91], pages 90–99. CODEN ASF-PDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Bonet:1998:ESB

[BEGJ98] M. L. Bonet, J. L. Esteban, N. Galesi, and J. Johannsen. Exponential separations between restricted resolution and cutting planes proof systems. In IEEE [IEE98], pages 638– 647. CODEN ASFPDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Benczur:1995:RCW

[Ben95] A. A. Benczur. A representation of cuts within 6/5 times the edge connectivity with applications. In IEEE [IEE95], pages 92–102. CO-DEN ASFPDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Babai:1990:CHP

[BF90] L. Babai and L. Fortnow. A characterization of Hash P by arithmetic straight line programs. In IEEE [IEE90], pages 26–34. CODEN ASF-PDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Babai:1991:ART

[BF91] L. Babai and K. Friedl. Approximate representation theory of finite groups. In IEEE [IEE91], pages 733–742. CO-DEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991.

IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Blum:1996:PTA

[BFKV96] A. Blum, A. Frieze, R. Kannan, and S. Vempala. A polynomial-time algorithm for learning noisy linear threshold functions. In IEEE [IEE96], pages 330–338. CODEN ASF-PDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Babai:1990:NET

[BFL90] L. Babai, L. Fortnow, and C. Lund. Nondeterministic exponential time has two-prover interactive protocols. In IEEE [IEE90], pages 16–25. CODEN ASFPDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Buhrman:1995:UAS

[BFT95] H. Buhrman, L. Fortnow, and L. Torenvliet. Using autoreducibility to separate complexity classes. In IEEE [IEE95], pages 520–527. CODEN ASF-PDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Broder:1996:GAD

[BFU96] A. Z. Broder, A. M. Frieze, and E. Upfal. A general approach to dynamic packet routing with bounded buffers. In IEEE [IEE96], pages 390–399. CODEN ASFPDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Blum:1992:TCT

[BG92] M. Blum and O. Goldreich.

Towards a computational theory of statistical tests. In IEEE [IEE92], pages 406–416. CODEN ASFPDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Bonet:1999:SPS

[BG99] M. L. Bonet and N. Galesi.

A study of proof search algorithms for resolution and polynomial calculus. In IEEE [IEE99], pages 422–431. CO-DEN ASFPDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Bellare:1990:RIP

[BGG90] M. Bellare, O. Goldreich, and S. Goldwasser. Randomness in interactive proofs. In IEEE [IEE90], pages 563–572. CO-DEN ASFPDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Brieden:1998:ADR

[BGK^{+}98] A. Brieden, P. Gritzmann, R. Kannan, V. Klee, L. Lovasz, and M. Simonovits. Approximation of diameters: randomization doesn’t help. In IEEE [IEE98], pages 244–251. CODEN ASFPDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Beimel:1995:LBM

[BGP95] A. Beimel, A. Gal, and M. Paterson. Lower bounds for monotone span programs.

In IEEE [IEE95], pages 674–681. CODEN ASF-PDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Bellare:1995:FBP

[BGS95] M. Bellare, O. Goldreich, and M. Sudan. Free bits, PCPs and non-approximability-towards tight results. In IEEE [IEE95], pages 422–431. CODEN ASF-PDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Barve:1995:ACP

[BGV95] R. D. Barve, E. F. Grove, and J. S. Vitter. Applicationcontrolled paging for a shared cache. In IEEE [IEE95], pages 204–213. CODEN ASF-PDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Broder:1998:IRW

[BH98] A. Broder and M. R. Henzinger. Information retrieval on the Web. In IEEE [IEE98], page 6. CODEN ASFPDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Babai:1990:DFG

[BHK^{+}90] L. Babai, G. Hetyei, W. M. Kantor, A. Lubotzky, and A. Seress. On the diameter of finite groups. In IEEE [IEE90], pages 857–865. CODEN ASF-PDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Beigel:1995:FDF

[BHK95] R. Beigel, W. Hurwood, and N. Kahale. Fault diagnosis in a flash. In IEEE [IEE95], pages 571–580. CODEN ASF-PDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Beame:1994:LBH

[BIK^{+}94] P. Beame, R. Impagliazzo, J. Krajicek, T. Pitassi, and P. Pudlak. Lower bounds on Hilbert’s Nullstellensatz and propositional proofs. In Goldwasser [Gol94], pages 794–806. CODEN ASFPDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Bellare:1997:DPR

[BIN97] M. Bellare, R. Impagliazzo, and M. Naor. Does parallel repetition lower the error in computationally sound protocols? In IEEE [IEE97], pages 374–383. CODEN ASF-PDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Beame:1995:IDL

[BIP95] P. Beame, R. Impagliazzo, and T. Pitassi. Improved depth lower bounds for small distance connectivity. In IEEE [IEE95], pages 692–701. CO-DEN ASFPDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Bang-Jensen:1997:ECA

[BJJ97] J. Bang-Jensen and T. Jordan.

Edge-connectivity augmentation preserving simplicity. In IEEE [IEE97], pages 486–495. CODEN ASFPDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Berkman:1990:STL

[BJK^{+}90] O. Berkman, J. Jaja, S. Krishnamurthy, R. Thurimella, and U. Vishkin. Some triply-logarithmic parallel algorithms. In IEEE [IEE90], pages 871–881. CODEN ASF-PDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Blum:1993:LIK

[BK93] A. Blum and R. Kannan. Learning an intersection of k halfspaces over a uniform distribution. In IEEE [IEE93], pages 312–320. CODEN ASF-PDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Baruah:1991:LSP

[BKM^{+}91] S. Baruah, G. Koren, B. Mishra, A. Raghunathan, L. Rosier, and D. Shasha. On-line scheduling in the presence of overload. In IEEE [IEE91], pages 100–110. CODEN ASF-PDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Bhatia:1995:LTS

[BKN95] R. Bhatia, S. Khuller, and J. Naor. The loading time scheduling problem. In IEEE [IEE95], pages 72–81. CODEN ASFPDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Blum:1992:DTB

[BKRS92] A. Blum, H. Karloff, Y. Rabani, and M. Saks. A decomposition theorem and bounds for randomized server problems. In IEEE [IEE92], pages 197– 207. CODEN ASFPDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Blumofe:1994:SMC

[BL94] R. D. Blumofe and C. E.

Leiserson. Scheduling multithreaded computations by work stealing. In Goldwasser [Gol94], pages 356–368. CO-DEN ASFPDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Blaser:1998:BPM

[Bla98] M. Blaser. Bivariate polynomial multiplication. In IEEE [IEE98], pages 186–191. CO-DEN ASFPDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Blaser:1999:LBR

[Bla99] M. Blaser. A 5/2n^{2}-lower bound for the rank of n × n-matrix multiplication over arbitrary fields. In IEEE [IEE99], pages 45–50. CODEN ASFPDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Blomer:1991:CSR

[Blo91] J. Blomer. Computing sums of radicals in polynomial time. In IEEE [IEE91], pages 670–677. CODEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Blomer:1992:HDR

[Blo92] J. Blomer. How to denest Ramanujan’s nested radicals. In IEEE [IEE92], pages 447– 456. CODEN ASFPDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Blum:1990:SDF

[Blu90a] A. Blum. Separating distributionfree and mistake-bound learning models over the Boolean domain. In IEEE [IEE90], pages 211–218. CODEN ASF-PDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Blum:1990:STA

[Blu90b] A. Blum. Some tools for approximate 3-coloring. In IEEE [IEE90], pages 554–562. CODEN ASFPDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Bshouty:1995:SLA

[BM95] N. H. Bshouty and Y. Mansour. Simple learning algorithms for decision trees and multivariate polynomials. In IEEE [IEE95], pages 304–311. CODEN ASF-PDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Bornstein:1997:PEO

[BMMR97] C. Bornstein, B. Maggs, G. Miller, and R. Ravi. Parallelizing elimination orders with linear fill. In IEEE [IEE97], pages 274–283. CODEN ASF-PDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Boykin:1999:UFT

[BMP^{+}99] P. O. Boykin, T. Mor, M. Pulver, V. Roychowdhury, and F. Vatan. On universal and fault-tolerant quantum computing: a novel basis and a new constructive proof of universality for Shor’s basis. In IEEE [IEE99], pages 486– 494. CODEN ASFPDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Ben-Or:1994:ACT

[BO94] M. Ben-Or. Algebraic computation trees in characteristic p < 0. In Goldwasser [Gol94], pages 534–539. CODEN ASF-PDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Bouziane:1998:PRA

[Bou98] Z. Bouziane. A primitive recursive algorithm for the general Petri net reachability problem. In IEEE [IEE98], pages 130– 136. CODEN ASFPDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Bini:1992:IPP

[BP92] D. Bini and V. Pan. Improved parallel polynomial division and its extensions. In IEEE [IEE92], pages 131–136. CODEN ASFPDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Bafna:1993:GRS

[BP93] V. Bafna and P. A. Pevzner. Genome rearrangements and sorting by reversals. In IEEE [IEE93], pages 148–157. CO-DEN ASFPDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Beame:1996:SIR

[BP96] P. Beame and T. Pitassi. Simplified and improved resolution lower bounds. In IEEE [IEE96], pages 274–282. CO-DEN ASFPDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Basu:1994:CAC

[BPR94] S. Basu, R. Pollack, and M.- F. Roy. On the combinatorial and algebraic complexity of quantifier elimination. In Goldwasser [Gol94], pages 632– 641. CODEN ASFPDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Bonet:1997:NFI

[BPR97] M. L. Bonet, T. Pitassi, and R. Raz. No feasible interpolation for TC^{0}-Frege proofs. In IEEE [IEE97], pages 254– 263. CODEN ASFPDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Buss:1990:PCA

[BPT90] S. Buss, C. H. Papadimitriou, and N. J. Tsitsiklis. On the predictability of coupled automata: an allegory about chaos. In IEEE [IEE90], pages 788–793. CODEN ASF-PDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Bartal:1992:DSP

[BR92] Y. Bartal and A. Rosen. The distributed k-server problem — a competitive distributed translator for k-server algorithms. In IEEE [IEE92], pages 344–353. CODEN ASF-PDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Bellare:1994:REO

[BR94] M. Bellare and J. Rompel. Randomness-efficient oblivious sampling. In Goldwasser [Gol94], pages 276–287. CO-DEN ASFPDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Bruck:1990:PTF

[BS90] J. Bruck and R. Smolensky. Polynomial threshold functions, AC functions and spectrum norms. In IEEE [IEE90], pages 632–641. CODEN ASF-PDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Bender:1994:PTE

[BS94] M. A. Bender and D. K. Slonim. The power of team exploration: two robots can learn unlabeled directed graphs. In Goldwasser [Gol94], pages 75– 85. CODEN ASFPDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Bshouty:1993:ELM

[Bsh93] N. H. Bshouty. Exact learning via the Monotone theory. In IEEE [IEE93], pages 302– 311. CODEN ASFPDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Ben-Sasson:1999:RCH

[BSI99] E. Ben-Sasson and R. Impagliazzo. Random CNFs are hard for the polynomial calculus. In IEEE [IEE99], pages 415– 421. CODEN ASFPDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Buchsbaum:1992:DSB

[BST92] A. L. Buchsbaum, R. Sundar, and R. E. Tarjan. Data structural bootstrapping, linear path compression, and catenable heap ordered double ended queues. In IEEE [IEE92], pages 40–49. CODEN ASFPDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428.

LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Beame:1998:TST

[BST98] P. Beame, M. Saks, and J. S. Thathachar. Time-space tradeoffs for branching programs. In IEEE [IEE98], pages 254– 263. CODEN ASFPDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Beigel:1991:ACC

[BT91] R. Beigel and J. Tarui. On ACC (circuit complexity). In IEEE [IEE91], pages 783–792. CODEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Beame:1990:CST

[BTY90] P. Beame, M. Tompa, and P. Yan. Communication-space tradeoffs for unrestricted protocols. In IEEE [IEE90], pages 420–428. CODEN ASF-PDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Barve:1999:TFM

[BV99] R. D. Barve and J. S. Vitter.

A theoretical framework for memory-adaptive algorithms. In IEEE [IEE99], pages 273– 284. CODEN ASFPDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Blum:1994:PRC

[BW94] M. Blum and H. Wasserman.

Program result-checking: a theory of testing meets a test of theory. In Goldwasser [Gol94], pages 382–392. CODEN ASF-PDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Callahan:1993:OPA

[Cal93] P. B. Callahan. Optimal parallel all-nearest-neighbors using the well-separated pair decomposition. In IEEE [IEE93], pages 332–340. CODEN ASF-PDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Conforti:1992:CLP

[CC92] M. Conforti and G. Cornuejols. A class of logic problems solvable by linear programming. In IEEE [IEE92], pages 670– 675. CODEN ASFPDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Cole:1993:OFP

[CCG^{+}93] R. Cole, M. Crochemore, Z. Galil, L. Gasieniec, R. Eariharan, S. Muthukrishnan, K. Park, and W. Rytter. Optimally fast parallel algorithms for preprocessing and pattern matching in one and two dimensions. In IEEE [IEE93], pages 248–258. CODEN ASF-PDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Charikar:1998:AFM

[CCG^{+}98] M. Charikar, C. Chekuri, A. Goel, S. Guha, and S. Plotkin. Approximating a finite metric by a small number of tree metrics. In IEEE [IEE98], pages 379–388. CO-DEN ASFPDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Conforti:1997:FEH

[CCKV97] M. Conforti, G. Cornuejols, A. Kapoor, and K. Vuskovic. Finding an even hole in a graph. In IEEE [IEE97], pages 480–485. CODEN ASF-PDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Cachin:1998:OTM

[CCM98] C. Cachin, C. Crepeau, and J. Marcil. Oblivious transfer with a memory-bounded receiver. In IEEE [IEE98], pages 493–502. CODEN ASF-PDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Cai:1998:TSA

[CDZ98] Mao-Cheng Cai, Xiaotie Deng, and Wenan Zang. A TDI system and its application to approximation algorithms. In IEEE [IEE98], pages 227–231. CODEN ASFPDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Chazelle:1990:CCC

[CEG^{+}90] B. Chazelle, H. Edelsbrunner, L. J. Guibas, R. Pollack, R. Seidel, M. Sharir, and J. Snoeyink. Counting and cutting cycles of lines and rods in space. In IEEE [IEE90], pages 242–251. CODEN ASF-PDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Cai:1994:EAC

[CFKL94] Jin-Yi Cai, W. H. Fuchs, D. Kozen, and Zicheng Lui. Efficient average-case algorithms for the modular group. In Goldwasser [Gol94], pages 143– 152. CODEN ASFPDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Canetti:1990:BTB

[CG90] R. Canetti and O. Goldreich. Bounds on tradeoffs between randomness and communication complexity. In IEEE [IEE90], pages 766–775. CODEN ASFPDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Chang:1993:BQA

[CG93] R. Chang and W. I. Gasarch.

On bounded queries and approximation. In IEEE [IEE93], pages 547–556. CODEN ASF-PDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Canetti:1996:IMC

[CG96a] R. Canetti and R. Gennaro.

Incoercible multiparty computation. In IEEE [IEE96], pages 504–513. CODEN ASF-PDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Cunningham:1996:OPM

[CG96b] W. H. Cunningham and J. F. Geelen. The optimal pathmatching problem. In IEEE [IEE96], pages 78–85. CODEN ASFPDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Carter:1998:TOB

[CG98] L. Carter and K. S. Gatlin. Towards an optimal bit-reversal permutation program. In IEEE [IEE98], pages 544–553. CO-DEN ASFPDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Charikar:1999:ICA

[CG99] M. Charikar and S. Guha. Improved combinatorial algorithms for the facility location and k-median problems. In IEEE [IEE99], pages 378– 388. CODEN ASFPDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Colussi:1990:ECS

[CGG90] L. Colussi, Z. Galil, and R. Giancarlo. On the exact complexity of string matching. In IEEE [IEE90], pages 135–144. CODEN ASFPDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Cryan:1998:ETC

[CGG98] M. Cryan, L. A. Goldberg, and P. W. Goldberg. Evolutionary trees can be learned in polynomial time in the twostate general Markov model. In IEEE [IEE98], pages 436– 445. CODEN ASFPDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Chor:1990:PCI

[CGGK90] B. Chor, M. Gereb-Graus, and E. Kushilevitz. Private computations over the integers. In IEEE [IEE90], pages 335–344. CODEN ASF-PDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Chor:1995:PIR

[CGKS95] B. Chor, O. Goldreich, E. Kushilevitz, and M. Sudan. Private information retrieval. In IEEE [IEE95], pages 41–50. CODEN ASFPDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Cole:1992:TBE

[CH92] R. Cole and R. Hariharan. Tighter bounds on the exact complexity of string matching. In IEEE [IEE92], pages 600– 609. CODEN ASFPDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Chazelle:1990:TSP

[Cha90] B. Chazelle. Triangulating a simple polygon in linear time. In IEEE [IEE90], pages 220–230. CODEN ASF-PDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990.

Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Chazelle:1991:OCH

[Cha91] B. Chazelle. An optimal convex hull algorithm and new results on cuttings. In IEEE [IEE91], pages 29–38. CODEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Chaudhuri:1993:SFA

[Cha93a] S. Chaudhuri. Sensitive functions and approximate problems. In IEEE [IEE93], pages 186–193. CODEN ASF-PDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Chazelle:1993:GDR

[Cha93b] B. Chazelle. Geometric discrepancy revisited. In IEEE [IEE93], pages 392–399. CO-DEN ASFPDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Chazelle:1994:SAL

[Cha94] B. Chazelle. A spectral approach to lower bounds. In Goldwasser [Gol94], pages 674– 682. CODEN ASFPDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Chazelle:1997:FDA

[Cha97] B. Chazelle. A faster deterministic algorithm for minimum spanning trees. In IEEE [IEE97], pages 22–31. CODEN ASFPDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Chan:1998:RSH

[Cha98] T. M. Chan. Random sampling, halfspace range reporting, and construction of (≤ k)-levels in three dimensions. In IEEE [IEE98], pages 586– 595. CODEN ASFPDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Chan:1999:DPC

[Cha99] T. M. Chan. Dynamic planar convex hull operations in nearlogarithmic amortized time. In IEEE [IEE99], pages 92–99. CODEN ASFPDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Chen:1999:FDE

[CHH99] Zhi-Zhong Chen, Xin He, and Chun-Hsi Huang. Finding double Euler trails of planar graphs in linear time [CMOS VLSI circuit design]. In IEEE [IEE99], pages 319–329. CO-DEN ASFPDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Chaudhuri:1993:TLB

[CHLT93] S. Chaudhuri, M. Herlihy, N. A. Lynch, and M. R. Tuttle. A tight lower bound for k-set agreement. In IEEE [IEE93], pages 206–215. CODEN ASF-PDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Clote:1993:PCH

[CIK93] P. Clote, A. Ignjatovic, and B. Kapron. Parallel computable higher type functionals. In IEEE [IEE93], pages 72– 81. CODEN ASFPDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Cheng:1990:NRD

[CJ90] S. W. Cheng and R. Janardan. New results on dynamic planar point location. In IEEE [IEE90], pages 96– 105. CODEN ASFPDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Chalasani:1996:AOP

[CJS96] P. Chalasani, S. Jha, and I. Saias. Approximate option pricing. In IEEE [IEE96], pages 244–253. CODEN ASF-PDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Chang:1990:ASM

[CL90] W. I. Chang and E. L. Lawler.

Approximate string matching in sublinear expected time. In IEEE [IEE90], pages 116–124. CODEN ASFPDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Clarkson:1992:SED

[Cla92] K. L. Clarkson. Safe and effective determinant evaluation. In IEEE [IEE92], pages 387– 395. CODEN ASFPDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Clarkson:1994:MOS

[Cla94] K. L. Clarkson. More outputsensitive geometric algorithms. In Goldwasser [Gol94], pages 695–702. CODEN ASFPDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Cole:1995:RBN

[CMS95] R. Cole, B. Maggs, and R. Sitaraman. Routing on butterfly networks with random faults. In IEEE [IEE95], pages 558–570. CODEN ASF-PDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Cai:1997:IWC

[CN97] Jin-Yi Cai and A. P. Nerurkar. An improved worst-case to average-case connection for lattice problems. In IEEE [IEE97], pages 468–477. CO-DEN ASFPDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Cohen:1992:AMF

[Coh92] E. Cohen. Approximate max flow on small depth networks. In IEEE [IEE92], pages 648– 658. CODEN ASFPDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Cohen:1993:FAC

[Coh93] E. Cohen. Fast algorithms for constructing t-spanners and paths with stretch t. In IEEE [IEE93], pages 648–658. CO-DEN ASFPDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Cohen:1994:EST

[Coh94] E. Cohen. Estimating the size of the transitive closure in linear time. In Goldwasser [Gol94], pages 190–200. CO-DEN ASFPDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Cohen:1997:LNP

[Coh97] E. Cohen. Learning noisy perceptrons by a perceptron in polynomial time. In IEEE [IEE97], pages 514–523. CO-DEN ASFPDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Cosmadakis:1993:LRM

[Cos93] S. S. Cosmadakis. Logical reducibility and monadic NP. In IEEE [IEE93], pages 52– 61. CODEN ASFPDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Cover:1996:UDC

[Cov96] T. M. Cover. Universal data compression and portfolio selection. In IEEE [IEE96], pages 534–538. CODEN ASF-PDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Charatonik:1994:SCP

[CP94] W. Charatonik and L. Pacholski. Set constraints with projections are in NEXPTIME. In Goldwasser [Gol94], pages 642–653. CODEN ASFPDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Cole:1990:OAF

[CR90] R. Cole and A. Raghunathan. Online algorithms for finger searching. In IEEE [IEE90], pages 480–489. CODEN ASF-PDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Chaudhuri:1992:CPP

[CR92a] S. Chaudhuri and J. Radhakrishnan. The complexity of parallel prefix problems on small domains. In IEEE [IEE92], pages 638–647. CODEN ASF-PDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Chvatal:1992:MGS

[CR92b] V. Chvatal and B. Reed. Mick gets some (the odds are on his side) (satisfiability). In IEEE [IEE92], pages 620–627. CODEN ASFPDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Chen:1993:UDP

[CR93] S. Chen and J. H. Reif. Using difficulty of prediction to decrease computation: fast sort, priority queue and convex hull on entropy bounded inputs.

In IEEE [IEE93], pages 104– 112. CODEN ASFPDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Charikar:1998:FCD

[CR98] M. Charikar and B. Raghavachari. The finite capacity dial-a-ride problem. In IEEE [IEE98], pages 458–467. CODEN ASF-PDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Coppersmith:1992:LBD

[CS92] D. Coppersmith and B. Schieber. Lower bounds on the depth of monotone arithmetic computations. In IEEE [IEE92], pages 288–295. CODEN ASF-PDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Cai:1995:RHC

[CS95] Jin-Yi Cai and D. Sivakumar. The resolution of a Hartmanis conjecture. In IEEE [IEE95], pages 362–371. CO-DEN ASFPDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Czumaj:1997:RAP

[CS97] A. Czumaj and V. Stemann.

Randomized allocation processes. In IEEE [IEE97], pages 194–203. CODEN ASF-PDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Cardoze:1998:PMS

[CS98] D. E. Cardoze and L. J. Schulman. Pattern matching for spatial point sets. In IEEE [IEE98], pages 156–165. CO-DEN ASFPDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Cai:1997:CDC

[CSS97] Jin-Yi Cai, D. Sivakumar, and M. Strauss. Constant depth circuits and the Lutz hypothesis. In IEEE [IEE97], pages 595–604. CODEN ASF-PDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Cheriyan:1996:AMS

[CT96] J. Cheriyan and R. Thurimella. Approximating minimum-size k-connected spanning subgraphs via matching. In IEEE [IEE96], pages 292–301. CO-DEN ASFPDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

deAlfaro:1998:CRG

[dAHK98] L. de Alfaro, T. A. Henzinger, and O. Kupferman. Concurrent reachability games. In IEEE [IEE98], pages 564–575. CODEN ASFPDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Dasgupta:1999:LMG

[Das99] S. Dasgupta. Learning mixtures of Gaussians. In IEEE [IEE99], pages 634–644. CO-DEN ASFPDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

deBerg:1990:HSR

[dBO90] M. de Berg and M. H. Overmars. Hidden surface removal for axis-parallel polyhedra. In IEEE [IEE90], pages 252–261. CODEN ASF-PDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Donald:1991:CCH

[DC91] B. R. Donald and D. R. Chang.

On the complexity of computing the homology type of a triangulation. In IEEE [IEE91], pages 650–661. CODEN ASF-PDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

DeSantis:1994:MFC

[DDPY94] A. De Santis, G. Di Crescenzo, G. Persiano, and M. Yung. On monotone formula closure of SZK. In Goldwasser [Gol94], pages 454–465. CODEN ASF-PDV. ISBN 0-8186-6582-3.

ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Dolev:1990:PSM

[DDWY90] D. Dolev, C. Dwork, O. Waarts, and M. Yung. Perfectly secure message transmission. In IEEE [IEE90], pages 36–45. CODEN ASFPDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Dey:1997:IBP

[Dey97] T. K. Dey. Improved bounds on planar k-sets and k-levels. In IEEE [IEE97], pages 156– 161. CODEN ASFPDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Dyer:1999:CIS

[DFJ99] M. Dyer, A. Frieze, and M. Jerrum. On counting independent sets in sparse graphs. In IEEE [IEE99], pages 210– 217. CODEN ASFPDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Dey:1995:OAC

[DG95] T. K. Dey and S. Guha. Optimal algorithms for curves on surfaces. In IEEE [IEE95], pages 266–274. CODEN ASF-PDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Dinitz:1998:SSU

[DGG98] Y. Dinitz, N. Garg, and M. X.

Goemans. On the single-source unsplittable flow problem. In IEEE [IEE98], pages 290–299. CODEN ASFPDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Dubiner:1990:FTP

[DGM90] M. Dubiner, Z. Galil, and E. Magen. Faster tree pattern matching. In IEEE [IEE90], pages 145–150. CODEN ASF-PDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Du:1990:APL

[DH90] D.-Z. Du and F. K. Hwang. An approach for proving lower bounds: solution of Gilbert– Pollak’s conjecture on Steiner ratio. In IEEE [IEE90], pages 76–85. CODEN ASF-PDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Dhagat:1994:PLI

[DH94] A. Dhagat and L. Hellerstein. PAC learning with irrelevant attributes. In Goldwasser [Gol94], pages 64–74. CODEN ASFPDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Dor:1996:APA

[DHZ96] D. Dor, S. Halperin, and U. Zwick. All pairs almost shortest paths. In IEEE [IEE96], pages 452–461. CO-DEN ASFPDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Dagum:1995:OAM

[DKLR95] P. Dagum, R. Karp, M. Luby, and S. Ross. An optimal algorithm for Monte Carlo estimation. In IEEE [IEE95], pages 142–149. CODEN ASF-PDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Deng:1991:HLU

[DKP91] X. Deng, T. Kameda, and C. Papadimitriou. How to learn an unknown environment. In IEEE [IEE91], pages 298–303. CODEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Dinur:1998:ACW

[DKS98] I. Dinur, G. Kindler, and S. Safra. Approximating-CVP to within almost-polynomial factors is NP-hard. In IEEE [IEE98], pages 99–109. CO-DEN ASFPDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

delaVega:1998:RAS

[dlVK98] W. Fernandez de la Vega and C. Kenyon. A randomized approximation scheme for metric MAX-CUT. In IEEE [IEE98], pages 468–471. CODEN ASF-PDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Dumer:1999:HAM

[DMS99] I. Dumer, D. Micciancio, and M. Sudan. Hardness of approximating the minimum distance of a linear code. In IEEE [IEE99], pages 475–484. CODEN ASFPDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Dwork:1999:MF

[DNRS99] C. Dwork, M. Naor, O. Reingold, and L. Stockmeyer. Magic functions. In IEEE [IEE99], pages 523–534. CO-DEN ASFPDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Deng:1990:EUG

[DP90] X. Deng and C. H. Papadimitriou. Exploring an unknown graph. In IEEE [IEE90], pages 355–361. CODEN ASF-PDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

DeSantis:1992:ZKP

[DP92] A. De Santis and G. Persiano. Zero-knowledge proofs of knowledge without interaction. In IEEE [IEE92], pages 427–436. CODEN ASF-PDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Durr:1996:DPU

[DS96] C. Durr and M. Santha. A decision procedure for unitary linear quantum cellular automata. In IEEE [IEE96], pages 38–45. CODEN ASF-PDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Dubiner:1992:APP

[DZ92] M. Dubiner and U. Zwick. Amplification and percolation (probabilistic Boolean functions). In IEEE [IEE92], pages 258–267. CODEN ASF-PDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Dor:1996:MSR

[DZ96] D. Dor and U. Zwick. Median selection requires (2 + ϵ)n comparisons. In IEEE [IEE96], pages 125–134. CO-DEN ASFPDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Du:1991:BHE

[DZF91] D.-Z. Du, Y. Zhang, and Q. Feng. On better heuristic for Euclidean Steiner minimum trees. In IEEE [IEE91], pages 431–439. CODEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Eberly:1992:EBM

[Ebe92] W. Eberly. On efficient band matrix arithmetic. In IEEE [IEE92], pages 457–463. CO-DEN ASFPDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Emiris:1991:GAR

[EC91] I. Emiris and J. Canny. A general approach to removing degeneracies. In IEEE [IEE91], pages 405–413. CO-DEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Edelsbrunner:1995:ADN

[Ede95] H. Edelsbrunner. Algebraic decomposition of non-convex polyhedra. In IEEE [IEE95], pages 248–257. CODEN ASF-PDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Egidi:1993:CTA

[Egi93] L. Egidi. The complexity of the theory of p-adic numbers. In IEEE [IEE93], pages 412– 421. CODEN ASFPDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Etzioni:1996:EIG

[EHJ^{+}96] O. Etzioni, S. Hanks, T. Jiang, R. M. Karp, O. Madani, and O. Waarts. Efficient information gathering on the Internet. In IEEE [IEE96], pages 234– 243. CODEN ASFPDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Edmonds:1991:CCT

[EIRS91] J. Edmonds, R. Impagliazzo, S. Rudich, and J. Sgall. Communication complexity towards lower bounds on circuit depth. In IEEE [IEE91], pages 249–257. CODEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Emerson:1991:TAM

[EJ91] E. A. Emerson and C. S. Jutla. Tree automata, mucalculus and determinacy. In IEEE [IEE91], pages 368–377. CODEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428.

LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Ergun:1996:ACP

[EKR96] F. Ergun, S. R. Kumar, and R. Rubinfeld. Approximate checking of polynomials and functional equations. In IEEE [IEE96], pages 592–601. CO-DEN ASFPDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Erickson:1994:OSC

[ELL94] K. B. Erickson, R. E. Ladner, and A. LaMarca. Optimizing static calendar queues. In Goldwasser [Gol94], pages 732–743. CODEN ASFPDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Even:1990:CSD

[ELW90] S. Even, A. Litman, and P. Winkler. Computing with snakes in directed networks of automata. In IEEE [IEE90], pages 740–745. CODEN ASF-PDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990.

Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Even:1995:DCA

[ENRS95] G. Even, J. Naor, S. Rao, and B. Schieber. Divideand-conquer approximation algorithms via spreading metrics. In IEEE [IEE95], pages 62–71. CODEN ASF-PDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Even:1996:AAS

[ENZ96] G. Even, J. S. Naor, and L. Zosin. An 8-approximation algorithm for the subset feedback vertex set problem. In IEEE [IEE96], pages 310–319. CODEN ASFPDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Eppstein:1991:DTD

[Epp91] D. Eppstein. Dynamic threedimensional linear programming. In IEEE [IEE91], pages 488–494. CODEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Eppstein:1992:STS

[Epp92] D. Eppstein. Sparsification — a technique for speeding up dynamic graph algorithms. In IEEE [IEE92], pages 60– 69. CODEN ASFPDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Eppstein:1994:FKS

[Epp94] D. Eppstein. Finding the k shortest paths. In Goldwasser [Gol94], pages 154–165. CODEN ASFPDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Eppstein:1999:SPE

[Epp99] D. Eppstein. Setting parameters by example. In IEEE [IEE99], pages 309–318. CO-DEN ASFPDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Erickson:1996:NLB

[Eri96] J. Erickson. New lower bounds for halfspace emptiness. In IEEE [IEE96], pages 472–481. CODEN ASFPDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Erickson:1993:BLB

[ES93a] J. Erickson and R. Seidel. Better lower bounds on detecting affine and spherical degeneracies. In IEEE [IEE93], pages 528–536. CODEN ASF-PDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Evans:1993:SPA

[ES93b] W. Evans and L. J. Schulman. Signal propagation, with application to a lower bound on the depth of noisy formulas. In IEEE [IEE93], pages 594– 603. CODEN ASFPDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Edelsbrunner:1991:QTA

[ET91] H. Edelsbrunner and T. S. Tan. A quadratic time algorithm for the minmax length triangulation. In IEEE [IEE91], pages 414–423. CODEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

El-Yaniv:1992:CAF

[EYFKT92] R. El-Yaniv, A. Fiat, R. Karp, and G. Turpin. Competitive analysis of financial games. In IEEE [IEE92], pages 327– 333. CODEN ASFPDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Farach:1997:OST

[Far97] M. Farach. Optimal suffix tree construction with large alphabets. In IEEE [IEE97], pages 137–143. CODEN ASF-PDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Farach-Colton:1999:ANN

[FCI99] M. Farach-Colton and P. Indyk. Approximate nearest neighbor algorithms for Hausdorff metrics via embeddings.

In IEEE [IEE99], pages 171– 179. CODEN ASFPDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Feige:1993:RTS

[Fei93] U. Feige. A randomized timespace tradeoff of O˜(mR∘) for USTCON. In IEEE [IEE93], pages 238–246. CODEN ASF-PDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Feige:1999:NSP

[Fei99] U. Feige. Noncryptographic selection protocols. In IEEE [IEE99], pages 142–152. CO-DEN ASFPDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Fiat:1991:CAL

[FFK^{+}91] A. Fiat, D. P. Foster, H. Karloff, Y. Rabani, Y. Ravid, and S. Viswanathan. Competitive algorithms for layered graph traversal. In IEEE [IEE91], pages 288–297. CODEN ASF-PDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Fenner:1992:ICH

[FFK92] S. Fenner, L. Fortnow, and S. A. Kurtz. The isomorphism conjecture holds relative to an oracle. In IEEE [IEE92], pages 30–39. CODEN ASFPDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Farach:1998:OMB

[FFM98] M. Farach, P. Ferragina, and S. Muthukrishnan. Overcoming the memory bottleneck in suffix tree construction. In IEEE [IEE98], pages 174–183. CODEN ASFPDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Ferragina:1995:OLS

[FG95] P. Ferragina and R. Grossi.

Optimal on-line search and sublinear time update in string matching. In IEEE [IEE95], pages 604–612. CODEN ASF-PDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Feige:1991:ACA

[FGL^{+}91] U. Feige, S. Goldwasser, L. Lovasz, S. Safra, and M. Szegedy. Approximating clique is almost NP-complete. In IEEE [IEE91], pages 2–12. CODEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Frankel:1997:ORP

[FGMY97] Y. Frankel, P. Gemmell, P. D. MacKenzie, and Moti Yung. Optimal-resilience proactive public-key cryptosystems. In IEEE [IEE97], pages 384–393. CODEN ASFPDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Franklin:1993:EGG

[FGY93] M. Franklin, Z. Galil, and M. Yung. Eavesdropping games: a graph-theoretic approach to privacy in distributed systems. In IEEE [IEE93], pages 670–679. CO-DEN ASFPDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Formann:1990:DGP

[FHH^{+}90] M. Formann, T. Hagerup, J. Haralambides, M. Kaufmann, F. T. Leighton, A. Simvonis, E. Welzl, and G. Woeginger. Drawing graphs in the plane with high resolution. In IEEE [IEE90], pages 86–95. CODEN ASFPDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Fischlin:1997:LBS

[Fis97] M. Fischlin. Lower bounds for the signature size of incremental schemes. In IEEE [IEE97], pages 438–447. CO-DEN ASFPDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Frieze:1996:LLT

[FJK96] A. Frieze, M. Jerrum, and R. Kannan. Learning linear transformations. In IEEE [IEE96], pages 359–368. CO-DEN ASFPDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Frieze:1996:RLA

[FK96] A. Frieze and R. Kannan. The regularity lemma and approximation schemes for dense problems. In IEEE [IEE96], pages 12–20. CODEN ASF-PDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Feige:1998:HFL

[FK98] U. Feige and J. Kilian. Heuristics for finding large independent sets, with applications to coloring semi-random graphs. In IEEE [IEE98], pages 674– 683. CODEN ASFPDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Freund:1995:EAL

[FKM^{+}95] Y. Freund, M. Kearns, Y. Mansour, D. Ron, R. Rubinfeld, and R. E. Schapire. Efficient algorithms for learning to play repeated games against computationally bounded adversaries. In IEEE [IEE95], pages 332–341. CODEN ASF-PDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Feder:1991:ACC

[FKN91] T. Feder, E. Kushilevitz, and M. Naor. Amortized communication complexity. In IEEE [IEE91], pages 239–248. CO-DEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Feigenbaum:1999:ADA

[FKSV99] J. Feigenbaum, S. Kannan, M. Strauss, and M. Viswanathan. An approximate L^{1}-difference algorithm for massive data streams. In IEEE [IEE99], pages 501–511. CODEN ASF-PDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Frieze:1998:FMC

[FKV98] A. Frieze, R. Kannan, and S. Vempala. Fast Monte-Carlo algorithms for finding low-rank approximations. In IEEE [IEE98], pages 370–378. CODEN ASFPDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Fleischer:1999:AFM

[Fle99]  L. K. Fleischer.  Approx-

imating fractional multicommodity flow independent of the number of commodities. In IEEE [IEE99], pages 24– 31. CODEN ASFPDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Frigo:1999:COA

[FLPR99] M. Frigo, C. E. Leiserson, H. Prokop, and S. Ramachandran. Cache-oblivious algorithms. In IEEE [IEE99], pages 285–297. CODEN ASF-PDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Feige:1990:MNI

[FLS90] U. Feige, D. Lapidot, and A. Shamir. Multiple noninteractive zero knowledge proofs based on a single random string. In IEEE [IEE90], pages 308–317. CODEN ASF-PDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Fiat:1997:TOP

[FM97] A. Fiat and M. Mendel. Truly online paging with locality of reference. In IEEE [IEE97], pages 326–335. CODEN ASF-PDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Fich:1990:P

[FMP90] F. E. Fich, J. I. Munro, and P. V. Poblete. Permuting. In IEEE [IEE90], pages 372–379. CODEN ASF-PDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Fiat:1995:CAT

[FMRW95] A. Fiat, Y. Mansour, A. Rosen, and O. Waarts. Competitive access time via dynamic storage rearrangement. In IEEE [IEE95], pages 392–401. CO-DEN ASFPDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Frandsen:1993:DWP

[FMS93] G. S. Frandsen, P. B. Miltersen, and S. Skyum. Dynamic word problems. In IEEE [IEE93], pages 470–479. CODEN ASFPDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Feige:1992:EAH

[FR92] U. Feige and P. Raghavan. Exact analysis of hot-potato routing. In IEEE [IEE92], pages 553–562. CODEN ASF-PDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Feldman:1999:DSN

[FR99] J. Feldman and M. Ruhl. The Directed Steiner Network problem is tractable for a constant number of terminals. In IEEE [IEE99], pages 299–308. CODEN ASFPDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Frank:1990:AGM

[Fra90] A. Frank. Augmenting graphs to meet edge-connectivity requirements. In IEEE [IEE90], pages 708–718. CODEN ASF-PDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Frederickson:1991:ADS

[Fre91] G. N. Frederickson. Ambivalent data structures for dynamic 2-edge-connectivity and k smallest spanning trees. In IEEE [IEE91], pages 632–641. CODEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Friedman:1992:BEP

[Fri92] J. Friedman. On the bit extraction problem. In IEEE [IEE92], pages 314–319. CODEN ASF-PDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Fiat:1990:CSA

[FRR90] A. Fiat, Y. Rabani, and Y. Ravid. Competitive k-server algorithms. In IEEE [IEE90], pages 454–463. CODEN ASF-PDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Felperin:1992:TWR

[FRU92] S. Felperin, P. Raghavan, and E. Upfal. A theory of wormhole routing in parallel computers. In IEEE [IEE92], pages 563– 572. CODEN ASFPDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Feldmann:1991:DSP

[FST91] S. Feldmann, J. Sgall, and S.- H. Teng. Dynamic scheduling on parallel machines. In IEEE [IEE91], pages 111–120.

CODEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Farach:1994:OET

[FT94] M. Farach and M. Thorup. Optimal evolutionary tree comparison by sparse dynamic programming. In Goldwasser [Gol94], pages 770–779. CO-DEN ASFPDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Fulk:1990:RSI

[Ful90] M. A. Fulk. Robust separations in inductive inference. In IEEE [IEE90], pages 405–410. CODEN ASFPDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Furer:1995:IHR

[Fur95] M. Furer. Improved hardness results for approximating the chromatic number. In IEEE [IEE95], pages 414–421. CO-DEN ASFPDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Fredman:1990:TDA

[FW90] M. L. Fredman and D. E. Willard. Trans-dichotomous algorithms for minimum spanning trees and shortest paths. In IEEE [IEE90], pages 719– 725. CODEN ASFPDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Gabow:1993:FCS

[Gab93] H. N. Gabow. A framework for cost-scaling algorithms for submodular flow problems. In IEEE [IEE93], pages 449–458. CODEN ASFPDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Gacs:1997:RCA

[Gac97] P. Gacs. Reliable cellular automata with self-organization. In IEEE [IEE97], pages 90– 99. CODEN ASFPDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Gal:1991:LBC

[Gal91] A. Gal. Lower bounds for the complexity of reliable Boolean circuits with noisy gates. In IEEE [IEE91], pages 594–601. CODEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Gamarnik:1998:SAQ

[Gam98] D. Gamarnik. Stability of adversarial queues via fluid models. In IEEE [IEE98], pages 60– 70. CODEN ASFPDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Gartner:1992:SAA

[Gar92] B. Gartner. A subexponential algorithm for abstract optimization problems. In IEEE [IEE92], pages 464–472. CO-DEN ASFPDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number

Garg:1996:AMT

[Gar96] N. Garg. A 3-approximation for the minimum tree spanning k vertices. In IEEE [IEE96], pages 302–309. CO-DEN ASFPDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Gazit:1991:DPA

[Gaz91] H. Gazit. A deterministic parallel algorithm for planar graphs isomorphism. In IEEE [IEE91], pages 723–732. CO-DEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Ge:1993:TEM

[Ge93] Guoqiang Ge. Testing equalities of multiplicative representations in polynomial time. In IEEE [IEE93], pages 422– 426. CODEN ASFPDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Goldreich:1991:FTC

[GGL91] O. Goldreich, S. Goldwasser, and N. Linial. Fault-tolerant computation in the full information model. In IEEE [IEE91], pages 447–457. CO-DEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Goldreich:1998:TM

[GGLR98] O. Goldreich, S. Goldwassert, E. Lehman, and D. Ron. Testing monotonicity. In IEEE [IEE98], pages 426–435. CO-DEN ASFPDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Goldreich:1996:PTC

[GGR96] O. Goldreich, S. Goldwasser, and D. Ron. Property testing and its connection to learning and approximation. In IEEE [IEE96], pages 339–348. CODEN ASFPDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Gupta:1991:CPI

[GI91] A. Gupta and R. Impagliazzo. Computing planar intertwines. In IEEE [IEE91], pages 802–811. CODEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Goel:1999:SLB

[GI99] A. Goel and P. Indyk. Stochastic load balancing and related problems. In IEEE [IEE99], pages 579–586. CODEN ASF-PDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Giesbrecht:1992:FAM

[Gie92] M. Giesbrecht. Fast algorithms for matrix normal forms. In IEEE [IEE92], pages 121–130. CODEN ASFPDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Goldreich:1990:SPA

[GIL^{+}90] O. Goldreich, R. Impagliazzo, L. Levin, R. Venkatesan, and D. Zuckerman. Security preserving amplification of hardness. In IEEE [IEE90], pages 318–326. CODEN ASF-PDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Gillman:1993:CBR

[Gil93] D. Gillman. A Chernoff bound for random walks on expander graphs. In IEEE [IEE93], pages 680–691. CODEN ASF-PDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Goldman:1999:AAP

[GIP99] D. Goldman, S. Istrail, and C. H. Papadimitriou. Algorithmic aspects of protein structure similarity. In IEEE [IEE99], pages 512–521. CO-DEN ASFPDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Grigoriev:1991:AAN

[GK91] D. Grigoriev and M. Karpinski. An approximation algorithm for the number of zeros or arbitrary polynomials over GF(q). In IEEE [IEE91], pages 662–669. CODEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Garg:1998:FSA

[GK98] N. Garg and J. Konemann. Faster and simpler algorithms for multicommodity flow and other fractional packing problems. In IEEE [IEE98], pages 300–309. CODEN ASF-PDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Garay:1993:SLT

[GKP93] J. A. Garay, S. Kutten, and D. Peleg. A sub-linear time distributed algorithm for minimum-weight spanning trees. In IEEE [IEE93], pages 659–668. CODEN ASF-PDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Grigni:1995:ASP

[GKP95] M. Grigni, E. Koutsoupias, and C. Papadimitriou. An approximation scheme for planar graph TSP. In IEEE [IEE95], pages 640–645. CO-DEN ASFPDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Goldman:1990:EIC

[GKS90] S. A. Goldman, M. J. Kearns, and R. E. Schapire. Exact identification of circuits using fixed points of amplification functions. In IEEE [IEE90], pages 193–202. CODEN ASF-PDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Grigoriev:1995:ILB

[GKV95] D. Grigoriev, M. Karpinski, and N. Vorobjov. Improved lower bound on testing membership to a polyhedron by algebraic decision trees. In IEEE [IEE95], pages 258–265. CO-DEN ASFPDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Gottlob:1998:CAC

[GLS98] G. Gottlob, N. Leone, and F. Scarcello. The complexity of acyclic conjunctive queries. In IEEE [IEE98], pages 706– 715. CODEN ASFPDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Guruswami:1998:TCN

[GLST98] V. Guruswami, D. Lewin, M. Sudan, and L. Trevisan. A tight characterization of NP with 3 query PCPs. In IEEE [IEE98], pages 8–17. CODEN ASFPDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Gradel:1992:HTC

[GM92] E. Gradel and G. L. Mc-Colm. Hierarchies in transitive closure logic, stratified Datalog and infinitary logic. In IEEE [IEE92], pages 167– 176. CODEN ASFPDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Goldberg:1997:CRG

[GM97] L. A. Goldberg and P. D.

MacKenzie. Contention resolution with guaranteed constant expected delay. In IEEE [IEE97], pages 213–222. CO-DEN ASFPDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Gil:1991:TTN

[GMV91] J. Gil, Y. Matias, and U. Vishkin. Towards a theory of nearly constant time parallel algorithms. In IEEE [IEE91], pages 698–710. CO-DEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Galil:1995:RMC

[GMY95] Z. Galil, A. Mayer, and Moti Yung. Resolving message complexity of Byzantine Agreement and beyond. In IEEE [IEE95], pages 724–733. CO-DEN ASFPDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Gupta:1999:CTE

[GNRS99] A. Gupta, I. Newman, Y. Rabinovich, and A. Sinclair. Cuts, trees and l_{1}-embeddings of graphs. In IEEE [IEE99], pages 399–408. CODEN ASF-PDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Goldwasser:1994:P

[Gol94] Shafi Goldwasser, editor. Proceedings: 35th Annual Symposium on Foundations of Computer Science, November 20– 22, 1994, Santa Fe, New Mexico. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. CODEN ASFPDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Goldwasser:1997:NDC

[Gol97] S. Goldwasser. New directions in cryptography: twenty some years later (or cryptography and complexity theory: a match made in heaven). In IEEE [IEE97], pages 314– 324. CODEN ASFPDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Goodrich:1991:UAA

[Goo91] M. T. Goodrich. Using approximation algorithms to design parallel algorithms that may ignore processor allocation. In IEEE [IEE91], pages 711–722. CODEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Gottlob:1993:NTC

[Got93] G. Gottlob. NP trees and Carnap’s modal logic. In IEEE [IEE93], pages 42–51. CODEN ASFPDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Goldreich:1991:QKC

[GP91] O. Goldreich and E. Petrank.

Quantifying knowledge complexity. In IEEE [IEE91], pages 59–68. CODEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Galil:1992:TAI

[GP92] Z. Galil and K. Park. Truly alphabet-independent twodimensional pattern matching. In IEEE [IEE92], pages 247– 256. CODEN ASFPDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Gupta:1993:GAC

[GP93] V. Gupta and V. Pratt. Gates accept concurrent behavior. In IEEE [IEE93], pages 62–71. CODEN ASFPDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Greene:1994:MIH

[GPY94] D. Greene, M. Parnas, and F. Yao. Multi-index hashing for information retrieval. In Goldwasser [Gol94], pages 722– 731. CODEN ASFPDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Goldberg:1997:BFD

[GR97a] A. V. Goldberg and S. Rao. Beyond the flow decomposition barrier. In IEEE [IEE97], pages 2–11. CODEN ASF-PDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Goldberg:1997:FUU

[GR97b] A. V. Goldberg and S. Rao.

Flows in undirected unit capacity networks. In IEEE [IEE97], pages 32–34. CODEN ASFPDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Grigoriev:1998:ECL

[GR98] D. Grigoriev and A. A.

Razborov. Exponential complexity lower bounds for depth 3 arithmetic circuits in algebras of functions over finite fields. In IEEE [IEE98], pages 269–278. CODEN ASF-PDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Grigoriev:1998:TTL

[Gri98] D. Grigoriev. Tseitin’s tautologies and lower bounds for Nullstellensatz proofs. In IEEE [IEE98], pages 648–652. CO-DEN ASFPDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Grolmusz:1992:SCC

[Gro92] V. Grolmusz. Separating the communication complexities of MOD m and MOD p circuits. In IEEE [IEE92], pages 278– 287. CODEN ASFPDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Grohe:1996:EFV

[Gro96] M. Grohe. Equivalence in finite-variable logics is complete for polynomial time. In IEEE [IEE96], pages 264–273. CODEN ASFPDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Goldreich:1995:LPQ

[GRS95] O. Goldreich, R. Rubinfeld, and M. Sudan. Learning polynomials with queries: The highly noisy case. In IEEE [IEE95], pages 294–303. CO-DEN ASFPDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Guruswami:1998:IDR

[GS98] V. Guruswami and M. Sudan. Improved decoding of Reed–Solomon and algebraicgeometric codes. In IEEE [IEE98], pages 28–37. CODEN ASFPDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Garg:1994:FSC

[GSV94] N. Garg, H. Saran, and V. V.

Vazirani. Finding separator cuts in planar graphs within twice the optimal. In Goldwasser [Gol94], pages 14–23. CODEN ASFPDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Grolmusz:1998:LBM

[GT98] V. Grolmusz and G. Tardos. Lower bounds for (MOD p– MOD m) circuits. In IEEE [IEE98], pages 279–288. CO-DEN ASFPDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Gilleron:1993:SSS

[GTT93] R. Gilleron, S. Tison, and M. Tommasi. Solving systems of set constraints with negated subset relationships. In IEEE [IEE93], pages 372– 380. CODEN ASFPDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Goodrich:1993:EMC

[GTVV93] M. T. Goodrich, Jyh-Jong Tsay, D. E. Vengroff, and J. S. Vitter. External-memory computational geometry. In IEEE [IEE93], pages 714–723. CODEN ASFPDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Gurevich:1990:MDP

[Gur90] Y. Gurevich. Matrix decomposition problem is complete for the average case. In IEEE [IEE90], pages 802–811. CO-DEN ASFPDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Gutierrez:1998:SWE

[Gut98] C. Gutierrez. Satisfiability of word equations with constants is in exponential space. In IEEE [IEE98], pages 112– 119. CODEN ASFPDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Grigoriev:1994:CLB

[GV94] D. Grigoriev and N. Vorobjov. Complexity lower bounds for computation trees with elementary transcendental function gates. In Goldwasser [Gol94], pages 548–552. CO-DEN ASFPDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Gartner:1994:RSA

[GZ94] B. Gartner and G. M. Ziegler. Randomized simplex algorithms on Klee-Minty cubes. In Goldwasser [Gol94], pages 502–510. CODEN ASFPDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Haken:1995:CBS

[Hak95] A. Haken. Counting bottlenecks to show monotone P ≠ NP. In IEEE [IEE95], pages 36–40. CODEN ASF-PDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Hall:1995:AFS

[Hal95] L. A. Hall. Approximability of flow shop scheduling. In IEEE [IEE95], pages 82–91. CODEN ASFPDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Haastad:1993:SE

[Hås93] J. Håstad. The shrinkage exponent is 2. In IEEE [IEE93], pages 114–123. CODEN ASF-PDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Haastad:1996:CHA

[Hås96] J. Håstad. Clique is hard to approximate within n^{1−ϵ}. In IEEE [IEE96], pages 627–636. CODEN ASFPDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Havlicek:1997:COW

[Hav97] J. Havlicek. Computable obstructions to wait-free computability. In IEEE [IEE97], pages 80–89. CODEN ASF-PDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Henzinger:1994:FDC

[Hen94]  M. R. Henzinger.  Fully

dynamic cycle-equivalence in graphs. In Goldwasser [Gol94], pages 744–755. CODEN ASF-PDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Hampapuram:1993:OBW

[HF93] H. Hampapuram and M. L. Fredman. Optimal bi-weighted binary trees and the complexity of maintaining partial sums. In IEEE [IEE93], pages 480– 485. CODEN ASFPDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Haastad:1990:PSD

[HG90] J. Håstad and M. Goldmann. On the power of small-depth threshold circuits. In IEEE [IEE90], pages 610–618. CO-DEN ASFPDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Horng:1990:SNR

[HH90] G. Horng and M.-D. A. Huang.

Simplifying nested radicals and solving polynomials by radicals in minimum depth. In IEEE [IEE90], pages 847–856. CODEN ASFPDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Henzinger:1995:CSF

[HHK95] M. R. Henzinger, T. A. Henzinger, and P. W. Kopke. Computing simulations on finite and infinite graphs. In IEEE [IEE95], pages 453–462. CO-DEN ASFPDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Huang:1991:EAR

[HI91] M.-D. Huang and D. Ierardi.

Efficient algorithms for the Riemann–Roch problem and for addition in the Jacobian of a curve. In IEEE [IEE91], pages 678–687. CODEN ASF-PDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Huang:1993:CRP

[HI93] Ming-Deh Huang and D. Ierardi. Counting rational points on curves over finite fields. In IEEE [IEE93], pages 616– 625. CODEN ASFPDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Hirshfeld:1994:PTA

[HJM94] Y. Hirshfeld, M. Jerrum, and F. Moller. A polynomialtime algorithm for deciding equivalence of normed contextfree processes. In Goldwasser [Gol94], pages 623–631. CO-DEN ASFPDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Haastad:1993:TLB

[HJP93] J. Håstad, S. Jukna, and P. Pudlak. Top-down lower bounds for depth 3 circuits. In IEEE [IEE93], pages 124– 129. CODEN ASFPDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Henzinger:1995:FDB

[HK95] M. R. Henzinger and V. King.

Fully dynamic biconnectivity and transitive closure. In IEEE [IEE95], pages 664–672. CO-DEN ASFPDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Hoffmann:1991:AGT

[HKK91] F. Hoffmann, M. Kaufmann, and K. Kriegel. The art gallery theorem for polygons with holes. In IEEE [IEE91], pages 39–48. CODEN ASF-PDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Helmbold:1992:ATN

[HLL92] D. P. Helmbold, N. Littlestone, and P. M. Long. Apple tasting and nearly one-sided learning. In IEEE [IEE92], pages 493– 502. CODEN ASFPDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Hass:1997:CCK

[HLP97] J. Hass, J. C. Lagarias, and N. Pippenger. The computational complexity of knot and link problems. In IEEE [IEE97], pages 172–181. CO-DEN ASFPDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Haastad:1998:SIR

[HN98] J. Håstad and M. Naslund. The security of individual RSA bits. In IEEE [IEE98], pages 510– 519. CODEN ASFPDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Hannenhalli:1995:TMM

[HP95] S. Hannenhalli and P. A. Pevzner. Transforming men into mice (polynomial algorithm for genomic distance problem). In IEEE [IEE95], pages 581–592. CODEN ASF-PDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Har-Peled:1999:TWP

[HP99] S. Har-Peled. Taking a walk in a planar arrangement. In IEEE [IEE99], pages 100–110. CODEN ASFPDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Harel:1990:DPN

[HR90] D. Harel and D. Raz. Deciding properties of nonregular programs. In IEEE [IEE90], pages 652–661. CODEN ASF-PDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Hsu:1991:LTA

[HR91] T.-S. Hsu and V. Ramachandran. A linear time algorithm for triconnectivity augmentation. In IEEE [IEE91], pages 548–559. CODEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Hagerup:1992:WMH

[HR92] T. Hagerup and R. Raman.

Waste makes haste: tight bounds for loose parallel sorting. In IEEE [IEE92], pages 628–637. CODEN ASF-PDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Henzinffer:1996:CVC

[HRG96] M. R. Henzinffer, S. Rao, and H. N. Gabow. Computing vertex connectivity: new bounds from old techniques. In IEEE [IEE96], pages 462–471. CO-DEN ASFPDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Halldorsson:1993:DVU

[HRS93] M. M. Halldorsson, J. Radhakrishnan, and K. V. Subrahmanyam. Directed vs. undirected monotone contact networks for threshold functions. In IEEE [IEE93], pages 604– 613. CODEN ASFPDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Halperin:1993:NQB

[HS93a] D. Halperin and M. Sharir. Near-quadratic bounds for the motion planning problem for a polygon in a polygonal environment. In IEEE [IEE93], pages 382–391. CODEN ASF-PDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Hershberger:1993:ECE

[HS93b] J. Hershberger and S. Suri.

Efficient computation of Euclidean shortest paths in the plane. In IEEE [IEE93], pages 508–517. CODEN ASF-PDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Hsu:1992:FCT

[Hsu92] T. Hsu. On four-connecting a triconnected graph. In IEEE [IEE92], pages 70–79. CODEN ASFPDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Herlihy:1991:LCL

[HSW91] M. Herlihy, N. Shavit, and O. Waarts. Low contention linearizable counting. In IEEE [IEE91], pages 526–535. CO-DEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428.

LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Huang:1996:SSP

[HW96] Ming-Deh Huang and Yiu-Chung Wong. Solving systems of polynomial congruences modulo a large prime. In IEEE [IEE96], pages 115– 124. CODEN ASFPDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Hemaspaandra:1997:MPB

[HW97] E. Hemaspaandra and G. Wechsung. The minimization problem for Boolean formulas. In IEEE [IEE97], pages 575–584. CODEN ASFPDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

IEEE:1990:PAS

[IEE90] IEEE, editor. Proceedings: 31st Annual Symposium on Foundations of Computer Science: October 22–24, 1990, St. Louis, Missouri. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1990. CODEN ASFPDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

IEEE:1991:PAS

[IEE91] IEEE, editor. Proceedings:

32nd annual Symposium on Foundations of Computer Science, San Juan, Puerto Rico, October 1–4, 1991. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1991. CODEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

IEEE:1992:ASF

[IEE92] IEEE, editor. 33rd Annual Symposium on Foundations of Computer Science: October 24–27, 1992, Pittsburgh, Pennsylvania: proceedings [papers]. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1992. CODEN ASFPDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0.

IEEE Computer Society Press Order Number 2900.

IEEE:1993:ASF

[IEE93] IEEE, editor. 34th Annual Symposium on Foundations of Computer Science: November 3–5, 1993, Palo Alto, California: proceedings [papers]. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. CODEN ASFPDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

IEEE:1995:ASF

[IEE95] IEEE, editor. 36th Annual Symposium on Foundations of Computer Science: October 23–25, 1995, Milwaukee, Wisconsin. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. CODEN ASFPDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

IEEE:1996:ASF

[IEE96] IEEE, editor. 37th Annual Symposium on Foundations of Computer Science: October 14–16, 1996, Burlington, Vermont. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. CODEN ASFPDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

IEEE:1997:ASF

[IEE97] IEEE, editor. 38th Annual Symposium on Foundations of Computer Science: October 20–22, 1997, Miami Beach, Florida. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. CODEN ASFPDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

IEEE:1998:ASF

[IEE98] IEEE, editor. 39th Annual Symposium on Foundations of Computer Science: proceedings: November 8–11, 1998, Palo Alto, California. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1998. CODEN ASFPDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

IEEE:1999:ASF

[IEE99] IEEE, editor. 40th Annual Symposium on Foundations of Computer Science: October 17–19, 1999, New York City, New York. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1999. CODEN ASFPDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Itkis:1994:FLS

[IL94] G. Itkis and L. Levin. Fast and lean self-stabilizing asynchronous protocols. In Goldwasser [Gol94], pages 226–239. CODEN ASFPDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Impagliazzo:1990:NBW

[Imp90] R. Impagliazzo. No better ways to generate hard NP instances than picking uniformly at random. In IEEE [IEE90], pages 812–821. CODEN ASF-PDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Impagliazzo:1995:HCD

[Imp95] R. Impagliazzo. Hard-core distributions for somewhat hard problems. In IEEE [IEE95], pages 538–545. CODEN ASF-PDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Indyk:1997:DSC

[Ind97] P. Indyk. Deterministic superimposed coding with applications to pattern matching. In IEEE [IEE97], pages 127– 136. CODEN ASFPDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Indyk:1998:FAS

[Ind98a] P. Indyk. Faster algorithms for string matching problems: matching the convolution bound. In IEEE [IEE98], pages 166–173. CODEN ASF-PDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Indyk:1998:ANN

[Ind98b] P. Indyk. On approximate nearest neighbors in non-Euclidean spaces. In IEEE [IEE98], pages 148–155. CO-DEN ASFPDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Indyk:1999:STA

[Ind99] P. Indyk. A sublinear time approximation scheme for clustering in metric spaces. In IEEE [IEE99], pages 154–159. CODEN ASFPDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Impagliazzo:1998:WPS

[IPZ98] R. Impagliazzo, R. Paturi, and F. Zane. Which problems have strongly exponential complexity? In IEEE [IEE98], pages 653–662. CODEN ASF-PDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Irani:1993:VIC

[IR93] S. Irani and Y. Rabani. On the value of information in coordination games. In IEEE [IEE93], pages 12–21. CODEN ASFPDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Irani:1990:CIG

[Ira90]  S. Irani.  Coloring induc-

tive graphs on-line. In IEEE [IEE90], pages 470–479. CO-DEN ASFPDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Istrail:1990:CGU

[Ist90] S. Istrail. Constructing generalized universal traversing sequences of polynomial size for graphs with small diameter. In IEEE [IEE90], pages 439–448. CODEN ASFPDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Impagliazzo:1999:NOC

[ISW99] R. Impagliazzo, R. Shaltiel, and A. Wigderson. Nearoptimal conversion of hardness into pseudo-randomness. In IEEE [IEE99], pages 181– 190. CODEN ASFPDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Impagliazzo:1998:RVT

[IW98] R. Impagliazzo and A. Wigderson. Randomness vs. time: de-randomization under a uniform assumption. In IEEE [IEE98], pages 734–743. CO-DEN ASFPDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Jackson:1994:EMQ

[Jac94]  J. Jackson.  An efficient

membership-query algorithm for learning DNF with respect to the uniform distribution. In Goldwasser [Gol94], pages 42– 53. CODEN ASFPDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Jain:1998:FAA

[Jai98] K. Jain. A factor 2 approximation algorithm for the generalized Steiner network problem. In IEEE [IEE98], pages 448– 457. CODEN ASFPDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Jayanti:1992:FTW

[JCT92] P. Jayanti, T. D. Chandra, and S. Toueg. Fault-tolerant wait-free shared objects. In IEEE [IEE92], pages 157–166. CODEN ASFPDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Jiang:1998:OQA

[JKL98] Tao Jiang, P. Kearney, and Ming Li. Orchestrating quartets: approximation and data correction. In IEEE [IEE98], pages 416–425. CODEN ASF-PDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Juedes:1993:CDH

[JL93] D. W. Juedes and J. H. Lutz. The complexity and distribution of hard problems. In IEEE [IEE93], pages 177–185. CODEN ASFPDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Johnson:1991:CCP

[JM91] D. B. Johnson and P. Metaxas. Connected components in O(lg^{3/2 mod V} mod ) parallel time for the CREW PRAM. In IEEE [IEE91], pages 688–697. CODEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Jerrum:1993:SAG

[JS93] M. Jerrum and G. B. Sorkin. Simulated annealing for graph bisection. In IEEE [IEE93], pages 94–103. CODEN ASF-PDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Jerrum:1992:MEA

[JV92] M. Jerrum and U. Vazirani.

A mildly exponential approximation algorithm for the permanent. In IEEE [IEE92], pages 320–326. CODEN ASF-PDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Jain:1999:PDA

[JV99] K. Jain and V. V. Vazirani.

Primal-dual approximation algorithms for metric facility location and k-median problems. In IEEE [IEE99], pages 2– 13. CODEN ASFPDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Keselman:1994:MAS

[KA94] D. Keselman and A. Amir.

Maximum agreement subtree in a set of evolutionary treesmetrics and efficient algorithms. In Goldwasser [Gol94], pages 758–769. CODEN ASF-PDV. ISBN 0-8186-6582-3.

ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Kahale:1991:BER

[Kah91] N. Kahale. Better expansion for Ramanujan graphs. In IEEE [IEE91], pages 398–404. CODEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Kahale:1992:SEL

[Kah92] N. Kahale. On the second eigenvalue and linear expansion of regular graphs. In IEEE [IEE92], pages 296–303. CODEN ASFPDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Kant:1992:DPG

[Kan92] G. Kant. Drawing planar graphs using the lmc-ordering. In IEEE [IEE92], pages 101– 110. CODEN ASFPDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Kannan:1994:MCP

[Kan94] R. Kannan. Markov chains and polynomial time algorithms. In Goldwasser [Gol94], pages 656– 671. CODEN ASFPDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Karker:1993:RSM

[Kar93] D. R. Karker. Random sampling in matroids, with applications to graph connectivity and minimum spanning trees. In IEEE [IEE93], pages 84– 93. CODEN ASFPDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Klein:1990:ATM

[KARR90] P. Klein, A. Agrawal, R. Ravi, and S. Rao. Approximation through multicommodity flow. In IEEE [IEE90], pages 726–737. CODEN ASF-PDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Kapron:1991:NCM

[KC91] B. Kapron and S. A. Cook. A new characterization of Mehlhorn’s polynomial time functionals. In IEEE [IEE91], pages 342–347. CODEN ASF-PDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Kosaraju:1990:TPT

[KD90] S. R. Kosaraju and A. L. Delcher. A tree-partitioning technique with applications to expression evaluation and term matching. In IEEE [IEE90], pages 163–172. CODEN ASF-PDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Kearns:1998:TTI

[Kea98] M. Kearns. A tutorial on theoretical issues in probabilistic artificial intelligence. In IEEE [IEE98], page 4. CODEN ASFPDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Kaminski:1990:FMA

[KF90] M. Kaminski and N. Francez.

Finite-memory automata. In IEEE [IEE90], pages 683–688. CODEN ASFPDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Kilian:1994:CBI

[Kil94] J. Kilian. On the complexity of bounded-interaction and noninteractive zero-knowledge proofs. In Goldwasser [Gol94], pages 466–477. CODEN ASF-PDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

King:1999:FDA

[Kin99] V. King. Fully dynamic algorithms for maintaining allpairs shortest paths and transitive closure in digraphs. In IEEE [IEE99], pages 81–89. CODEN ASFPDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Kenyon:1992:TPR

[KK92] C. Kenyon and R. Kenyon. Tiling a polygon with rectangles. In IEEE [IEE92], pages 610–619. CODEN ASF-PDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Karger:1994:RCS

[KK94] D. R. Karger and D. Koller. (de)randomized construction of small sample spaces in N C. In Goldwasser [Gol94], pages 252–263. CODEN ASFPDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Kimbrel:1996:NOP

[KK96] T. Kimbrel and A. R. Karlin. Near-optimal parallel prefetching and caching. In IEEE [IEE96], pages 540–549. CO-DEN ASFPDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Kolpakov:1999:FMR

[KK99] R. Kolpakov and G. Kucherov.

Finding maximal repetitions in a word in linear time. In IEEE [IEE99], pages 596–604. CODEN ASFPDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Kaklamanis:1990:ATB

[KKL^{+}90] C. Kaklamanis, A. R. Karlin, F. T. Leighton, V. Milenkovic, P. Raghavan, S. Rao, C. Thomborson, and A. Tsantilas. Asymptotically tight bounds for computing with faulty arrays of processors. In IEEE [IEE90], pages 285–296. CO-DEN ASFPDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Karger:1991:FHP

[KKP91] D. R. Karger, D. Koller, and S. J. Phillips. Finding the hidden path: time bounds for all-pairs shortest paths. In IEEE [IEE91], pages 560–568. CODEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no.

Kaklamanis:1993:UES

[KKR93] C. Kaklamanis, D. Krizanc, and S. Rao. Universal emulations with sublogarithmic slowdown. In IEEE [IEE93], pages 341–350. CODEN ASF-PDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Kannan:1996:SAM

[KL96] R. Kannan and Guangxing Li. Sampling according to the multivariate normal density. In IEEE [IEE96], pages 204–212. CODEN ASFPDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Klarlund:1991:PMC

[Kla91] N. Klarlund. Progress measures for complementation ωautomata with applications to temporal logic. In IEEE [IEE91], pages 358–367. CO-DEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Klein:1991:WUS

[Kle91] R. Klein. Walking an unknown street with bounded detour. In IEEE [IEE91], pages 304–313. CODEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Kleinberg:1994:LPM

[Kle94] J. M. Kleinberg. The localization problem for mobile robots. In Goldwasser [Gol94], pages 521–531. CODEN ASFPDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Kleinberg:1996:SSU

[Kle96] J. M. Kleinberg. Single-source unsplittable flow. In IEEE [IEE96], pages 68–77. CODEN ASFPDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Kleitman:1994:DRB

[KLM94] D. Kleitman, T. Leighton, and Y. Ma. On the design of reliable Boolean circuits that contain partially unreliable gates. In Goldwasser [Gol94], pages 332–346. CODEN ASFPDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Kutylowski:1994:FFP

[KLOW94] M. Kutylowski, K. Lorys, B. Oesterdiekhoff, and R. Wanka. Fast and feasible periodic sorting networks of constant depth. In Goldwasser [Gol94], pages 369–380. CODEN ASFPDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Kushilevitz:1994:RCM

[KMO94] E. Kushilevitz, S. Micali, and R. Ostrovsky. Reducibility and completeness in multiparty private computations. In Goldwasser [Gol94], pages 478– 489. CODEN ASFPDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Kamath:1994:TBO

[KMPS94] A. Kamath, R. Motwani, K. Palem, and P. Spirakis. Tail bounds for occupancy and the satisfiability threshold conjecture. In Goldwasser [Gol94], pages 592–603. CODEN ASF-PDV. ISBN 0-8186-6582-3.

ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Kleinberg:1997:SME

[KMRV97] J. Kleinberg, R. Motwani, P. Raghavan, and S. Venkatasubramanian. Storage management for evolving databases. In IEEE [IEE97], pages 353– 362. CODEN ASFPDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Karger:1994:AGC

[KMS94] D. Karger, R. Motwani, and M. Sudan. Approximate graph coloring by semidefinite programming. In Goldwasser [Gol94], pages 2–13. CODEN ASFPDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Khanna:1994:SVC

[KMSV94] S. Khanna, R. Motwani, M. Sudan, and U. Vazirani. On syntactic versus computational views of approximability. In Goldwasser [Gol94], pages 819–830. CODEN ASFPDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Kannan:1998:LSS

[KN98] R. Kannan and A. Nolte. Local search in smooth convex sets. In IEEE [IEE98], pages 218– 226. CODEN ASFPDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Kushilevitz:1997:RNS

[KO97] E. Kushilevitz and R. Ostrovsky. Replication is not needed: single database, computationally-private information retrieval. In IEEE [IEE97], pages 364–373. CO-DEN ASFPDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Kogan:1996:CPF

[Kog96] G. Kogan. Computing permanents over fields of characteristic 3: where and why it becomes difficult. In IEEE [IEE96], pages 108–114. CO-DEN ASFPDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Koiran:1993:WVB

[Koi93] P. Koiran. A weak version of the Blum, Shub and Smale model. In IEEE [IEE93], pages 486–495. CODEN ASF-PDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Koiran:1995:AVD

[Koi95]  P. Koiran.  Approximat-

ing the volume of definable sets. In IEEE [IEE95], pages 134–141. CODEN ASF-PDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Koiran:1997:RDA

[Koi97] P. Koiran. Randomized and deterministic algorithms for the dimension of algebraic varieties. In IEEE [IEE97], pages 36–45. CODEN ASF-PDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Kosaraju:1995:FAC

[Kos95]  S. Rao Kosaraju.  Faster

algorithms for the construction of parameterized suffix trees. In IEEE [IEE95], pages 631–638. CODEN ASF-PDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Koutsoupias:1999:WAK

[Kou99] E. Koutsoupias. Weak adversaries for the k-server problem. In IEEE [IEE99], pages 444– 449. CODEN ASFPDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Koscielski:1990:CUF

[KP90] A. Koscielski and L. Pacholski. Complexity of unification in free groups and free semi-groups. In IEEE [IEE90], pages 824–829. CODEN ASF-PDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Kaltofen:1992:PEP

[KP92] E. Kaltofen and V. Pan.

Processor-efficient parallel solution of linear systems. II. The positive characteristic and singular cases. In IEEE [IEE92], pages 714–723. CODEN ASF-PDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Kortsarz:1993:CDS

[KP93] G. Kortsarz and D. Peleg.

On choosing a dense subgraph. In IEEE [IEE93], pages 692– 701. CODEN ASFPDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Koutsoupias:1994:BCA

[KP94] E. Koutsoupias and C. H. Papadimitriou. Beyond competitive analysis [on-line algorithms]. In Goldwasser [Gol94], pages 394–400. CODEN ASF-PDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Kalyanasundaram:1995:SPC

[KP95a] B. Kalyanasundaram and K. Pruhs. Speed is as powerful as clairvoyance [scheduling problems]. In IEEE [IEE95], pages 214–221. CODEN ASF-PDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Krause:1995:CBF

[KP95b] M. Krause and P. Pudlak. On computing Boolean functions by sparse real polynomials. In IEEE [IEE95], pages 682–691. CODEN ASF-PDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Kutten:1995:TFL

[KP95c] S. Kutten and D. Peleg. Tight fault locality. In IEEE [IEE95], pages 704–713. CODEN ASF-PDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Kalyanasundaram:1997:MFT

[KP97a] B. Kalyanasundaram and K. R. Pruhs. Minimizing flow time nonclairvoyantly. In IEEE [IEE97], pages 345–352. CODEN ASFPDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Khachiyan:1997:CIP

[KP97b] L. Khachiyan and L. Porkolab. Computing integral points in convex semi-algebraic sets. In IEEE [IEE97], pages 162– 171. CODEN ASFPDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Karlin:1992:MP

[KPR92] A. R. Karlin, S. J. Phillips, and P. Raghavan. Markov paging. In IEEE [IEE92], pages 208– 217. CODEN ASFPDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number

Kilian:1998:LBZ

[KPR98] J. Kilian, E. Petrank, and C. Rackoff. Lower bounds for zero knowledge on the Internet. In IEEE [IEE98], pages 484– 492. CODEN ASFPDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Kozen:1992:EIP

[KPS92] D. Kozen, J. Palsberg, and M. I. Schwartzbach. Efficient inference of partial types. In IEEE [IEE92], pages 363– 371. CODEN ASFPDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Kosaraju:1994:LTS

[KPS94] S. R. Kosaraju, J. K. Park, and C. Stein. Long tours and short superstrings. In Goldwasser [Gol94], pages 166– 177. CODEN ASFPDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Kenyon:1996:ASP

[KR96a] C. Kenyon and E. Remila. Approximate strip packing. In IEEE [IEE96], pages 31–36. CODEN ASFPDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Kleinberg:1996:SPE

[KR96b] J. Kleinberg and R. Rubinfeld. Short paths in expander graphs. In IEEE [IEE96], pages 86–95. CODEN ASF-PDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Kumar:1999:MCV

[KR99] V. S. Anil Kumar and H. Ramesh. Markovian coupling vs. conductance for the Jerrum-Sinclair chain. In IEEE [IEE99], pages 241–251. CO-DEN ASFPDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Kumar:1998:RSP

[KRRT98] R. Kumar, P. Raghavan, S. Rajagopalan, and A. Tomkins. Recommendation systems: a probabilistic analysis. In IEEE [IEE98], pages 664–673. CO-DEN ASFPDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Kleinberg:1999:FRL

[KRT99] J. Kleinberg, Y. Rabani, and E. Tardos. Fairness in routing and load balancing. In IEEE [IEE99], pages 568–578. CODEN ASFPDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Kearns:1990:EDF

[KS90] M. J. Kearns and R. E. Schapire. Efficient distributionfree learning of probabilistic concepts. In IEEE [IEE90], pages 382–391. CODEN ASF-PDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Klein:1993:LPP

[KS93] P. N. Klein and S. Subramanian. A linear-processor polylog-time algorithm for shortest paths in planar graphs. In IEEE [IEE93], pages 259–270. CODEN ASF-PDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Kumar:1996:EST

[KS96] S. R. Kumar and D. Sivakumar. Efficient self-testing/self-correction of linear recurrences. In IEEE [IEE96], pages 602– 611. CODEN ASFPDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Kolliopoulos:1997:IAA

[KS97] S. G. Kolliopoulos and C. Stein.

Improved approximation algorithms for unsplittable flow problems. In IEEE [IEE97], pages 426–436. CODEN ASF-PDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Klivans:1999:BHC

[KS99] A. R. Klivans and R. A. Servedio. Boosting and hardcore sets. In IEEE [IEE99], pages 624–633. CODEN ASF-PDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Kaplan:1994:TPC

[KST94] H. Kaplan, R. Shamir, and R. E. Tarjan. Tractability of parameterized completion problems on chordal and interval graphs: minimum fillin and physical mapping. In Goldwasser [Gol94], pages 780– 791. CODEN ASFPDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Kim:1999:LEO

[KST99] Jeong Han Kim, D. R. Simon, and P. Tetali. Limits on the efficiency of oneway permutation-based hash functions. In IEEE [IEE99], pages 535–542. CODEN ASF-PDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Kzeinberg:1995:DPD

[KT95] J. Kzeinberg and E. Tardos.

Disjoint paths in densely embedded graphs. In IEEE [IEE95], pages 52–61. CODEN ASFPDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Katoh:1999:LLT

[KT99a] N. Katoh and T. Tokuyama.

Lovasz’s lemma for the threedimensional K-level of concave surfaces and its applications. In IEEE [IEE99], pages 389– 398. CODEN ASFPDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Kleinberg:1999:AAC

[KT99b] J. Kleinberg and E. Tardos.

Approximation algorithms for classification problems with pairwise relationships: metric labeling and Markov random fields. In IEEE [IEE99], pages 14–23. CODEN ASF-PDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Kanevsky:1991:LMF

[KTDC91] A. Kanevsky, R. Tamassia, G. Di Battista, and J. Chen. On-line maintenance of the four-connected components of a graph. In IEEE [IEE91], pages 793–801. CODEN ASF-PDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Katoh:1992:MMS

[KTI92] N. Katoh, T. Tokuyama, and K. Iwano. On minimum and maximum spanning trees of linearly moving points. In IEEE [IEE92], pages 396–405. CODEN ASFPDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Kunde:1991:CRD

[Kun91] M. Kunde. Concentrated regular data streams on grids: sorting and routing near to the bisection bound. In IEEE [IEE91], pages 141–150. CO-DEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Kannan:1990:IEH

[KW90] S. Kannan and T. Warnow.

Inferring evolutionary history from DNA sequences. In IEEE [IEE90], pages 362–371. CO-DEN ASFPDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Krause:1991:VRC

[KW91] M. Krause and S. Waack. Variation ranks of communication matrices and lower bounds for depth two circuits having symmetric gates with unbounded fan-in. In IEEE [IEE91], pages 777–782. CODEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Kondacs:1997:PQF

[KW97] A. Kondacs and J. Watrous.

On the power of quantum finite state automata. In IEEE [IEE97], pages 66–75. CODEN ASFPDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Karp:1995:BVI

[KWZ95] R. M. Karp, O. Waarts, and G. Zweig. The bit vector intersection problem. In IEEE [IEE95], pages 621–630. CO-DEN ASFPDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Karloff:1997:AAM

[KZ97] H. Karloff and U. Zwick. A 7/8-approximation algorithm for MAX 3SAT? In IEEE [IEE97], pages 406–415. CO-DEN ASFPDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Lagergren:1990:EPA

[Lag90] J. Lagergren. Efficient parallel algorithms for treedecomposition and related problems. In IEEE [IEE90], pages 173–182. CODEN ASF-PDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Latombe:1995:CRC

[Lat95] J.-C. Latombe. Controllability, recognizability, and complexity issues in robot motion planning. In IEEE [IEE95], pages 484–500. CODEN ASF-PDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Leivant:1998:CNT

[Lei98] D. Leivant. A characterization of NC by tree recurrence. In IEEE [IEE98], pages 716– 724. CODEN ASFPDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Lund:1990:AMI

[LFKN90] C. Lund, L. Fortnow, H. Karloff, and N. Nisan. Algebraic methods for interactive proof systems. In IEEE [IEE90], pages 2–10. CODEN ASF-PDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Li:1990:TDS

[Li90] M. Li. Towards a DNA sequencing theory (learning a string). In IEEE [IEE90], pages 125–134. CODEN ASF-PDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Linial:1994:GGS

[LLR94] N. Linial, E. London, and Y. Rabinovich. The geometry of graphs and some of its algorithmic applications. In Goldwasser [Gol94], pages 577– 591. CODEN ASFPDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Lipton:1997:CSU

[LMN97] R. J. Lipton, P. J. Martino, and A. Neitzke. On the complexity of a set-union problem. In IEEE [IEE97], pages 110– 115. CODEN ASFPDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Leighton:1991:HFT

[LMP91] T. Leighton, Y. Ma, and C. G.

Plaxton. Highly fault-tolerant sorting circuits. In IEEE [IEE91], pages 458–469. CO-DEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Leighton:1992:FTS

[LMS92] T. Leighton, B. Maggs, and R. Sitaraman. On the fault tolerance of some popular bounded-degree networks. In IEEE [IEE92], pages 542–552. CODEN ASFPDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Lincoln:1990:DPP

[LMSS90] P. Lincoln, J. Michell, A. Scedrov, and N. Shankar. Decision problems for propositional linear logic. In IEEE [IEE90], pages 662–671. CODEN ASF-PDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Lovasz:1991:SPD

[LNNW91] L. Lovasz, M. Naor, I. Newman, and A. Wigderson. Search problems in the decision tree model. In IEEE [IEE91], pages 576–585. CO-DEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Lokam:1995:SMM

[Lok95] S. V. Lokam. Spectral methods for matrix rigidity with applications to size-depth tradeoffs and communication complexity. In IEEE [IEE95], pages 6–15. CODEN ASF-PDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Leighton:1990:FSC

[LP90] T. Leighton and C. G. Plaxton. A (fairly) simple circuit that (usually) sorts. In IEEE [IEE90], pages 264–274. CODEN ASFPDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Lund:1994:ICO

[LPR94] C. Lund, S. Phillips, and N. Reingold. IP over connection-oriented networks and distributional paging. In Goldwasser [Gol94], pages 424– 434. CODEN ASFPDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Luby:1995:MCA

[LRS95] M. Luby, D. Randall, and A. Sinclair. Markov chain algorithms for planar lattice structures. In IEEE [IEE95], pages 150–159. CODEN ASF-PDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Leiserson:1993:ECA

[LRT93] C. E. Leiserson, S. Rao, and S. Toledo. Efficient out-of-core algorithms for linear relaxation using blocking covers. In IEEE [IEE93], pages 704– 713. CODEN ASFPDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Lovasz:1990:MRM

[LS90] L. Lovasz and M. Simonovits. The mixing rate of Markov chains, an isoperimetric inequality, and computing the volume. In IEEE [IEE90], pages 346–354. CODEN ASF-PDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Lapidot:1991:FPM

[LS91a] D. Lapidot and A. Shamir. Fully parallelized multi prover protocols for NEXP-time. In IEEE [IEE91], pages 13–18. CODEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Leighton:1991:EAD

[LS91b] T. Leighton and E. J. Schwabe. Efficient algorithms for dynamic allocation of distributed memory. In IEEE [IEE91], pages 470–479. CODEN ASF-PDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Lenhof:1992:EKC

[LS92a] H.-P. Lenhof and M. Smid.

Enumerating the k closest pairs optimally. In IEEE [IEE92], pages 380–386. CO-DEN ASFPDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Lovasz:1992:RCV

[LS92b] L. Lovasz and M. Simonovits.

On the randomized complexity of volume and diameter. In IEEE [IEE92], pages 482– 492. CODEN ASFPDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Luo:1990:CCA

[LT90] Z.-Q. Luo and J. N. Tsitsiklis. Communication complexity of algebraic computation. In IEEE [IEE90], pages 758–765. CODEN ASF-PDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Luczak:1999:RNC

[LU99] M. J. Luczak and E. Upfal. Reducing network congestion and blocking probability through balanced allocation. In IEEE [IEE99], pages 587– 595. CODEN ASFPDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Luks:1992:CSM

[Luk92] E. M. Luks. Computing in solvable matrix groups. In IEEE [IEE92], pages 111–120. CODEN ASFPDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Lipton:1999:CS

[LV99] R. J. Lipton and A. Viglas. On the complexity of SAT. In IEEE [IEE99], pages 459– 464. CODEN ASFPDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Lynch:1990:PSA

[Lyn90] J. F. Lynch. Probabilities of sentences about very sparse random graphs. In IEEE [IEE90], pages 689–696. CO-DEN ASFPDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Maggs:1997:ELD

[MadHVW97] B. M. Maggs, F. Meyer auf der Heide, B. Vocking, and M. Westermann. Exploiting locality for data management in systems of limited bandwidth. In IEEE [IEE97], pages 284– 293. CODEN ASFPDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Matousek:1991:RPH

[Mat91] J. Matousek. Reporting points in halfspaces. In IEEE [IEE91], pages 207–215. CODEN ASF-PDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Matousek:1998:GCA

[Mat98] J. Matousek. Geometric computation and the art of sampling (tutorial). In IEEE [IEE98], page 2. CODEN ASFPDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Micali:1994:CP

[Mic94] S. Micali. CS proofs. In Goldwasser [Gol94], pages 436– 453. CODEN ASFPDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Micciancio:1998:SVL

[Mic98] D. Micciancio. The shortest vector in a lattice is hard to approximate to within some constant. In IEEE [IEE98], pages 92–98. CODEN ASF-PDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Mitchell:1993:RTP

[Mit93] S. A. Mitchell. Refining a triangulation of a planar straightline graph to eliminate large angles. In IEEE [IEE93], pages 583–592. CODEN ASF-PDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Mitzenmacher:1996:LBD

[Mit96] M. Mitzenmacher. Load balancing and density dependent jump Markov processes. In IEEE [IEE96], pages 213–222. CODEN ASFPDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Mihail:1995:EAO

[MKR95] M. Mihail, C. Kaklamanis, and Satish Rao. Efficient access to optical bandwidth wavelength routing on directed fiber trees, rings, and trees of rings. In IEEE [IEE95], pages 548–557. CODEN ASF-PDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

McAllister:1993:CPL

[MKS93] M. McAllister, D. Kirkpatrick, and J. Snoeyink. A compact piecewise-linear Voronoi diagram for convex sites in the plane. In IEEE [IEE93], pages 573–582. CODEN ASF-PDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Matousek:1991:FTD

[MMP^{+}91] J. Matousek, N. Miller, J. Pach, M. Sharir, S. Sifrony, and E. Welzl. Fat triangles determine linearly many holes (computational geometry). In IEEE [IEE91], pages 49–58. CODEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Mitchell:1998:LCB

[MMS98] J. Mitchell, M. Mitchell, and A. Scedrov. A linguistic characterization of bounded oracle computation and probabilistic polynomial time. In IEEE [IEE98], pages 725–733. CODEN ASFPDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Morgenstern:1991:ECN

[Mor91] M. Morgenstern. Explicit construction of natural bounded concentrators. In IEEE [IEE91], pages 392–397. CO-DEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Maler:1990:TBC

[MP90] O. Maler and A. Pnueli. Tight bounds on the complexity of cascaded decomposition of automata. In IEEE [IEE90], pages 672–682. CODEN ASF-PDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Marcinkowski:1992:UHC

[MP92] J. Marcinkowski and L. Pacholski. Undecidability of the Horn-clause implication problem. In IEEE [IEE92], pages 354–362. CODEN ASF-PDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Mitchell:1992:CSL

[MPA92] J. S. B. Mitchell, C. Piatko, and E. M. Arkin. Computing a shortest k-link path in a polygon. In IEEE [IEE92], pages 573–582. CODEN ASF-PDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Mansour:1998:JCQ

[MPS98] Y. Mansour and B. Patt-Shamir. Jitter control in QoS networks. In IEEE [IEE98], pages 50–59. CODEN ASF-PDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Miltersen:1992:ACM

[MPT92] P. B. Miltersen, M. Paterson, and J. Tarui. The asymptotic complexity of merging networks. In IEEE [IEE92], pages 236–246. CODEN ASF-PDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Mahajan:1995:DSP

[MR95] S. Mahajan and H. Ramesh. Derandomizing semidefinite programming based approximation algorithms. In IEEE [IEE95], pages 162–169. CO-DEN ASFPDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Madras:1996:FGB

[MR96] N. Madras and D. Randall.

Factoring graphs to bound mixing rates. In IEEE [IEE96], pages 194–203. CODEN ASF-PDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Munro:1997:SRB

[MR97] J. I. Munro and V. Raman.

Succinct representation of balanced parentheses, static trees and planar graphs. In IEEE [IEE97], pages 118–126. CO-DEN ASFPDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Muthukrishnan:1999:OSM

[MRSG99] S. Muthukrishnan, R. Rajaraman, A. Shaheen, and J. E. Gehrke. Online scheduling to minimize average stretch. In IEEE [IEE99], pages 433– 443. CODEN ASFPDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Micali:1999:VRF

[MRV99] S. Micali, M. Rabin, and S. Vadhan. Verifiable random functions. In IEEE [IEE99], pages 120–130. CODEN ASF-PDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Margaritis:1995:RSS

[MS95] D. Margaritis and S. S. Skiena. Reconstructing strings from substrings in rounds. In IEEE [IEE95], pages 613–620. CO-DEN ASFPDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Morris:1999:RWT

[MS99] B. Morris and A. Sinclair. Random walks on truncated cubes and sampling 0–1 knapsack solutions. In IEEE [IEE99], pages 230–240. CODEN ASF-PDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Martel:1990:APA

[MSP90] C. Martel, R. Subramonian, and A. Part. Asynchronous PRAMs are (almost) as good as synchronous PRAMs. In IEEE [IEE90], pages 590–599. CODEN ASFPDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Maass:1991:CPS

[MSS91] W. Maass, G. Schnitger, and E. D. Sontag. On the computational power of sigmoid versus Boolean threshold circuits. In IEEE [IEE91], pages 767–776. CODEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Maass:1990:CLC

[MT90] W. Maass and G. Turan. On the complexity of learning from counterexamples and membership queries. In IEEE [IEE90], pages 203–210. CODEN ASF-PDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Miller:1991:UGA

[MTV91] G. L. Miller, S.-H. Teng, and S. A. Vavasis. A unified geometric approach to graph separators. In IEEE [IEE91], pages 538–547. CODEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Mulmuley:1991:RMSb

[Mul91a] K. Mulmuley. Randomized multidimensional search trees: further results in dynamic sampling. In IEEE [IEE91], pages 216–227. CODEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Mulmuley:1991:RMSa

[Mul91b] K. Mulmuley. Randomized multidimensional search trees: lazy balancing and dynamic shuffling. In IEEE [IEE91], pages 180–196. CODEN ASF-PDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Mulmuley:1992:RGA

[Mul92]  K. Mulmuley.  Random-

ized geometric algorithms and pseudo-random generators. In IEEE [IEE92], pages 90–100. CODEN ASFPDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Miltersen:1999:DAM

[MV99] P. B. Miltersen and N. V.

Vinodchandran. Derandomizing Arthur–Merlin games using hitting sets. In IEEE [IEE99], pages 71–80. CODEN ASFPDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Matousek:1991:DAB

[MWW91] J. Matousek, E. Welzl, and L. Wernisch. Discrepancy and in-approximations for bounded VC-dimension. In IEEE [IEE91], pages 424–430. CO-DEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Mayers:1998:QCI

[MY98] D. Mayers and A. Yao. Quantum cryptography with imperfect apparatus. In IEEE [IEE98], pages 503–509. CO-DEN ASFPDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Nayak:1999:OLB

[Nay99] A. Nayak. Optimal lower bounds for quantum automata and random access codes. In IEEE [IEE99], pages 369–376. CODEN ASFPDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Neff:1990:SPP

[Nef90] C. A. Neff. Specified precision polynomial root isolation is in NC. In IEEE [IEE90], pages 152–162. CODEN ASF-PDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Naor:1990:FAO

[NGM90] D. Naor, D. Gusfield, and C. Martel. A fast algorithm for optimally increasing the edgeconnectivity. In IEEE [IEE90], pages 698–707. CODEN ASF-PDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Naor:1991:OFS

[NR91] M. Naor and R. M. Roth.

Optimal file sharing in distributed networks. In IEEE [IEE91], pages 515–525. CO-DEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Neff:1994:ACR

[NR94] C. A. Neff and J. H. Reif. An O(n^{1+ϵ} log b) algorithm for the complex roots problem. In Goldwasser [Gol94], pages 540– 547. CODEN ASFPDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Naor:1995:STA

[NR95] M. Naor and O. Reingold.

Synthesizers and their application to the parallel construction of pseudo-random functions. In IEEE [IEE95], pages 170–181. CODEN ASF-PDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Naor:1997:NTC

[NR97] M. Naor and O. Reingold. Number-theoretic constructions of efficient pseudorandom functions. In IEEE [IEE97], pages 458–467. CO-DEN ASFPDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Nisan:1994:PHB

[NRS94] N. Nisan, S. Rudich, and M. Saks. Products and help bits in decision trees. In Goldwasser [Gol94], pages 318– 329. CODEN ASFPDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Naor:1997:IAS

[NS97] J. Naor and B. Schieber. Improved approximations for shallow-light spanning trees. In IEEE [IEE97], pages 536– 541. CODEN ASFPDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Naor:1995:SNO

[NSS95] M. Naor, L. J. Schulman, and A. Srinivasan. Splitters and near-optimal derandomization. In IEEE [IEE95], pages 182–191. CODEN ASF-PDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Nisan:1992:UCS

[NSW92] N. Nisan, E. Szemeredi, and A. Wigderson. Undirected connectivity in O(log^{1.5} n) space. In IEEE [IEE92], pages 24– 29. CODEN ASFPDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Naor:1994:LCA

[NW94] M. Naor and A. Wool. The load, capacity and availability of quorum systems. In Goldwasser [Gol94], pages 214– 225. CODEN ASFPDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Nisan:1995:LBA

[NW95] N. Nisan and A. Wigderson. Lower bounds on arithmetic circuits via partial derivatives. In IEEE [IEE95], pages 16–25. CODEN ASF-PDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Naor:1997:AAD

[NZ97] J. Naor and L. Zosin. A 2approximation algorithm for the directed multiway cut problem. In IEEE [IEE97], pages 548–553. CODEN ASF-PDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Ogihara:1995:SPH

[Ogi95] M. Ogihara. Sparse P-hard sets yield space-efficient algorithms. In IEEE [IEE95], pages 354–361. CODEN ASF-PDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Okasaki:1995:ALE

[Oka95] C. Okasaki. Amortization, lazy evaluation, and persistence: lists with catenation via lazy linking. In IEEE [IEE95], pages 646–654. CO-DEN ASFPDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Olshevsky:1998:USA

[OP98] V. Olshevsky and V. Pan.

A unified superfast algorithm for boundary rational tangential interpolation problems and for inversion and factorization of dense structured matrices. In IEEE [IEE98], pages 192– 201. CODEN ASFPDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Orlitsky:1995:CC

[OR95] A. Orlitsky and J. R. Roche.

Coding for computing. In IEEE [IEE95], pages 502–511. CODEN ASFPDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Orlitsky:1991:ICB

[Orl91] A. Orlitsky. Interactive communication: balanced distributions, correlated files, and average-case complexity. In IEEE [IEE91], pages 228–238. CODEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Papadimitriou:1990:GTL

[Pap90] C. H. Papadimitriou. On graph-theoretic lemmata and complexity classes. In IEEE [IEE90], pages 794–801. CO-DEN ASFPDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Papadimitriou:1991:SST

[Pap91] C. H. Papadimitriou. On selecting a satisfying truth assignment. In IEEE [IEE91], pages 163–169. CODEN ASF-PDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Plaxton:1995:TBD

[Pla95] C. G. Plaxton. Tight bounds for a distributed selection game with applications to fixedconnection machines. In IEEE [IEE95], pages 114–122. CO-DEN ASFPDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Plandowski:1999:SWE

[Pla99] W. Plandowski. Satisfiability of word equations with constants is in PSPACE. In IEEE [IEE99], pages 495–500. CODEN ASFPDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Pnueli:1997:TDT

[Pnu97] A. Pnueli. Two decades of temporal logic: Achievements and challenges. In IEEE [IEE97], page 78. CODEN ASFPDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Poon:1993:SBG

[Poo93] C. K. Poon. Space bounds for graph connectivity problems on node-named JAGs and node-ordered JAGs. In IEEE [IEE93], pages 218–227. CO-DEN ASFPDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Plaxton:1992:ILB

[PPS92] C. G. Plaxton, B. Poonen, and T. Suel. Improved lower bounds for Shellsort. In IEEE [IEE92], pages 226–235. CO-DEN ASFPDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Paturi:1998:IET

[PPSZ98] R. Paturi, P. Pudlik, M. E. Saks, and F. Zane. An improved exponential-time algorithm for k-SAT. In IEEE [IEE98], pages 628–637. CO-DEN ASFPDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Paterson:1990:FCS

[PPZ90] M. S. Paterson, N. Pippenger, and U. Zwick. Faster circuits and shorter formulae for multiple addition, multiplication and symmetric Boolean functions. In IEEE [IEE90], pages 642–650. CODEN ASF-PDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Paturi:1997:SCL

[PPZ97] R. Paturi, P. Pudlak, and F. Zane. Satisfiability coding lemma. In IEEE [IEE97], pages 566–574. CODEN ASF-PDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Pneuli:1990:DRS

[PR90] A. Pneuli and R. Rosner. Distributed reactive systems are hard to synthesize. In IEEE [IEE90], pages 746–757. CO-DEN ASFPDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Plaxton:1996:FFT

[PR96] C. G. Plaxton and R. Rajaraman. Fast fault-tolerant concurrent access to shared objects. In IEEE [IEE96], pages 570–579. CODEN ASF-PDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Pagter:1998:OTS

[PR98] J. Pagter and T. Rauhe. Optimal time-space trade-offs for sorting. In IEEE [IEE98], pages 264–268. CODEN ASF-PDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Peleg:1999:NTL

[PR99] D. Peleg and V. Rubinovich. A near-tight lower bound on the time complexity of distributed MST construction. In IEEE [IEE99], pages 253–261. CODEN ASFPDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Papadimitriou:1994:MPG

[PRST94] C. H. Papadimitriou, P. Raghavan, M. Sudan, and H. Tamaki. Motion planning on a graph. In Goldwasser [Gol94], pages 511–520. CODEN ASFPDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Pan:1992:PCT

[PRT92] V. Y. Pan, J. H. Reif, and S. R.

Tate. The power of combining the techniques of algebraic and numerical computing: improved approximate multipoint polynomial evaluation and improved multipole algorithms. In IEEE [IEE92], pages 703– 713. CODEN ASFPDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Paturi:1990:TCP

[PS90] R. Paturi and M. E. Saks. On threshold circuits for parity. In IEEE [IEE90], pages 397– 404. CODEN ASFPDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Paterson:1995:CRB

[PS95] M. Paterson and A. Srinivasan. Contention resolution with bounded delay. In IEEE [IEE95], pages 104–113. CO-DEN ASFPDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Plotkin:1991:FAA

[PST91] S. A. Plotkin, D. B. Shmoys, and E. Tardos. Fast approximation algorithms for fractional packing and covering problems. In IEEE [IEE91], pages 495–504. CODEN ASF-PDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Petrank:1996:KC

[PT96] E. Petrank and G. Tardos. On the knowledge complexity of N P. In IEEE [IEE96], pages 494–503. CODEN ASF-PDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Pach:1998:WCN

[PT98] J. Pach and G. Toth. Which crossing number is it, anyway? [computational geometry]. In IEEE [IEE98], pages 617–626. CODEN ASFPDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Pitassi:1992:CHC

[PU92] T. Pitassi and A. Urquhart.

The complexity of the Hajos calculus. In IEEE [IEE92], pages 187–196. CODEN ASF-PDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Paterson:1991:SMF

[PZ91] M. S. Paterson and U. Zwick.

Shrinkage of de Morgan formulae under restriction. In IEEE [IEE91], pages 324–333. CODEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991.

IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Reinhardt:1997:MNU

[RA97] K. Reinhardt and E. Allender. Making nondeterminism unambiguous. In IEEE [IEE97], pages 244–253. CO-DEN ASFPDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Rabani:1996:PCM

[Rab96a] Y. Rabani. Path coloring on the mesh. In IEEE [IEE96], pages 400–409. CODEN ASF-PDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Rabin:1996:CHA

[Rab96b] M. O. Rabin. Computationally hard algebraic problems. In IEEE [IEE96], pages 284– 289. CODEN ASFPDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Radhakrishnan:1991:BBT

[Rad91]  J. Radhakrishnan.  Better

bounds for threshold formulas. In IEEE [IEE91], pages 314–323. CODEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Radzik:1992:NMF

[Rad92] T. Radzik. Newton’s method for fractional combinatorial optimization. In IEEE [IEE92], pages 659–669. CODEN ASF-PDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Radzik:1993:FAG

[Rad93] T. Radzik. Faster algorithms for the generalized network flow problem. In IEEE [IEE93], pages 438–448. CODEN ASF-PDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Rauch:1992:FDB

[Rau92] M. Rauch. Fully dynamic biconnectivity in graphs. In IEEE [IEE92], pages 50–59. CODEN ASFPDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Ravi:1994:RRR

[Rav94] R. Ravi. Rapid rumor ramification: approximating the minimum broadcast time. In Goldwasser [Gol94], pages 202– 213. CODEN ASFPDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Reif:1993:ARR

[Rei93] J. H. Reif. An O(n log^{3} n) algorithm for the real root problem. In IEEE [IEE93], pages 626–635. CODEN ASF-PDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Reif:1995:EPS

[Rei95] J. H. Reif. Efficient parallel solution of sparse eigenvalue and eigenvector problems. In IEEE [IEE95], pages 123–132. CO-DEN ASFPDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Raz:1997:SMN

[RM97] R. Raz and P. McKenzie. Separation of the monotone NC hierarchy. In IEEE [IEE97], pages 234–243. CODEN ASF-PDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Raz:1999:ERE

[RRV99] R. Raz, O. Reingold, and S. Vadhan. Error reduction for extractors. In IEEE [IEE99], pages 191–201. CODEN ASF-PDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Reischuk:1991:RCN

[RS91] R. Reischuk and B. Schmeltz.

Reliable computation with noisy circuits and decision trees — a general n log n lower bound. In IEEE [IEE91], pages 602–611. CODEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no.

91CH3062-7. Computer Society order no. 2445.

Raz:1993:LRC

[RS93] R. Raz and B. Spieker. On the “log rank”-conjecture in communication complexity. In IEEE [IEE93], pages 168–176. CODEN ASFPDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Rajagopalan:1996:VI

[RS96] S. Rajagopalan and L. J. Schulman. Verifying identities. In IEEE [IEE96], pages 612–616. CODEN ASFPDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Radhakrishnan:1998:IBA

[RS98] J. Radhakrishnan and A. Srinivasan. Improved bounds and algorithms for hypergraph twocoloring. In IEEE [IEE98], pages 684–693. CODEN ASF-PDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Regan:1995:PGM

[RSC95] K. W. Regan, D. Sivakumar, and Jin-Yi Cai. Pseudorandom generators, measure theory, and natural proofs. In IEEE [IEE95], pages 26–35. CODEN ASFPDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Rabinovich:1992:QDS

[RSW92] Y. Rabinovich, A. Sinclair, and A. Wigderson. Quadratic dynamical systems. In IEEE [IEE92], pages 304–313. CO-DEN ASFPDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Ranade:1994:NTB

[RSW94] A. Ranade, S. Schleimer, and D. S. Wilkerson. Nearly tight bounds for wormhole routing. In Goldwasser [Gol94], pages 347–355. CODEN ASFPDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Rabani:1998:LDM

[RSW98] Y. Rabani, A. Sinclair, and R. Wanka. Local divergence of Markov chains and the analysis of iterative load-balancing schemes. In IEEE [IEE98], pages 694–703. CODEN ASF-PDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Radhakrishnan:1997:TBD

[RTS97] J. Radhakrishnan and A. Ta-Shma. Tight bounds for depthtwo superconcentrators. In IEEE [IEE97], pages 585–594. CODEN ASFPDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Reif:1990:CCO

[RTY90] J. H. Reif, J. D. Tygar, and A. Yoshida. The computability and complexity of optical beam tracing. In IEEE [IEE90], pages 106–114. CODEN ASF-PDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Rubinfeld:1994:RFE

[Rub94] R. Rubinfeld. On the robustness of functional equations. In Goldwasser [Gol94], pages 288–299. CODEN ASFPDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Rajagopalan:1993:PDR

[RV93] S. Rajagopalan and V. V. Vazirani. Primal-dual RNC approximation algorithms for (multi)- set (multi)-cover and covering integer programs. In IEEE [IEE93], pages 322–331. CO-DEN ASFPDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Russell:1998:PIL

[RZ98] A. Russell and D. Zuckerman.

Perfect information leader election in log n + O(1) rounds. In IEEE [IEE98], pages 576– 583. CODEN ASFPDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Sahai:1999:NMN

[Sah99] A. Sahai. Non-malleable noninteractive zero knowledge and adaptive chosen-ciphertext security. In IEEE [IEE99], pages 543–553. CODEN ASF-PDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Schwarzkopf:1991:DMG

[Sch91] O. Schwarzkopf. Dynamic maintenance of geometric structures made easy. In IEEE [IEE91], pages 197–206. CO-DEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Schwentick:1994:GCM

[Sch94] T. Schwentick. Graph connectivity and monadic NP. In Goldwasser [Gol94], pages 614– 622. CODEN ASFPDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Schenk:1995:FAA

[Sch95] E. Schenk. Faster approximate agreement with multiwriter registers. In IEEE [IEE95], pages 714–723. CO-DEN ASFPDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Schonhage:1998:MCT

[Sch98] A. Schonhage. Multiplicative complexity of Taylor shifts and a new twist of the substitution method. In IEEE [IEE98], pages 212–215. CODEN ASF-PDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Schoning:1999:PAK

[Sch99] T. Schoning. A probabilistic algorithm for k-SAT and constraint satisfaction problems. In IEEE [IEE99], pages 410– 414. CODEN ASFPDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Senizergues:1998:DBE

[Sen98] G. Senizergues. Decidability of bisimulation equivalence for equational graphs of finite out-degree. In IEEE [IEE98], pages 120–129. CODEN ASF-PDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Shamir:1990:IIP

[Sha90]  A. Shamir.  IP=PSPACE

(interactive proof=polynomial space). In IEEE [IEE90], pages 11–15. CODEN ASF-PDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Sharir:1993:ATU

[Sha93] M. Sharir. Almost tight upper bounds for lower envelopes in higher dimensions. In IEEE [IEE93], pages 498–507. CO-DEN ASFPDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Shor:1991:HPB

[Sho91] P. W. Shor. How to pack better than best fit: tight bounds for average-case online bin packing. In IEEE [IEE91], pages 752–759. CODEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Shor:1994:AQC

[Sho94] P. W. Shor. Algorithms for quantum computation: discrete logarithms and factoring. In Goldwasser [Gol94], pages 124–134. CODEN ASF-PDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Shor:1996:FTQ

[Sho96] P. W. Shor. Fault-tolerant quantum computation. In IEEE [IEE96], pages 56–65. CODEN ASFPDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Simon:1994:PQC

[Sim94] D. R. Simon. On the power of quantum computation. In Goldwasser [Gol94], pages 116– 123. CODEN ASFPDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Skutella:1998:SRP

[Sku98] M. Skutella. Semidefinite relaxations for parallel machine scheduling. In IEEE [IEE98], pages 472–481. CODEN ASF-PDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Smolensky:1990:IAF

[Smo90] R. Smolensky. On interpolation by analytic functions with special properties and some weak lower bounds on the size of circuits with symmetric gates. In IEEE [IEE90], pages 628–631. CODEN ASF-PDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Smolensky:1993:RLD

[Smo93]  R. Smolensky.  On repre-

sentations by low-degree polynomials. In IEEE [IEE93], pages 130–138. CODEN ASF-PDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Spielman:1996:HFT

[Spi96] D. A. Spielman. Highly faulttolerant parallel computation. In IEEE [IEE96], pages 154– 163. CODEN ASFPDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Shallcross:1993:NEP

[SPLK93] D. Shallcross, V. Pan, and Yu Lin-Kriz. The NC equivalence of planar integer linear programming and Euclidean GCD. In IEEE [IEE93], pages 557–564. CODEN ASF-PDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Srinivasan:1997:IAE

[Sri97] A. Srinivasan. Improved approximations for edge-disjoint paths, unsplittable flow, and related routing problems. In IEEE [IEE97], pages 416–425. CODEN ASFPDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Shoup:1991:LBP

[SS91] V. Shoup and R. Smolensky. Lower bounds for polynomial evaluation and interpolation problems. In IEEE [IEE91], pages 378–383. CO-DEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Sipser:1994:EC

[SS94] M. Sipser and D. A. Spielman. Expander codes. In Goldwasser [Gol94], pages 566– 576. CODEN ASFPDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Sander:1997:DPP

[SS97] T. Sander and M. A. Shokrollahi. Deciding properties of polynomials without factoring. In IEEE [IEE97], pages 46– 55. CODEN ASFPDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Sinha:1994:EOB

[ST94] R. K. Sinha and J. S.

Thathachar. Efficient oblivious branching programs for threshold functions. In Goldwasser [Gol94], pages 309–317. CODEN ASFPDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Spielmat:1996:SPW

[ST96] D. A. Spielmat and Shang-Hua Teng. Spectral partitioning works: planar graphs and finite element meshes. In IEEE [IEE96], pages 96–105. CODEN ASFPDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Sudan:1998:PCP

[ST98] M. Sudan and L. Trevisan. Probabilistically checkable proofs with low amortized query complexity. In IEEE [IEE98], pages 18–27. CODEN ASF-PDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Subramanian:1995:MCR

[Sub95] C. R. Subramanian. Minimum coloring random and semi-random graphs in polynomial expected time. In IEEE [IEE95], pages 463–472. CO-DEN ASFPDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Sudan:1996:MLD

[Sud96] M. Sudan. Maximum likelihood decoding of Reed Solomon codes. In IEEE [IEE96], pages 164–172. CO-DEN ASFPDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Sundar:1991:LBD

[Sun91] R. Sundar. A lower bound for the dictionary problem under a hashing model. In IEEE [IEE91], pages 612–621. CODEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Saran:1991:FCW

[SV91] H. Saran and V. V. Vazirani. Finding k-cuts within twice the optimal. In IEEE [IEE91], pages 743–751. CO-DEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Sahinalp:1996:EAD

[SV96] S. C. Sahinalp and U. Vishkin.

Efficient approximate and dynamic matching of patterns using a labeling paradigm. In IEEE [IEE96], pages 320–328. CODEN ASFPDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Sahai:1997:CPP

[SV97] A. Sahai and S. P. Vadhan. A complete promise problem for statistical zero-knowledge. In IEEE [IEE97], pages 448– 457. CODEN ASFPDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Smith:1998:GST

[SW98] W. D. Smith and N. C. Wormald. Geometric separator theorems and applications. In IEEE [IEE98], pages 232– 243. CODEN ASFPDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Shmoys:1991:SPM

[SWW91] D. B. Shmoys, J. Wein, and D. P. Williamson. Scheduling parallel machines on-line. In IEEE [IEE91], pages 131–140. CODEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Sander:1999:NIC

[SYY99] T. Sander, A. Young, and Moti Yung. Non-interactive cryptocomputing for NC^{1}. In IEEE [IEE99], pages 554–566. CODEN ASFPDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Sturtivant:1990:EIB

[SZ90] C. Sturtivant and Z.-L. Zhang.

Efficiently inverting bijections given by straight line programs. In IEEE [IEE90], pages 327–334. CODEN ASF-PDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Srinivasan:1994:CVW

[SZ94] A. Srinivasan and D. Zuckerman. Computing with very weak random sources. In Goldwasser [Gol94], pages 264–275. CODEN ASFPDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Saks:1995:RD

[SZ95] M. Saks and S. Zhou. RSPACE(S)_{e} DSPACE(S^{3/2}). In IEEE [IEE95], pages 344–353. CO-DEN ASFPDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Shoshan:1999:APS

[SZ99] A. Shoshan and U. Zwick. All pairs shortest paths in undirected graphs with integer weights. In IEEE [IEE99], pages 605–614. CODEN ASF-PDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Szegedy:1994:NNL

[Sze94] M. Szegedy. A note on the θ number of Lovasz and the generalized Delsarte bound. In Goldwasser [Gol94], pages 36– 39. CODEN ASFPDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Szegedy:1998:ATI

[Sze98] M. Szegedy. Algorithms to tile the infinite grid with finite clusters. In IEEE [IEE98], pages 137–145. CODEN ASF-PDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Tamaki:1992:ESE

[Tam92] H. Tamaki. Efficient selfembedding of butterfly networks with random faults. In IEEE [IEE92], pages 533–541. CODEN ASFPDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Thorup:1997:USS

[Tho97] M. Thorup. Undirected single source shortest paths in linear time. In IEEE [IEE97], pages 12–21. CODEN ASF-PDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Thorup:1998:MGP

[Tho98] M. Thorup. Map graphs in polynomial time. In IEEE [IEE98], pages 396–405. CO-DEN ASFPDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Toda:1990:CFM

[Tod90] S. Toda. The complexity of finding medians. In IEEE [IEE90], pages 778–787. CO-DEN ASFPDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Toledo:1992:MNL

[Tol92] S. Toledo. Maximizing nonlinear concave functions in fixed dimension. In IEEE [IEE92], pages 676–685. CO-DEN ASFPDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Torng:1995:UAP

[Tor95] E. Torng. A unified analysis of paging and caching. In IEEE [IEE95], pages 194–203. CO-DEN ASFPDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Trevisan:1996:GAL

[TSSW96] L. Trevisan, G. B. Sorkin, M. Sudan, and D. P. Williamson. Gadgets, approximation, and linear programming. In IEEE [IEE96], pages 617–626. CO-DEN ASFPDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Turan:1994:CBF

[TV94] G. Turan and F. Vatan. On the computation of Boolean functions by analog circuits of bounded fan-in. In Goldwasser [Gol94], pages 553–564. CODEN ASFPDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Therien:1996:TLS

[TW96] D. Therien and T. Wilke.

Temporal logic and semidirect products: an effective characterization of the until hierarchy. In IEEE [IEE96], pages 256–263. CODEN ASF-PDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

Teng:1993:ASS

[TY93] Shang-Hua Teng and F. Yao. Approximating shortest superstrings. In IEEE [IEE93], pages 158–165. CODEN ASF-PDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Umans:1997:HCS

[UL97] C. Umans and W. Lenhart. Hamiltonian cycles in solid grid graphs. In IEEE [IEE97], pages 496–505. CODEN ASF-PDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Umans:1998:MED

[Uma98] C. Umans. The minimum equivalent DNF problem and shortest implicants. In IEEE [IEE98], pages 556–563. CO-DEN ASFPDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Umans:1999:HAM

[Uma99] C. Umans. Hardness of approximating Σ_{2}^{p} minimization problems. In IEEE [IEE99], pages 465–474. CODEN ASF-PDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Unger:1998:CAB

[Ung98] W. Unger. The complexity of the approximation of the bandwidth problem. In IEEE [IEE98], pages 82–91. CODEN ASFPDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Varadarajan:1997:ASP

[VA97] K. R. Varadarajan and P. K.

Agarwal. Approximating shortest paths on a nonconvex polyhedron. In IEEE [IEE97], pages 182–191. CO-DEN ASFPDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Vaidya:1990:RPC

[Vai90]  P. M. Vaidya.  Reducing

the parallel complexity of certain linear programming problems. In IEEE [IEE90], pages 583–589. CODEN ASF-PDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Valiant:1995:CC

[Val95] L. G. Valiant. Cognitive computation. In IEEE [IEE95], pages 2–3. CODEN ASF-PDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Varadarajan:1998:DCA

[Var98] K. R. Varadarajan. A divideand-conquer algorithm for min-cost perfect matching in the plane. In IEEE [IEE98], pages 320–329. CODEN ASF-PDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428.

LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

vanDam:1998:QOI

[vD98] W. van Dam. Quantum oracle interrogation: getting all information for almost half the price. In IEEE [IEE98], pages 362–367. CODEN ASF-PDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Vempala:1997:RSB

[Vem97] S. Vempala. A random sampling based algorithm for learning the intersection of halfspaces. In IEEE [IEE97], pages 508–513. CODEN ASF-PDV. ISBN 0-8186-8197-7 (paperback), 0-8186-8198-5 (casebound), 0-8186-8199-3 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 .S92 1997. IEEE catalog number 97CB36150. IEEE Computer Society Press order number PR08197.

Vempala:1998:RPN

[Vem98] S. Vempala. Random projection: a new approach to VLSI layout. In IEEE [IEE98], pages 389–395. CODEN ASF-PDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Verma:1991:TUH

[Ver91] R. M. Verma. A theory of using history for equational systems with applications. In IEEE [IEE91], pages 348–357. CODEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Vallee:1990:LRA

[VF90] B. Vallee and P. Flajolet. The lattice reduction algorithm of Gauss: an average case analysis. In IEEE [IEE90], pages 830–839. CODEN ASF-PDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Vigoda:1999:IBS

[Vig99] E. Vigoda. Improved bounds for sampling colorings. In IEEE [IEE99], pages 51–59. CODEN ASFPDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Vishwanathan:1990:ROG

[Vis90] S. Vishwanathan. Randomized online graph coloring. In IEEE [IEE90], pages 464–469. CODEN ASFPDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Vitter:1991:OPD

[VK91] J. S. Vitter and P. Krishnan. Optimal prefetching via data compression. In IEEE [IEE91], pages 121–130. CO-DEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Vocking:1999:HAH

[Voc99] B. Vocking. How asymmetry helps load balancing. In IEEE [IEE99], pages 131–141. CODEN ASFPDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Vazirani:1996:EAC

[VSR96] V. V. Vazirani, H. Saran, and B. S. Rajan. An efficient algorithm for constructing minimal trellises for codes over finite Abelian groups. In IEEE [IEE96], pages 144–153. CO-DEN ASFPDV. ISBN 0-7803-3762-X (casebound), 0-8186-7594-2 (softbound), 0-8186-7596-9 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1996. IEEE catalog number 96CH35973. IEEE Computer Society Press order number PR07594.

VandenBussche:1992:COC

[VVAG92] J. Van den Bussche, D. Van Gucht, M. Andries, and M. Gyssens. On the completeness of object-creating query languages. In IEEE [IEE92], pages 372–379. CODEN ASF-PDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

vonzurGathen:1991:EEF

[vzG91] J. von zur Gathen. Efficient exponentiation in finite field. In IEEE [IEE91], pages 384–391. CODEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

vonzurGathen:1995:FPC

[vzGS95] J. von zur Gathen and I. Shparlinski. Finding points on curves over finite fields. In IEEE [IEE95], pages 284–292. CO-DEN ASFPDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Wagener:1992:OPH

[Wag92] H. Wagener. Optimal parallel hull construction for simple polygons in O(log log n) time. In IEEE [IEE92], pages 593– 599. CODEN ASFPDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Wang:1991:ATD

[Wan91] W. Wang. An asynchronous two-dimensional self-correcting cellular automaton. In IEEE [IEE91], pages 278–285. CO-DEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Watrous:1995:ODQ

[Wat95]  J. Watrous.  On one-

dimensional quantum cellular automata. In IEEE [IEE95], pages 528–537. CODEN ASF-PDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Watrous:1999:QCS

[Wat99a] J. Watrous. On quantum and classical space-bounded processes with algebraic transition amplitudes. In IEEE [IEE99], pages 341–351. CO-DEN ASFPDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Watrous:1999:PCR

[Wat99b] J. Watrous. PSPACE has constant-round quantum interactive proof systems. In IEEE [IEE99], pages 112–119. CODEN ASFPDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Weihe:1994:MFP

[Wei94] K. Weihe. Maximum (s,t)- flows in planar networks in O(—V—log—V—) time. In Goldwasser [Gol94], pages 178– 189. CODEN ASFPDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Weihe:1999:EDR

[Wei99] K. Weihe. Edge-disjoint routing in plane switch graphs in linear time. In IEEE [IEE99], pages 330–339. CODEN ASF-PDV. ISBN 0-7695-0409-4 (softbound), 0-7803-5955-0 (casebound), 0-7695-0411-6 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1999. IEEE Catalog Number 99CB37039.

Wu:1991:CCP

[WK91] I.-C. Wu and H. T. Kung.

Communication complexity for parallel divide-and-conquer. In IEEE [IEE91], pages 151–162. CODEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Woods:1998:USE

[Woo98] A. R. Woods. Unsatisfiable systems of equations, over a finite field. In IEEE [IEE98], pages 202–211. CODEN ASF-PDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.

Yannakakis:1995:PDT

[Yan95] M. Yannakakis. Perspectives on database theory. In IEEE [IEE95], pages 224–246. CO-DEN ASFPDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Yao:1990:ATC

[Yao90] A. C.-C. Yao. ON ACC and threshold circuits. In IEEE [IEE90], pages 619–627. CO-DEN ASFPDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Yao:1992:ADT

[Yao92] A. Chi-Chih Yao. Algebraic decision trees and Euler characteristics. In IEEE [IEE92], pages 268–277. CODEN ASF-PDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Yao:1993:QCC

[Yao93] A. Chi-Chih Yao. Quantum circuit complexity. In IEEE [IEE93], pages 352–361. CO-DEN ASFPDV. ISBN 0-8186-4370-6 (paperback), 0-8186-4371-4 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1993. IEEE Catalog Number 93CH3368-8. IEEE Computer Society Press Order Number 4372-02.

Yao:1994:LBM

[Yao94] A. C.-C. Yao. A lower bound for the monotone depth of connectivity. In Goldwasser [Gol94], pages 302–308. CO-DEN ASFPDV. ISBN 0-8186-6582-3. ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Yap:1992:FUR

[Yap92] C. K. Yap. Fast unimodular reduction: planar integer lattices. In IEEE [IEE92], pages 437–446. CODEN ASF-PDV. ISBN 0-8186-2901-0 (microfiche), 0-8186-2900-2 (paperback). ISSN 0272-5428. LCCN QA 76 S979 1992. IEEE Catalog Number 92CH3188-0. IEEE Computer Society Press Order Number 2900.

Cai:1994:CMP

[yCLZ94] Jin yi Cai, R. J. Lipton, and Y. Zalcstein. The complexity of the membership problem for 2-generated commutative semigroups of rational matrices. In Goldwasser [Gol94], pages 135–142. CODEN ASF-PDV. ISBN 0-8186-6582-3.

ISSN 0272-5428. LCCN QA 76 S979 1994. IEEE catalog number 94CH35717. IEEE Computer Society Press Order Number 6580-02.

Yao:1995:SMR

[YDS95] F. Yao, A. Demers, and S. Shenker. A scheduling model for reduced CPU energy. In IEEE [IEE95], pages 374–382. CODEN ASF-PDV. ISBN 0-7803-3121-4 (casebound), 0-8186-7183-1 (softbound), 0-8186-7184-X (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1995. IEEE catalog number 95CB35834.

Zuckerman:1990:GWR

[Zuc90] D. Zuckerman. General weak random sources. In IEEE [IEE90], pages 534–543. CO-DEN ASFPDV. ISBN 0-8186-2082-X (paperback), 0-8186-6082-1 (microfiche). ISSN 0272-5428. LCCN TK7885.A1 S92 1990. Formerly called the Annual Symposium on Switching and Automata Theory. IEEE catalog number 90CH29256. Computer Society order no. 2082.

Zuckerman:1991:SBU

[Zuc91] D. Zuckerman. Simulating BPP using a general weak random source. In IEEE [IEE91], pages 79–89. CODEN ASFPDV. ISBN 0-8186-2445-0. ISSN 0272-5428. LCCN TK7885.A1 S92 1991. IEEE Catalog no. 91CH3062-7. Computer Society order no. 2445.

Zwick:1998:APS

[Zwi98] U. Zwick. All pairs shortest paths in weighted directed graphs-exact and almost exact algorithms. In IEEE [IEE98], pages 310–319. CODEN ASF-PDV. ISBN 0-8186-9172-7 (softbound), 0-7803-5229-7 (casebound), 0-8186-9174-3 (microfiche). ISSN 0272-5428. LCCN QA267 .S95 1998 Sci-Eng. IEEE Catalog Number 98CB36280. IEEE Computer Society Press Order Number PR9172.